### v1

# Chaos Engineering: A Multi-Vocal Literature Review

Abstract.
Organizations, particularly medium and large enterprises, typically today rely heavily on complex, distributed systems to deliver critical services and products. However, the growing complexity of these systems poses challenges in ensuring service availability, performance, and reliability. Traditional resilience testing methods often fail to capture modern systems’ intricate interactions and failure modes. Chaos Engineering addresses these challenges by proactively testing how systems in production behave under turbulent conditions, allowing developers to uncover and resolve potential issues before they escalate into outages. Though chaos engineering has received growing attention from researchers and practitioners alike, we observed a lack of a comprehensive literature review. Hence, we performed a Multivocal Literature Review (MLR) on chaos engineering to fill this research gap by systematically analyzing 88 academic and grey literature sources published from January 2019 to April 2024. We first used the selected sources to derive a unified definition of chaos engineering and to identify key capabilities, components, and adoption drivers. We also developed a taxonomy for chaos engineering and compared the relevant tools using it. Finally, we analyzed the state of the current chaos engineering research and identified several open research issues.
Chaos Engineering, Fault Injection, Distributed System, Resilience, Multivocal Literature Review, Microservice
copyright: noneisbn: 978-1-4503-XXXX-X/18/06ccs: Software and its engineering Software testing and debugging
1. Introduction
Cloud-based applications are being adopted across various domains, including healthcare, finance, construction, education, and defense (Alabbadi, 2011; Dang et al., 2019; Bello et al., 2021). While cloud computing promises to offer benefits such as cost savings, flexibility, and faster time to market (Buyya et al., 2009), the growing complexity of cloud infrastructure poses a significant challenge to maintaining dependable systems, leading to an increased risk of technical failures and costly downtime (Song and Sohn, 2022; Li et al., 2011; Wang et al., 2015; Al-Said Ahmad et al., 2024). The consequences of these failures can be severe, as seen in significant financial losses incurred for companies, including the 2007-2013 outages of 38 cloud service providers, resulting in \$669 million in losses, and Facebook’s six-hour outage in 2021, which resulted in an estimated \$160 million loss for the global economy  (Gill, 2022). The 2017 ITIC (Information Technology Intelligence Consulting) survey also reported that 98% of organizations experienced downtime with financial losses exceeding \$100,000 per hour (Inc., 2023).
An effective approach to overcoming the challenges of ensuring reliable and resilient cloud-based systems is chaos engineering, which proactively tests system resilience by introducing controlled failures into production environments (Bhaskar, 2022; Fawcett, 2020; Palani and Gupta, 2023). This approach was pioneered by Netflix in 2008 to fortify its video streaming infrastructure. Since then, chaos engineering has grown and extended its influence, evolving from its original focus on cloud services and microservice architecture to broader applications across industries (Bairyev, 2023). Today, this approach is widely employed in critical sectors like fintech and healthcare, with major technology companies such as LinkedIn, Facebook, Google, Microsoft, Amazon, Slack, and others adopting it to enhance their system resilience, improve customer experience, and drive business growth.  (Inc., 2023; Patel, 2022; Roa, 2022; Fawcett, 2020; Al-Said Ahmad et al., 2024; Bremmers, 2021; Bairyev, 2023).
Figure 1 illustrates chaos engineering’s adoption through Moore’s ’Crossing the Chasm’ model (Moore and McKenna, 1999). Adoption begins with Innovators (early risk-takers) and Early Adopters (those seeking competitive advantage), followed by the Early Majority (who adopt once proven), the Late Majority (who wait until widely supported), and finally, Laggards (the last to adopt, often only when essential). The critical hurdle, Moore’s Chasm, is where many approaches fail due to complexity and unclear benefits. Crossing this chasm leads to mainstream adoption and integration across the SDLC (Software Development Lifecycle) and industries (Mukkara, 2022).
Figure 1. Chaos Engineering in the Context of Geoffrey Moore’s Technology Adoption Lifecycle Model from Crossing the Chasm  (Mukkara, 2022).
[…]
Chaos Engineering on the Technology Adoption Lifecycle
We observed many researchers and practitioners have explored chaos engineering in various contexts, discussing its implementations and challenges through research papers and grey literature such as online blogs, technical papers, and forums. However, despite its increasing adoption, much of the existing knowledge remains dispersed across the literature. To the authors’ best understanding, no secondary research has been published to synthesize this scattered information, such as comprehensive reviews or survey papers. Hence, to address this research gap, in this paper, we present a Multivocal Literature Review (MLR) (Garousi et al., 2019) on chaos engineering, synthesizing a broad spectrum of perspectives from academic and grey literature sources. Unlike traditional Systematic Literature Reviews (SLRs), which often overlook grey literature, an MLR includes insights from peer-reviewed academic studies alongside industry reports, technical blogs, and white papers (Garousi et al., 2019, 2016). In particular, as chaos engineering emerged from the industry and most tools are from the industry, the grey literature is a highly relevant and valuable source of information.
In this MLR, we selected and analyzed 88 sources, spanning academic and grey literature sources. Our findings provide a detailed overview of chaos engineering, including its core principles, quality requirements, and foundational components defined by researchers and industry experts. We also examine the motivations behind adopting chaos engineering, the specific resilience challenges it addresses, and the practices and evaluation methods that support its effective implementation. Additionally, we introduce a taxonomy that can help practitioners and decision-makers make informed decisions when comparing and selecting chaos engineering tools. By analyzing the benefits and shortcomings of current approaches in the chaos engineering domain, this MLR contributes to a thorough understanding of chaos engineering and provides guidelines for future studies to advance this critical field.
The remainder of the paper is organized as follows. Section  2 presents the research methodology used in this study, while Section 3 provides a comprehensive definition of chaos engineering, along with its functionalities, quality requirements, and core components. Section 4 focuses on the motivations for chaos engineering and the specific challenges it addresses. Section 5 introduces a taxonomy of chaos engineering, covering tools, practices, and evaluation methods within the field. Section 6 examines the current state of chaos engineering in research and industry. Section  7 discusses the implications of our findings, highlighting key insights, identifying open issues, and addressing potential threats to validity. Finally, Section 8 summarizes the related literature surveys, and Section  9 concludes the paper.
2. Research Methodology
We employed the guidelines proposed by Garousi et al. (Garousi et al., 2019) to systematically collect and analyze multivocal literature on chaos engineering. We also drew methodological inspiration from other multivocal literature reviews, such as Buck et al. (Buck et al., 2021) and Islam et al. (Islam et al., 2019). Figure  2 illustrates our MLR methodology, comprising several steps divided into three phases: planning and designing the review protocol, executing the review, and documenting the findings. The remainder of this section provides a detailed discussion of each step.
2.1. Research Identification
We employed a search strategy to identify appropriate literature grounded in a series of research questions in Table  1. The first three questions cover the various aspects of chaos engineering, including its definitions, key features, benefits, challenges, use cases, tools and techniques, and best and bad practices. The final question provides an overview of the present landscape of chaos engineering literature.
Figure 2. Overview of the MLR Process(Islam et al., 2019).
[…]
Multivocal Literature Review Process.
Table 1. Research Questions and Motivation
Research Questions Motivation
RQ1. How is chaos engineering defined in the literature? RQ1.1. Which core principles and primary functionalities are most commonly associated with chaos engineering? RQ1.2. What are the core components of a chaos engineering platform, and how do they support its core functions? RQ1.3. What quality requirements are essential for an effective chaos engineering platform, and how are they quantified? This question clarifies the definition of chaos engineering, exploring its core principles, key components, and quality requirements. These elements build a foundation for understanding, implementing, and evaluating chaos engineering solutions.
RQ2. What challenges does chaos engineering intend to solve? This question helps identify key challenges, guiding organizations to areas where chaos engineering can best enhance resilience and reliability.
RQ3. What kinds of solutions have been proposed for chaos engineering? RQ3.1. What tools and techniques are utilized, proposed, designed, and implemented by researchers and practitioners? RQ3.2. What best and bad practices are documented in adopting chaos engineering? RQ3.3. How are the chaos engineering approaches evaluated? This question examines the tools and techniques in chaos engineering, highlighting key practices, common pitfalls, and methods for evaluating and enhancing solutions across various contexts.
RQ4. What is the state of chaos engineering in research and industry? RQ4.1. What is the extent and diversity of research conducted on chaos engineering, and what trends can be observed over time? RQ4.2. Who are the thought leaders and key contributors in the field of chaos engineering, and what intersections of thought do they represent? Exploring this question allows us to assess the growth and direction of chaos engineering, highlighting current trends, influential contributors, and areas that need further exploration, ensuring continued advancement in the field.
2.2. Search Strategy
This section outlines the search strategy, detailing the data sources, queries, and the complete search process.
2.2.1. Sources of data
We conducted automated searches in four major digital libraries—Scopus, IEEE Xplore, ACM Digital Library, and Web of Science—using both peer-reviewed scholarly articles and grey literature as sources. These libraries serve as primary sources for software engineering studies (Chen et al., 2010). We performed multi-keyword searches targeting titles, abstracts, and keywords for papers published between January 2019 and April 2024. Additionally, Google Scholar was used to pinpoint any relevant literature that was not included in the previous searches. We used Google Search for grey literature, similar to other multi-vocal literature reviews (MLRs)  (Garousi and Mäntylä, 2016; Garousi et al., 2017; Islam et al., 2019).
2.2.2. Search Terms
Following the SLR guidelines (Kitchenham et al., 2007), we constructed search terms based on the research questions. The first two authors tested the search strings to assess their applicability for addressing the defined research questions. Multiple pilot searches were conducted to refine the search terms, ensuring the inclusion of all relevant papers—boolean operators (AND, OR) combined keywords, synonyms, and related concepts. Table 4 shows the final search strings used across the four digital libraries. We adapted and applied the Google Search search string “(Chaos AND Engineering)” for grey literature to capture relevant non-peer-reviewed sources.
2.3. Eligibility Criteria
Following some existing MLR studies (Islam et al., 2019; Butijn et al., 2020; Garousi et al., 2019), we applied a specific set of inclusion and exclusion criteria, as detailed in Table  2
Table 2. Inclusion and Exclusion Criteria.
Inclusion Criteria Exclusion Criteria
IC1: Articles must be in English and be readily accessible in full text. EC1: Short articles or brief papers (less than six pages).
IC2: Articles must include robust validation methods, including working prototypes or tools with proper references, especially for grey sources. EC2: Articles that lack the relevant focus on chaos engineering.
IC3: Articles that report practices and challenges in implementing chaos engineering, including case studies, best practices, or guidelines. EC3: Articles that predominantly discuss physical infrastructure or purely hardware aspects without direct connection to chaos engineering.
IC4: Articles that provide insights into the trends and evolution of chaos engineering practice. EC4: Studies enhancing specific algorithms or features of a single tool without a broader chaos engineering context unless they provide insights into the field’s evolution or state of tools.
IC5: Articles that identify key contributors and diverse perspectives within chaos engineering. EC5: Duplicate articles from all sources.
IC6: Articles must cover all identified phases of the chaos engineering pipeline, ensuring a holistic view of the field. EC6: Articles lacking empirical evidence or practical applications to support theoretical propositions.
2.4. Study Selection
The libraries’ advanced query modes were used to maximize search coverage. The search was conducted in April 2024, and the results were saved in BibTeX and CSV formats. Duplicates were removed before proceeding with further analysis. Figure  3 illustrates the selection process for academic and grey literature, showing the databases searched and the number of selected papers at each stage. Different approaches were followed to select grey and academic literature.
2.4.1. Academic Literature Selection
Searches in Scopus, IEEE Xplore, ACM DL, and Web of Science yielded 1,290 papers. After duplicate removal, 222 unique papers remained. Applying the inclusion and exclusion criteria, the selection was narrowed to 42 relevant papers. An additional search was conducted using Google Scholar, which added two more papers to the final count, bringing the total to 44 papers.
Figure 3. MLR Study Selection Process.
[…]
MULTIVOCAL_REVIEW_Final
Table 3. Criteria for Qualitative Assessment of Grey Literature.
Criteria Questions
Authority (Q1). Is the organization publishing the work credible?
(Q2). Are there other publications by the author in this domain of study?
(Q3). How closely is the author connected to a reputable institution?
Methodology (Q4). Is the objective of the source evident and well-stated?
(Q5). Can the author be considered an expert in this field?
Date (Q6). Is the publication date of the source within the past five years?
Table 4. Search Strings and Results.
Search Strings ACM DL IEEE Xplore Scopus Web of Science Total
"chaos engineering" OR "chaos test" OR "chaos experiment" OR "chaos toolkit" OR "chaos platform" OR "chaos mesh" OR "chaos integration" OR "chaos architecture" OR "chaos infrastructure" OR "chaos injection" OR "chaos tool" 95 63 121 47 326
"chaos engineering" AND ("trends" OR "developments" OR "progression") 29 0 17 0 46
"chaos engineering" AND ("leader" OR "pioneer" OR "expert" OR "contributor") 35 3 4 0 42
"chaos engineering" AND ("phases" OR "stages" OR "steps" OR"pipeline") 62 5 11 0 78
"chaos engineering" AND ("challenges" OR "issues" OR "problems") 79 16 34 4 133
"chaos engineering" AND ("applications" OR "benefits") 77 20 37 8 142
"chaos engineering" AND ("solutions" OR "approaches" OR "methods") 72 20 55 4 151
"chaos engineering" AND ("tools" OR "techniques") 79 24 36 8 147
"chaos engineering" AND ("best practices" OR "bad practices" OR "implementation" OR "Culture") 68 8 16 5 97
"chaos engineering" AND ("evaluation" OR "metrics") 75 24 26 3 128
Total 671 183 357 79 1290
2.4.2. Grey Literature Selection
We followed the approach of other multi-vocal and grey literature reviews  (Garousi and Mäntylä, 2016; Garousi et al., 2017; Islam et al., 2019; Butijn et al., 2020; Garousi et al., 2019) and used Google Search to gather grey literature. As described by these reviews, we examined the first ten pages of results, as this was deemed sufficient to identify the most pertinent literature since Google’s algorithm emphasizes the most relevant results on the first pages  (Soldani et al., 2018; Verdecchia et al., 2019; Islam et al., 2019). We continued scanning further only if necessary, stopping when no relevant new articles appeared  (Islam et al., 2019). To avoid bias, we searched in incognito mode, following the recommendations of Rahman et al. (Rahman et al., 2023). We incorporated additional quality assessment criteria beyond the inclusion and exclusion criteria, as outlined in Table  3. We picked six key criteria: the credibility of the publishing organization (Q1), the author’s prior publications in the field (Q2), the author’s institutional affiliation (Q3), clarity of purpose (Q4), the author’s expertise (Q5), and recency of the source (Q6). The responses to these criteria were classified as ‘Yes’ (1), ‘Partly’ (0.5), and ‘No’ (0), with scores assigned accordingly for each category. These criteria were adapted from prior studies on grey and multi-vocal literature reviews  (Soldani et al., 2018; Butijn et al., 2020; Kumara et al., 2021), ensuring a thorough evaluation of the sources. As a result, we selected 44 documents, contributing to 88 unique articles (Figure  3) for extracting and synthesizing data. Table  5 illustrates the selected sources’ distribution across various venues.
Table 5. Selected Study for Data Synthesis and Qualitative Evaluation.
Academic Literature Grey Literature
Websites & Blogs Whitepaper
(Fogli et al., 2023; Torkura et al., 2020; Zhang et al., 2019; Dedousis et al., 2023; Simonsson et al., 2021; Zhang et al., 2021a; Cotroneo et al., 2022; Camacho et al., 2022; Torkura et al., 2021; Poltronieri et al., 2022; Zhang et al., 2023; Nikolaidis et al., 2021; Frank et al., 2021b, a; Pierce et al., 2021; Soldani and Brogi, 2021; Al-Said Ahmad et al., 2024; Vu et al., 2022; Chen et al., 2024; Nikolaidis et al., 2023; Ma et al., 2023; Meiklejohn et al., 2021; Frank et al., 2023; Sondhi et al., 2021; Kassab et al., 2022; Jernberg et al., 2020; Konstantinou et al., 2021; Pulcinelli et al., 2023; Ji et al., 2023; Ikeuchi et al., 2020; Chen et al., 2022; Naqvi et al., 2022; Malik et al., 2023; Bedoya et al., 2023; Kesim et al., 2020; Basiri et al., 2019; Park et al., 2019; Klinaku et al., 2022; Rivera et al., 2023; Camilli et al., 2022; Yu et al., 2021; Siwach et al., 2022) = 44 (Inc., 2023; Wickramasinghe, 2023; Gill, 2021; Bremmers, 2021; Patel, 2022; Jakkaraju, 2020; Hawkins, 2020; Nombela, 2023; Le, 2022; National Australia Bank, 2020; Shah, 2021; Krishnamurthy, 2021; Moon, 2022; Katirtzis, 2022; Wachtel, 2023; Durai, 2022; Anwar and Arunachalam, 2019; Tomka, 2024; Dua, 2024; Bocetta, 2019; Mooney, 2023; Wu, 2021; Lardo, 2019; Miles, 2019a; Bhaskar, 2022; Azure Readiness, 2023; Gill, 2022; Fawcett, 2020; Manna, 2021; Gianchandani et al., 2022; Bairyev, 2023; Sbiai et al., 2023; Abdul, 2024; Roa, 2022; Kadikar, 2023; Green, 2023; Aharon, 2024; Krivas and Portela, 2020; Kostic, 2024; Vinisky, 2024; Colyer, 2019) = 41 (Palani and Gupta, 2023; Tivey et al., 2019; Starr et al., 2022) = 3
2.5. Data Extraction, Synthesis, and Analysis
In alignment with the MLR process, we read and assessed the selected papers, extracting and summarizing data to address our research questions.
2.5.1. Data Extraction
To systematically collect essential details, such as author names, paper titles, publication venues, and years, we used a structured data extraction form based on guidelines from Kitchenham et al. (Kitchenham, 2004) and Garousi and Felderer (Garousi and Felderer, 2017). This information was organized in a spreadsheet.
2.5.2. Synthesis and Analysis
We applied thematic analysis to synthesize insights across our research questions, ensuring that coding remained objective and independent of the data extraction structure or biases (Stewart, 2024). Full papers were subsequently imported into ATLAS.ti, a qualitative data analysis tool that allowed for direct text coding. This process adhered to established thematic analysis steps (Braun and Clarke, 2006), providing consistency and depth in our approach. We used a step-by-step approach to systematically identify codes, clusters, and classifications related to our research questions (Kumara et al., 2021):
  • Familiarization with the Data: We began by reading the selected papers to understand the data comprehensively, enabling us to identify the key elements relevant to our research questions.
  • Pilot Study: We randomly selected twenty sources for initial analysis, with the first two authors of the paper independently analyzing each to establish preliminary codes using structural and descriptive coding approaches (Saldaña, 2021). This phase involved creating codes by conceptualizing information related to the research questions. Through iterative review and refinement, researchers adjusted codes to clarify emerging theories and form thematic clusters.
  • Inter-Rater Consensus Analysis for the Pilot Study: After finalizing the coding process, the first two authors of the paper conducted an inter-rater review, assessing each other’s codes to establish consensus on naming, types, and categorization. This collaborative effort helped to refine codes and maintain their consistency.
  • Complete Dataset Coding: The remaining sources were subsequently coded based on the agreement reached during the pilot study. The first author of this paper conducted this step, and the second author reviewed the resulting codes.
  • Inter-Rater Consensus Analysis for the Complete Study: After completing the individual coding, we performed a final inter-rater assessment to ensure the coding aligned with the established framework. Any discrepancies were discussed and resolved.
  • Generating the report: The results of this thematic analysis are presented in Sections 3,  4,  5,  6 and  7.
2.6. Replication Package
For the validation and reproducibility of this work, we have provided an extensive replication package online [^1] https://drive.google.com/drive/folders/1wOu11lB9wsxK7cQKzPL63GAjD9UiiF0s?usp=sharing. This package includes the complete compilation of sources, search queries, and thematic analysis performed using Atlas.ti [^2] https://atlasti.com/ (such as codes, groupings, quotations, and analyses).
3. Chaos Engineering: Definitions, Functionaries, and Elements (RQ1)
This section first examines practitioners’ definitions of chaos engineering. Next, it describes the phases, functional requirements, and non-functional requirements of chaos engineering. Finally, it identifies the main functional components of chaos engineering.
3.1. Definitions
Chaos engineering is a growing practice, but its definition remains inconsistent. Hence, we aim to establish a clear and shared definition by analyzing those proposed by different sources. Originally developed at Netflix, chaos engineering focuses on intentionally introducing faults within a system to assess and enhance its robustness  (Fogli et al., 2023; Torkura et al., 2020; Dedousis et al., 2023; Torkura et al., 2021). However, confusion often arises when distinguishing chaos engineering from traditional testing practices. While conventional testing remains vital, chaos engineering should be seen as a complement rather than a replacement (Green, 2023). Traditional testing and monitoring are crucial, yet they may not be sufficient to expose all possible failure modes in complex systems. Chaos engineering mitigates this issue by proactively identifying the weaknesses of a system that might go unnoticed during standard testing procedures but may surface as failures in production environments (Paul, 2019; Green, 2023).
Table 6. Selected Definitions of Chaos Engineering
Source Definition
Netflix
"the discipline of experimenting on a system to build confidence in its ability to withstand turbulent
conditions in production" (Team, 2024).
Gremlin
"a disciplined approach to identifying failures before they become outages by intentionally
’breaking things’ to build more resilient systems" (Inc., 2023).
AWS
"the process of stressing an application in testing or production environments by creating
disruptive events, such as server outages or API throttling, observing how the system responds,
and implementing improvements" (Starr et al., 2022; Kahil and Arenhage, 2021).
Google "the practice of intentionally introducing failures into production systems under load" (Google Cloud, 2024).
Microsoft "intentionally injecting faults into applications in a controlled manner to simulate real-world issues" (Microsoft, 2023).
IBM
"intentionally causing failures in production or pre-production environments
to understand their impact and improve defense strategies" (IBM, 2024).
Datadog
"a discipline that encourages experimenting in production and injecting controlled failures
into the system to understand how the system will react and improve reliability" (Pulido et al., 2021).
Harness "the science of injecting failure to gauge system resiliency" (Schillerstrom, 2022).
Twilio "intentionally injecting failures to test a system’s ability to self-heal" (Wong, 2017).
Uber "the practice of injecting controlled failures directly into production environments to test resilience" (Jones, 2020).
Splunk
"the practice of assessing the resilience of production systems by intentionally introducing
failures and running experiments" (Wickramasinghe, 2023).
Figure 4. Word Cloud of Key Terms in Chaos Engineering Definitions
Table 6 shows the excerpts from the definitions provided by a subset of the reviewed literature. We also created a word cloud to identify the key terms used in these definitions, as shown in Figure  4. The key themes emphasized in these definitions include system resilience, failure injection, testing in production environments, and controlled experiments. These recurring elements highlight chaos engineering’s focus on deliberately introducing failures in a controlled manner to assess how systems behave under stress and recover from disruptions. Based on these insights, we propose the following unified definition.
Chaos engineering is a resilience testing practice that intentionally injects controlled faults into software systems in production-like or actual production environments to simulate adverse real-world conditions. It enables uncovering hidden vulnerabilities in systems that hinder their ability to withstand potential real-world disruptions. It also enables the assessment and enhancement of systems’ capacity and operational readiness against such disruptions.
In line with the definition, chaos engineering relies on several guiding principles to design and carry out experiments, ensuring that systems are rigorously tested under controlled conditions (Basiri et al., 2016; Aisosa, 2023). Table 7 summarizes these principles.
Table 7. Principles and Measurement Metrics
Principle Description Articles Referenced
Establish Steady State Understand the system’s normal behavior and hypothesize how it will handle chaos (real-world disruptive incidents/failures). (Fogli et al., 2023; Torkura et al., 2020; Zhang et al., 2019; Dedousis et al., 2023; Simonsson et al., 2021; Inc., 2023; Bremmers, 2021; Gill, 2021; Wickramasinghe, 2023; Cotroneo et al., 2022; Nikolaidis et al., 2021; Frank et al., 2021b; Pierce et al., 2021; Frank et al., 2023; Konstantinou et al., 2021; Naqvi et al., 2022; Malik et al., 2023; Bedoya et al., 2023; Kesim et al., 2020; Basiri et al., 2019; Zhang et al., 2023; Siwach et al., 2022; Jernberg et al., 2020; Jakkaraju, 2020; Bocetta, 2019; Palani and Gupta, 2023; Gill, 2022; Fawcett, 2020; Manna, 2021; Bairyev, 2023; Sbiai et al., 2023; Green, 2023; Krivas and Portela, 2020; Kostic, 2024; Abdul, 2024; Katirtzis, 2022; Le, 2022; Colyer, 2019; Patel, 2022)
Setup Monitoring Infrastructure Set up tools to monitor the system for deviations from its normal behavior during controlled experiments. (Zhang et al., 2019; Simonsson et al., 2021; Inc., 2023; Bremmers, 2021; Patel, 2022; Gill, 2021; Wickramasinghe, 2023; Poltronieri et al., 2022; Zhang et al., 2021a; Nikolaidis et al., 2021; Pierce et al., 2021; Nikolaidis et al., 2023; Malik et al., 2023; Kesim et al., 2020; Siwach et al., 2022; Sbiai et al., 2023; Abdul, 2024; Bairyev, 2023; Green, 2023)
Conduct Chaos Experiments Design and execute experiments to simulate real-world failures. (Fogli et al., 2023; Torkura et al., 2020; Zhang et al., 2019; Dedousis et al., 2023; Simonsson et al., 2021; Inc., 2023; Bremmers, 2021; Gill, 2021; Pierce et al., 2021; Naqvi et al., 2022; Malik et al., 2023; Bedoya et al., 2023; Kesim et al., 2020; Jernberg et al., 2020; Jakkaraju, 2020; Palani and Gupta, 2023; Gill, 2022; Fawcett, 2020; Manna, 2021; Bairyev, 2023; Patel, 2022; Abdul, 2024; Roa, 2022)
Test and Refine Review results, apply improvements, and repeat to boost resilience. (Zhang et al., 2019; Simonsson et al., 2021; Inc., 2023; Bremmers, 2021; Gill, 2021; Wickramasinghe, 2023; Jernberg et al., 2020; Bairyev, 2023; Sbiai et al., 2023; Green, 2023; Kostic, 2024; Colyer, 2019; Roa, 2022; Basiri et al., 2019; Torkura et al., 2020; Fogli et al., 2023; Patel, 2022; Wu, 2021; Tomka, 2024)
Grow Blast Radius Gradually expand experiments to test increasingly complex systems and environments. (Torkura et al., 2020; Zhang et al., 2019; Simonsson et al., 2021; Inc., 2023; Gill, 2021; Wickramasinghe, 2023; Pierce et al., 2021; Frank et al., 2023; Poltronieri et al., 2022; Ikeuchi et al., 2020; Nikolaidis et al., 2021; Naqvi et al., 2022; Malik et al., 2023; Kesim et al., 2020; Zhang et al., 2023; Siwach et al., 2022; Jernberg et al., 2020; Jakkaraju, 2020; Fawcett, 2020; Manna, 2021; Bairyev, 2023; Sbiai et al., 2023; Tomka, 2024; Green, 2023; Kostic, 2024)
Expand to Production Run experiments in production to build confidence in the system’s ability to endure adverse real-world events. (Nombela, 2023; Krishnamurthy, 2021; Moon, 2022; Simonsson et al., 2021; Pierce et al., 2021; Konstantinou et al., 2021; Park et al., 2019; Fawcett, 2020)
3.2. Functionalities of Chaos Engineering
Based on our review of academic and grey literature, we identified five key features that define an effective chaos engineering solution: experimentation hub, collaborative workflow management, observability suite, security and governance, and integration and automation (see Figure  5). The rest of this section discusses each in detail.
3.2.1. Experimentation Hub
Experimentation is a crucial aspect of a chaos engineering platform, allowing users to meticulously plan and seamlessly conduct chaos experiments (Katirtzis, 2022; Abdul, 2024; Krivas and Portela, 2020). In this phase, users first define the scope, targets, and parameters of chaos experiments with precision, ensuring a clear focus for testing (Kostic, 2024; Bremmers, 2021; Kadikar, 2023; Wu, 2021; Cotroneo et al., 2022; Katirtzis, 2022; Torkura et al., 2020). The next step is to establish the steady state of the system - baseline performance metrics and normal behavior - which serves as a reference point for evaluating the impact of disruptions (Frank et al., 2023; Konstantinou et al., 2021; Jernberg et al., 2020; Siwach et al., 2022). Based on this steady state, users form hypotheses about how the system will behave under stress, guiding both the design of experiments and the setting of resilience expectations  (Kostic, 2024; Zhang et al., 2019; Wickramasinghe, 2023; Fogli et al., 2023; Dedousis et al., 2023). Once the experiment is defined, the chaos engineering platform injects faults, such as network latency, packet loss, or server failures. This enables users to assess the weaknesses and operational readiness of a system by evaluating its fault handling capabilities (Kostic, 2024; Chen et al., 2024; Fogli et al., 2023; Zhang et al., 2023). As these faults are injected and system behavior is observed, careful planning and execution help chaos experiments effectively identify and address system vulnerabilities, ultimately strengthening overall resilience (Nikolaidis et al., 2021; Colyer, 2019).
3.2.2. Collaborative Workflow Management
Chaos engineering is a collaborative practice that brings together stakeholders such as developers, testers, operations teams, and security experts. These individuals must work closely to design, execute, and analyze experiments to identify system vulnerabilities. Through effective collaborative workflow management, teams can maintain clear communication, coordinate tasks efficiently, and ensure that all parties are informed throughout the experimentation process (Siwach et al., 2022; Dua, 2024; Green, 2023).
Facilitating Team Communication and Coordination.
Chaos engineering relies on continuous communication between relevant parties to ensure alignment with the objectives of the experiment and the expected results (Gill, 2021). By forming cross-functional working groups, participants can discuss scenarios, hypothesize outcomes, and maintain a shared understanding of the objectives of the experiment. This approach ensures that everyone stays coordinated and clear on their roles during the execution and analysis phases. For example, in a simulated database failure, developers can focus on diagnosing and resolving the issue while the operations group works to reroute traffic and maintain service availability. At the same time, security experts assess the potential vulnerabilities that the failure exposes, and testers verify the system response to failures (Patel, 2022). This collaborative process helps avoid information silos and supports collective learning, leading to better preparation for future incidents (Green, 2023; Kadikar, 2023).
Notify Stakeholders about Experiment Progress.
Keeping stakeholders informed about the status of the chaos engineering experiment is crucial to transparency and effective decision making. First, well-designed notification systems trigger alerts based on anomalies, such as unexpected error rates or system performance changes  (Bremmers, 2021). Then, these alerts are automatically sent to stakeholders through integrated notification features. These alerts enable quick issue communication and prompt responses to performance degradation or system instability. In addition, they provide real-time insight into system behaviors, facilitating adjustments during the experiment. Finally, this continuous feedback loop refines the monitoring systems and improves incident preparedness through improved alert accuracy (Shah, 2021).
Maintain Documentation and Knowledge Repositories.
Building a consistent knowledge base for future reference is essential for the long-term success and continuous improvement of chaos engineering practices (Vinisky, 2024; Kadikar, 2023). By maintaining detailed records of experiment scenarios, findings, and recommendations, organizations create a valuable foundation for troubleshooting and informed decision-making (Kadikar, 2023; Gill, 2022). This process should include standardized templates for experiments and comprehensive postmortem reports, ensuring that the documentation is consistent and thorough (Siwach et al., 2022). A well-organized knowledge base preserves essential information and fosters a deeper understanding of system behavior under stress  (Gill, 2021). These insights from previous experiments can directly improve system design and operational practices (Vinisky, 2024). Through consistent documentation and structured knowledge sharing, chaos engineering platforms ensure that lessons learned are preserved, actively shaping future experiments, and strengthening overall system performance (Torkura et al., 2020).
3.2.3. Observability Suite
Chaos engineering platforms use an observability suite to monitor and analyze systems during experiments, allowing teams to detect issues early, identify root causes, and take corrective actions (Green, 2023; National Australia Bank, 2020; Chen et al., 2022). Key features discussed in the following subsections include real-time dashboards and experiment impact analysis to identify faults’ root causes.
Provide Real-Time Dashboards.
Real-time observability depends on dashboards, metrics, traces, and logs for the visualization of experiments. (National Australia Bank, 2020; Gatev, 2021; Pulcinelli et al., 2023). First, dashboards, integrated with monitoring tools, can show critical metrics such as latency, throughput, error rates, and resource usage, allowing engineers to monitor system performance and quickly detect problems (Wickramasinghe, 2023; Manna, 2021; Fawcett, 2020; Green, 2023). Compared to performance metrics, traces can capture the flow of requests through the system, revealing interactions among components and identifying bottlenecks (Chen et al., 2024; Zhang et al., 2021a). Distributed tracing tools allow engineers to analyze component interactions and inefficiencies during chaos experiments (Konstantinou et al., 2021; Frank et al., 2021b). Finally, logs record system events, supporting diagnostic efforts, especially in security contexts (Klinaku et al., 2022; Chen et al., 2022; Green, 2023). Log analysis tools enable engineers to trace anomalies and maintain a complete audit trail  (Torkura et al., 2021; Soldani and Brogi, 2021; Vu et al., 2022).
Analyze Impact and Identify Root Causes of Faults.
After gathering and visualizing data through real-time dashboards, the next crucial step is to analyze the results of chaos experiments and identify the root causes of any faults, using each pillar of observability in the process  (Yu et al., 2021; Gianchandani et al., 2022) The metrics collected during chaos experiments are compared with baseline performance to detect deviations that may indicate vulnerabilities. These deviations help identify how well the system handled stress and where failures occurred (Ji et al., 2023). Logs and trace data provide deeper insight into how individual components interacted and responded to stress. This detailed data helps identify the source of failures and component interactions that contribute to performance problems (Green, 2023). These metrics, logs, and traces can be combined to facilitate root cause analysis and offer qualitative and quantitative insights. For example, this helps engineers determine whether failures were caused by bottlenecks, resource mismanagement, or faulty component interactions (Yu et al., 2021). By understanding the root causes of failures, teams can implement targeted improvements to enhance and prevent future failures (Gill, 2022; Nombela, 2023; Gianchandani et al., 2022).
3.2.4. Security and Governance
Chaos engineering platforms require a robust security and governance framework to ensure that experiments are conducted safely and comply with regulatory standards (Fogli et al., 2023; Kostic, 2024). This section outlines vital mechanisms that protect system integrity while adhering to industry regulations, focusing on implementing safety controls, automated rollbacks, and ensuring regulatory compliance.
Implement Safety Controls and Automated Rollbacks.
Our literature identifies several vital mechanisms, including abort conditions, automated rollbacks, and blast radius control, which ensure that experiments are conducted safely (Patel, 2022).
  • Abort conditions: These ensure that chaos experiments are immediately stopped if predefined thresholds are exceeded (Abdul, 2024; Sbiai et al., 2023). For example, the experiment is automatically terminated if performance metrics, such as latency or availability, degrade beyond acceptable levels. This prevents further damage and maintains system stability by stopping experiments before they can cause widespread problems (Torkura et al., 2020).
  • Automated rollbacks: These allow systems to revert to a known stable state after an experiment (Fawcett, 2020; Miles, 2019b). In cases where a simulated failure causes service disruption, automated rollback mechanisms quickly restore the system to its pre-experiment configuration. This reduces downtime and prevents prolonged instability in the system (Gill, 2022).
  • Blast radius control: This mechanism limits the scope of chaos experiments to isolated parts of the system (Wickramasinghe, 2023; Manna, 2021). For example, teams might start by targeting a single microservice instead of testing the entire infrastructure. Gradually expanding the blast radius allows controlled testing, ensuring vulnerabilities can be identified without risking widespread system failures (Abdul, 2024). Additionally, Gremlin  (Inc., 2023) recommends using the Rumsfeld matrix, a framework to categorize system behaviors during chaos experiments. This approach enables teams to understand better and anticipate various types of system responses by classifying them as follows:
    • Known Knowns: Aspects of the system that are fully understood and predictable.
    • Known Unknowns: Recognized uncertainties that pose potential risks.
    • Unknown Knowns: Understood risks that may not be immediately apparent or considered.
    • Unknown Unknowns: Completely unforeseen issues that may emerge during experimentation.
Starting with known knowns, teams establish a baseline of predictable system behavior. They can then progressively move into more uncertain areas like Known Unknowns and Unknowns Knowns before finally addressing the most unpredictable Unknown Unknowns (Palani and Gupta, 2023). This step-by-step approach reduces system risks by building understanding before testing less predictable areas.
Ensure Regulatory Compliance.
Adhering to industry standards and legal frameworks is essential to ensure that systems meet established data protection and security requirements, including guidelines such as the General Data Protection Regulation (GDPR) and IEC 62443 (Helmke, 2019; Torkura et al., 2021). These frameworks provide organizations with critical guidance to safeguard sensitive information, maintain operational integrity, and protect vital infrastructures (Helmke, 2019; Kostic, 2024). To meet these compliance requirements, some organizations adopt Infrastructure as Code (IaC), which automates the provisioning and management of infrastructure (Torkura et al., 2020; Fogli et al., 2023). Tools such as Terraform and Ansible are commonly used in IaC to ensure consistent implementation of security policies, minimize human error, and maintain infrastructure in alignment with regulatory standards (Torkura et al., 2021; Vu et al., 2022). In addition to IaC, monitoring tools such as Datadog  (Abdul, 2024; Mooney, 2023), which detects compliance violations and configuration drift such as publicly exposed data, and Security Monkey, which monitors Secure Sockets Layer (SSL) certificates and other security settings to prevent vulnerabilities such as certificate expiration, are essential to maintain compliance (Dua, 2024; Tivey et al., 2019).
3.2.5. Integration and Automation with Continuous Integration and Continuous Delivery (CI/CD)
CI/CD pipelines are critical to optimizing the efficiency and consistency of chaos experiments (Rivera et al., 2023; Bairyev, 2023; Azure Readiness, 2023). When performed manually, tasks such as performing experiments and monitoring system behavior can be inefficient and prone to human error (Kostic, 2024; Jernberg et al., 2020; Bremmers, 2021; Chen et al., 2024). Automating these processes minimizes human intervention, ensuring tasks are performed reliably and consistently, thus reducing the likelihood of mistakes (Wickramasinghe, 2023; Camacho et al., 2022; Simonsson et al., 2021; Green, 2023). Integrating chaos experiments into CI/CD pipelines further enhances this reliability by automatically initiating tests as part of the routine development workflow (Roa, 2022). With each new deployment or configuration update, chaos tests run alongside other routine tests, enabling early detection of vulnerabilities. This proactive approach helps ensure issues are addressed in the development phase, resulting in a more comprehensive and practical testing process (Bremmers, 2021; Manna, 2021)
Figure 5. Key Functionalities Provided by Chaos Engineering.
[…]
Functional Requirement
3.3. Quality Requirements for Chaos Engineering Platform
We have identified 12 quality requirements for a chaos engineering platform from the selected literature: performance, recoverability, resilience, timeliness, observability, scalability, effectiveness, robustness, accuracy, fault tolerance, cost, and adaptability. We provide these requirements and the corresponding metrics for measuring them in Table 8.
Table 8. Quality Attributes and Measurement Metrics
Quality Attribute Measurement Metrics Articles Referenced Total
Performance Evaluate resource use and performance under attack, ensuring chaos experiments align with production metrics. (Fogli et al., 2023; Torkura et al., 2020; Zhang et al., 2019; Dedousis et al., 2023; Patel, 2022; Gill, 2021; Cotroneo et al., 2022; Nikolaidis et al., 2021; Vu et al., 2022; Klinaku et al., 2022) 10
Recoverability Measure mean time to repair (MTTR), restoration speed, and detection response. (Fogli et al., 2023; Torkura et al., 2020; Dedousis et al., 2023; Patel, 2022; Gill, 2021; Cotroneo et al., 2022; Camacho et al., 2022; Zhang et al., 2021a; Nikolaidis et al., 2021; Pierce et al., 2021; Vu et al., 2022; Frank et al., 2021b; Malik et al., 2023; Zhang et al., 2023; Tomka, 2024; Kadikar, 2023) 16
Resilience Tolerance for degraded performance under stress; success rate of automated restoration actions. (Fogli et al., 2023; Torkura et al., 2020; Zhang et al., 2019; Dedousis et al., 2023; Inc., 2023; Bremmers, 2021; Zhang et al., 2021a; Al-Said Ahmad et al., 2024; Vu et al., 2022; Nikolaidis et al., 2023; Frank et al., 2023; Kassab et al., 2022; Konstantinou et al., 2021; Nombela, 2023; Krishnamurthy, 2021; Bocetta, 2019; Manna, 2021; Kostic, 2024; Krivas and Portela, 2020; National Australia Bank, 2020; Azure Readiness, 2023) 20
Timeliness Time to restore service and detect interruptions; impact on user experience and disruption. (Camacho et al., 2022; Pierce et al., 2021; Vu et al., 2022; Frank et al., 2021b; Malik et al., 2023; Cotroneo et al., 2022; Basiri et al., 2019; Gill, 2021; Zhang et al., 2019; Durai, 2022; Shah, 2021; Katirtzis, 2022; Dua, 2024; Wu, 2021; Palani and Gupta, 2023; Manna, 2021) 16
Observability Assess logging completeness and monitoring uptime; measure alert latency and tracing success rate. (Fogli et al., 2023; Torkura et al., 2020; Dedousis et al., 2023; Inc., 2023; Bremmers, 2021; Patel, 2022; Gill, 2021; Camacho et al., 2022; Torkura et al., 2021; Zhang et al., 2021a; Vu et al., 2022; Frank et al., 2023; Malik et al., 2023; Kesim et al., 2020; Siwach et al., 2022; Konstantinou et al., 2021; Hawkins, 2020; Manna, 2021; Kostic, 2024; Kadikar, 2023; Azure Readiness, 2023) 20
Scalability Manage increasing attack rates and scale experiment complexity. (Fogli et al., 2023; Torkura et al., 2020; Zhang et al., 2019; Patel, 2022; Gill, 2021; Vu et al., 2022; Sondhi et al., 2021; Malik et al., 2023; Klinaku et al., 2022; Camilli et al., 2022; Nikolaidis et al., 2023; Kassab et al., 2022; Rivera et al., 2023; Krishnamurthy, 2021; Durai, 2022; Wu, 2021; Palani and Gupta, 2023; Manna, 2021; Azure Readiness, 2023; Vinisky, 2024; Kostic, 2024) 20
Effectiveness Measure achievement rates, end-user satisfaction, and time to achieve outcomes. (Sondhi et al., 2021; Chen et al., 2022; Malik et al., 2023; Wu, 2021) 4
Robustness Evaluate stability at peak load and error tolerance. (Vu et al., 2022; Ma et al., 2023; Yu et al., 2021; Wu, 2021; Bedoya et al., 2023; Chen et al., 2022; Vinisky, 2024) 7
Accuracy Evaluate system consistency through recovery accuracy and fault localization precision. (Green, 2023; Simonsson et al., 2021; Chen et al., 2022) 3
Fault Tolerance Automated failover success rate during faults; immediate recovery time after detected faults. (Fogli et al., 2023; Torkura et al., 2020; Zhang et al., 2019; Dedousis et al., 2023; Gill, 2021; Wickramasinghe, 2023; Zhang et al., 2021a; Camacho et al., 2022; Poltronieri et al., 2022; Nikolaidis et al., 2021; Ma et al., 2023; Meiklejohn et al., 2021; Sondhi et al., 2021; Azure Readiness, 2023; Tomka, 2024; Manna, 2021) 15
Cost Cost of chaos engineering and savings from preventing outages. (Gill, 2021; Patel, 2022; Kesim et al., 2020; Palani and Gupta, 2023; Klinaku et al., 2022; Inc., 2023; Bocetta, 2019; Basiri et al., 2019; National Australia Bank, 2020) 9
Adaptability Change methods and platforms for new systems. (Vu et al., 2022; Ma et al., 2023; Yu et al., 2021; Wu, 2021) 4
3.4. Key Components of Chaos Engineering
Before adopting a chaos engineering framework, organizations should carefully evaluate its key components to ensure that they align with their operational requirements and objectives. We have identified several essential components for an effective chaos engineering platform. Most of the reviewed studies incorporate a combination of these components, which we have categorized into five central units: experimentation, collaboration and communication, observability and analysis, security and governance, and automation and integration with CI/CD. This classification, derived from the functionalities discussed in Section  3.2, is shown in Figure  6. The remainder of this section provides a detailed analysis of each component within this framework.
Figure 6. Key Components of Chaos Engineering.
[…]
Key Components of Chaos Engineering.
3.5. Experimentation Unit
Successful coordination of chaos experiments depends on a systematic approach to both planning and execution (Zhang et al., 2021a; Nikolaidis et al., 2021). To this end, the experimentation unit is built around two foundational components: the experiment designer and the failure injection hub, which will be explored further in this section.
3.5.1. Experiment Designer
This module is crucial for structuring chaos experiments, allowing teams to define which systems to target (Zhang et al., 2021a; Camacho et al., 2022). After identifying the systems, the teams can choose the fault types most relevant to their testing scenarios, such as network failures or system crashes (Zhang et al., 2021a). Once the faults are defined, the experiment designer helps organize the timing and order in which they are introduced, ensuring that they reflect real-world conditions (Camacho et al., 2022). By enabling these customized failure simulations, the experiment designer ensures that the tests are well-structured, providing consistent data for postmortem analysis (Zhang et al., 2021a; Konstantinou et al., 2021). This structure allows teams to track how the behavior of the system evolves under different failure conditions. For example, teams might simulate a server crash during peak traffic to test failover capabilities and measure the impact on user experience (Zhang et al., 2019; Simonsson et al., 2021).
3.5.2. Fault Injection Hub
Once the experiment is designed and the faults are organized, the next phase is to inject these controlled faults and workloads into the system, simulating real-world scenarios (Fawcett, 2020; Palani and Gupta, 2023). The fault injection process, as outlined by Hseh et al (Hsueh et al., 1997), in Figure  7, involves multiple components working together: a controller, fault injector, fault library, workload generator, workload library, and target system (Nikolaidis et al., 2021; Frank et al., 2021b). Each of these components plays a critical role in ensuring the accuracy and smooth execution of the chaos experiment. Central to this process is the controller, which manages the introduction of faults and dynamically adjusts based on real-time system feedback, ensuring that the experiment remains aligned with its objectives (Torkura et al., 2020). As part of the hub, the fault injector introduces specific disruptions, such as hardware or software failures, selected from a diverse fault library that offers teams a variety of failure scenarios to choose from (Soldani and Brogi, 2021; Frank et al., 2021b). In parallel, the workload generator creates synthetic workloads based on predefined patterns stored in the workload library. This allows teams to simulate normal or peak traffic conditions to observe how the system handles varying stress levels (Fogli et al., 2023; Kesim et al., 2020). These faults and workloads are then applied to the target system, be it a production or preproduction environment, providing teams with valuable information on the weaknesses of the system and the recovery capabilities (Wickramasinghe, 2023; Zhang et al., 2021a).
Figure 7. Conceptual Schema of Fault Injection(Hsueh et al., 1997).
[…]
Conceptual schema of fault injection.
3.6. Collaboration and Communication Unit
Executing chaos experiments across teams requires strong communication and access to data from previous experiments (Kadikar, 2023; Patel, 2022; Siwach et al., 2022), which are supported by the alert and notification system and the knowledge base.
3.6.1. Alert and Notification System
The hub integrates chaos tools with existing alert systems to ensure smooth communication during chaos experiments. This integration ensures that teams are immediately notified of critical events such as experiment initiation, fault injection, and any unexpected behaviors during the process (Starr et al., 2022; Vinisky, 2024). These real-time notifications are distributed through multiple channels, including email and Slack, ensuring that all stakeholders receive timely updates. For example, if an issue arises, the system immediately alerts the relevant individuals, allowing them to respond quickly and address the problem effectively (Torkura et al., 2020; Shah, 2021). This prompt response mechanism is critical for maintaining system stability, as it helps mitigate potential risks before they escalate (Siwach et al., 2022).
3.6.2. Knowledge Base
Maintaining a well-organized archive of previous chaos experiments is crucial to continuously improve the resilience of the system and support real-time communication between users (Kadikar, 2023). This knowledge base is a central repository that stores comprehensive documentation that includes experiment designs, execution procedures, and results (Zhang et al., 2019; Torkura et al., 2020). Recalling past experiments, teams can leverage valuable insights from successful and unsuccessful tests, fostering a cycle of learning and enhancement. Detailed post-mortem reports, for example, allow teams to thoroughly analyze the results, identify improvement areas, and strengthen the resilience of the system (Inc., 2023; Palani and Gupta, 2023). Furthermore, using standardized templates ensures that this knowledge is systematically preserved, promoting consistency and reliability in future chaos experiments (Siwach et al., 2022; Zhang et al., 2019).
3.7. Observability and Analysis Unit
Monitoring system behavior during chaos experiments is crucial for understanding faults’ impact and identifying areas for improvement (Gianchandani et al., 2022; Nikolaidis et al., 2021; Yu et al., 2021). Essential components, such as data collector and data analyzer, play a vital role in this process, ensuring that the systems are evaluated and optimized effectively (Zhang et al., 2021a; Simonsson et al., 2021).
3.7.1. Data Collector
Effective data collection is essential during chaos experiments, capturing various types of information such as performance metrics, logs, and traces (Gatev, 2021). These data points clearly explain how the system responds under stress (Zhang et al., 2021a). The data collector gathers performance metrics to track key indicators such as response times, error rates, latency, and system throughput. These data are crucial to measuring the effectiveness of the system in detecting faults  (Cotroneo et al., 2022). Logs document significant events within the system, including errors and state transitions. These records help diagnose issues and improve the understanding of system behavior during failures  (Vu et al., 2022; Cotroneo et al., 2022). Tracing methods can monitor the flow of requests across distributed services, enabling the identification of bottlenecks and observing how traffic is managed under stress (Zhang et al., 2021a). All data collected are visualized through a centralized dashboard, providing a straightforward and organized interface that allows quick assessment of performance indicators and overall system health (Malik et al., 2023; Basiri et al., 2019; Camacho et al., 2022).
3.7.2. Data Analyzer
The information collected during chaos experiments is systematically analyzed to identify essential trends and patterns (Zhang et al., 2021a; Simonsson et al., 2021). This process involves impact analysis, which evaluates how failures influence system performance based on previously collected metrics (Pulcinelli et al., 2023). By analyzing these data, it is possible to determine how disruptions affect the system and identify its most vulnerable components  (Fawcett, 2020; Cotroneo et al., 2022). Additionally, root cause analysis leverages logs and traces to trace failures back to their origins, whether due to a configuration error, service malfunction, or hardware fault (Zhang et al., 2021a; Cotroneo et al., 2022).
3.8. Security and Governance Unit
Ensuring the secure execution of chaotic experiments in production environments is essential to prevent disruptions and unauthorized actions, achieved through adequate safety and access control mechanisms (Al-Said Ahmad et al., 2024; Green, 2023; Gianchandani et al., 2022). First, safety controls mitigate the risk of chaos experiments that cause severe outages by incorporating predefined thresholds, such as automated rollbacks, to restore systems to a steady state if an experiment causes prolonged disruptions (Zhang et al., 2021a; Gill, 2022). These controls allow the system to automatically abort experiments when severe impacts are detected, maintaining overall system stability (Kadikar, 2023). By limiting the blast radius to specific areas, safety controls ensure that experiments do not cause widespread failures (Simonsson et al., 2021; Gill, 2021), with fail-safe conditions ensuring immediate containment of issues (Bhaskar, 2022). Second, access control mechanisms further enhance security by restricting the ability to initiate, modify, or stop chaos experiments only to authorized personnel (Torkura et al., 2020; Sbiai et al., 2023). This ensures that only individuals with the appropriate permissions can execute or alter experiments, particularly in sensitive environments where unauthorized changes could lead to significant security and compliance risks  (Fogli et al., 2023; Tivey et al., 2019).
3.9. Automation and Integration Unit
This unit is vital in embedding chaos engineering into the software development lifecycle. It automates and integrates chaos experiments with existing systems to ensure consistent and efficient execution. First, the unit connects the chaos engineering platform with essential tools such as CI/CD pipelines, cloud platforms, and monitoring systems. This integration allows chaos experiments to be seamlessly incorporated into routine workflows such as deployments, updates, and system monitoring, enabling automatic execution as part of regular operations (Zhang et al., 2021a; Torkura et al., 2020). Second, the unit automates the execution of chaos experiments and includes a scheduling mechanism that allows tests to run at predefined intervals, such as during off-peak hours or after system updates (Cotroneo et al., 2022; Zhang et al., 2019). This ensures that chaos experiments are performed regularly and reliably, reducing the risk of human error and allowing organizations to test their systems without interrupting normal activities (Camacho et al., 2022; Simonsson et al., 2021).
4. Motivation Behind Chaos Engineering (RQ2)
This section delves into the challenges driving chaos engineering adoption in organizations. Figure 8 shows the technical and socio-technical challenges we identified from the reviewed literature.
Figure 8. Challenges that Drive Chaos Engineering Adoption.
[…]
Challenges that drive chaos engineering adoption.
4.1. Technical Challenges
Chaos engineering aims to tackle various technical challenges that undermine system resilience. The following subsections explore these challenges in detail.
Unanticipated System Failures
In a distributed system, one of the most critical risks is the occurrence of unexpected failures in various parts of the system (Ahmed and Wu, 2013). Given the highly interconnected nature of modern systems, a failure in one component can quickly cascade into widespread disruptions across multiple services or functions (Wickramasinghe, 2023). However, traditional testing methods often struggle to identify deep-rooted system vulnerabilities because they focus on known scenarios and isolated components, neglecting real-world systems’ dynamic and complex interactions (Soldani and Brogi, 2021). In contrast, chaos engineering adopts a proactive approach by deliberately introducing failures, such as server crashes, network interruptions, and resource exhaustion, in environments that closely resemble production. As a result, engineers can observe how the system behaves under stress, thus revealing hidden weaknesses that conventional testing methods might overlook. By focusing on unpredictable interactions between distributed components, chaos engineering effectively uncovers vulnerabilities that arise from the complexity of the system (Bairyev, 2023). For example, chaos experiments might simulate network partitioning between microservices, which can reveal issues related to data consistency and service availability (Bairyev, 2023; Fogli et al., 2023). These controlled disruptions allow engineers to better understand how different system components interact under turbulent conditions (Torkura et al., 2020).
Inadequate Fault Tolerance Mechanisms
Many distributed systems are built with fault tolerance mechanisms to handle routine issues, such as server failures or network delays (Inc., 2023). However, when larger-scale failures occur, e.g., multi-region outages or prolonged network disruptions, these mechanisms often fail to maintain system stability (Palani and Gupta, 2023; Silveira, 2023; Nallainathan, 2023). This shortfall occurs because mechanisms designed to handle more minor issues may not be able to handle the broader impact of significant failures, leading to severe performance degradation or even total system outages (Basiri et al., 2019; Gill, 2021). Chaos engineering addresses these shortcomings by targeting and testing fault tolerance mechanisms in scenarios replicating the most significant risks to system stability (Tivey et al., 2019). For example, chaos engineering can simulate the unavailability of critical services in one region to determine whether traffic is adequately redirected to the other areas or if the load balancers of the system can effectively distribute incoming requests between redundant resources (Bremmers, 2021; Aharon, 2024). Furthermore, by testing the ability of backup systems to maintain service levels during prolonged outages, chaos engineering reveals whether these fallback mechanisms are sufficient for extended failures (Inc., 2023; Wickramasinghe, 2023; Basiri et al., 2019). These focused experiments help ensure systems can handle severe disruptions without experiencing widespread failure (Bairyev, 2023).
Limited Performance Insights and Failure Assessment
Another significant challenge in distributed systems is understanding how they perform during high-stress or failure scenarios (Chen et al., 2006; Fidge, 1996). When systems operate under normal conditions, traffic bottlenecks, inefficient resource usage, or service communication failures often remain hidden (Fogli et al., 2023). This lack of visibility leaves teams unprepared for unexpected disruptions, which can lead to outages or performance degradation when systems are stressed (Konstantinou et al., 2021). Chaos engineering addresses this by introducing specific disruptions, such as intentionally increasing network latency or restricting memory resources on critical components (Siwach et al., 2022). These controlled failures are monitored through observability tools that provide real-time visibility into metrics such as latency, error rates, and resource consumption (Wachtel, 2023). By analyzing these metrics, teams gain detailed insight into how services handle stress - uncovering issues such as delayed responses or inefficient resource utilization that would otherwise remain hidden during normal operations (Bremmers, 2021; Palani and Gupta, 2023).
4.2. Socio-Technical Challenges
Chaos engineering can address various socio-technical challenges, expanding its impact beyond technical issues to encompass the complex interplay between people, processes, and technology (Miles, 2019b; Thomas et al., 2019; Patriarca et al., 2018).
Inconsistent Experiment Design and Execution
When experiment design and execution lack consistency, teams struggle to accurately assess the behavior of the system under stress (Farias et al., 2012; Caferra et al., 2023). Factors such as varying test timing, failure scenarios, or changing system states lead to unreliable data. This makes it difficult to identify critical vulnerabilities or adequately assess the system’s ability to handle disruptions (Klinaku et al., 2022). Without standardization, essential insights into system responses may be overlooked, creating false confidence in its resilience (Danilowicz and Nguyen, 2003). Chaos engineering addresses this by introducing a structured and repeatable approach, ensuring that predefined failure scenarios such as network issues or resource depletion are applied consistently across tests (Zhang et al., 2021a; Naqvi et al., 2022; Kostic, 2024; Vu et al., 2022). This uniformity eliminates variability, producing reliable, comparable data that allows teams to pinpoint weaknesses and confidently evaluate recovery mechanisms (Kadikar, 2023; Azure Readiness, 2023).
Lack of Collaboration and Knowledge Sharing
Limited collaboration and challenges in knowledge sharing between teams, such as development, operations, and security, can lead to communication gaps and the formation of silos, where vital information may not be effectively exchanged (Habeh et al., 2021; Riege, 2005). This results in misaligned responses and repeated errors during system failures, which ultimately slows recovery efforts and misses opportunities for improvement (de Vreede et al., 2016). Chaos engineering tackles this issue by fostering cross-functional collaboration, engaging all relevant teams in both the planning and execution of chaos experiments (Tomka, 2024; Tivey et al., 2019; Azure Readiness, 2023). For example, when simulating a specific failure, such as a service outage or database failure, development teams define the expected behavior of the system, operations teams monitor the system’s performance under stress, and security teams evaluate potential vulnerabilities exposed during the failure (Krishnamurthy, 2021). This multi-team involvement ensures that the insights from the experiment, whether an unexpected failure response or a missed alert, are immediately shared and understood by everyone (Green, 2023; Katirtzis, 2022; Dua, 2024). Furthermore, the results of each chaos experiment are documented and discussed in collaborative postmortem sessions, where all teams contribute to analyzing what went wrong, how to fix it, and how to prevent similar issues in the future (Siwach et al., 2022).
Lack of Risk Management
Effective risk management is critical in both production and non-production environments, yet many systems are exposed to unexpected failures due to insufficient preparation (Bremmers, 2021; Naqvi et al., 2022). In non-production settings, such as development or staging, failure scenarios are often inadequately tested, leading to a false sense of security (Wickramasinghe, 2023; Jernberg et al., 2020). These environments typically lack the complexity of live production systems, meaning that when software is deployed, it may not be fully equipped to handle adverse real-world events (Green, 2023). Consequently, when systems reach production, where they face real user traffic, external dependencies, and unpredictable conditions, even minor issues can escalate to significant outages, resulting in costly downtime and degraded user experiences (Gill, 2021; Zhang et al., 2019; Ji et al., 2023). Chaos engineering solves these challenges by enabling controlled experimentation that tests system behavior under simulated and real-world failure conditions (Torkura et al., 2020; Kostic, 2024). In production environments, controlled blast radiuses ensure that failures are confined to isolated parts of the system, thereby preventing widespread disruption (Palani and Gupta, 2023; Siwach et al., 2022; Gianchandani et al., 2022; Mooney, 2023; Bocetta, 2019). Predefined abort conditions and rollback procedures allow teams to stop experiments and recover quickly if the stability of the system is threatened (Abdul, 2024; Sbiai et al., 2023; Siwach et al., 2022).
Incident Management and Improvement
In complex, distributed systems, incident management faces significant challenges due to the unpredictability of failures and the limitations of reactive strategies (Gianchandani et al., 2022). In these approaches, teams respond to incidents only after they occur, focusing on diagnosing and resolving issues after disruptions have already impacted operations (Bremmers, 2021). This often leads to prolonged downtime and repeated failures, as reactive methods limit the opportunities to address these problems proactively (Patel, 2022). Chaos engineering offers a proactive approach by simulating targeted failures that stress critical parts of the system, revealing weaknesses in the incident response process (Anwar and Arunachalam, 2019; Siwach et al., 2022). For example, simulating a sudden network partition might show whether monitoring systems detect the issue promptly and if the system can reroute traffic as intended (Vinisky, 2024; Green, 2023; Gill, 2021). Similarly, inducing a Central Processing Unit (CPU) or memory exhaustion event could expose delays in triggering alerts or inefficiencies in the automated recovery process (Camacho et al., 2022; Kadikar, 2023; Palani and Gupta, 2023; Roa, 2022). By creating these controlled failure scenarios, chaos engineering helps teams identify where their incident response protocols fall short, whether in monitoring, alert escalation, or failover processes (Wu, 2021; Zhang et al., 2021a). This enables teams to proactively adjust their strategies, improving their ability to handle actual incidents faster, more accurately, and with minimal disruption (Dua, 2024; Rivera et al., 2023).
4.3. Benefits
Table 9 outlines the organizational benefits derived from the functionalities described in Section 3.2, addressing the previously identified challenges. These functions include defining chaos experiments, facilitating team communication, notifying stakeholders, maintaining documentation, and using real-time monitoring dashboards. The corresponding activities, such as establishing experimental parameters, analyzing system performance metrics, and fostering cross-functional collaboration, contribute to several organizational advantages. These include improved system understanding, enhanced error-handling capabilities, vital team coordination during stress conditions, and proactive system tuning. Furthermore, automation of tasks and real-time system monitoring can minimize disruptions and optimize system performance and user experience.
Table 9. Mapping of Key Activities Performed by Chaos Engineering with Organizational Benefits
Key Functions Activity Performed Benefits of Organization Articles Total
Define Chaos Experiments Set experiment scope, define normal conditions, hypothesize, and inject fault. Gain system insights, refine the design, establish a baseline, and test error handling. (Kostic, 2024; Gill, 2021; Manna, 2021; Inc., 2023; Wickramasinghe, 2023; Mooney, 2023; Fawcett, 2020; Vinisky, 2024; Al-Said Ahmad et al., 2024; Pierce et al., 2021; Torkura et al., 2020; Vu et al., 2022; Bedoya et al., 2023; Siwach et al., 2022; Zhang et al., 2019; Le, 2022; Palani and Gupta, 2023; Soldani and Brogi, 2021; Azure Readiness, 2023; Dedousis et al., 2023; Colyer, 2019; Patel, 2022; Shah, 2021; Green, 2023; Fogli et al., 2023; National Australia Bank, 2020; Cotroneo et al., 2022; Konstantinou et al., 2021; Zhang et al., 2023; Bairyev, 2023) 30
Facilitate Team Coordination Create groups for team collaboration. Strengthens coordination and insight sharing. (Gill, 2021; Manna, 2021; Dua, 2024; Kostic, 2024; Katirtzis, 2022; Krishnamurthy, 2021; Torkura et al., 2020; Green, 2023; Frank et al., 2021b; Kesim et al., 2020; Bocetta, 2019; Bremmers, 2021) 12
Notify Stakeholders Update stakeholders on progress and results. Promote transparency and prompt updates. (Gill, 2021; Manna, 2021; Dua, 2024; Kostic, 2024; Katirtzis, 2022; Krishnamurthy, 2021; Torkura et al., 2020; Green, 2023; Frank et al., 2021b; Kesim et al., 2020; Bocetta, 2019; Bremmers, 2021) 12
Maintain Knowledge Repositories Share findings for organizational learning. Create a knowledge base to analyze system stress behavior. (Gill, 2021; Manna, 2021; Dua, 2024; Kostic, 2024; Katirtzis, 2022; Krishnamurthy, 2021; Torkura et al., 2020; Green, 2023; Frank et al., 2021b; Kesim et al., 2020; Bocetta, 2019; Starr et al., 2022; Vinisky, 2024; Anwar and Arunachalam, 2019; Kadikar, 2023; Patel, 2022; Vu et al., 2022; Zhang et al., 2019, 2023; Rivera et al., 2023; Bremmers, 2021) 21
Provide Real-Time Dashboards for Visualization Track faults and visualize performance under stress. Assess system health for rapid anomaly detection and refinement. (Tivey et al., 2019; Kadikar, 2023; Pierce et al., 2021; Zhang et al., 2021a; Wickramasinghe, 2023; Gill, 2022; National Australia Bank, 2020; Simonsson et al., 2021; Al-Said Ahmad et al., 2024; Torkura et al., 2020; Gill, 2021; Kostic, 2024; Fogli et al., 2023; Bedoya et al., 2023; Manna, 2021; Krivas and Portela, 2020; Poltronieri et al., 2022; Camacho et al., 2022; Torkura et al., 2021; Pulcinelli et al., 2023; Chen et al., 2022; Yu et al., 2021; Gianchandani et al., 2022; Palani and Gupta, 2023; Anwar and Arunachalam, 2019; Krishnamurthy, 2021; Roa, 2022; Vinisky, 2024; Siwach et al., 2022; Green, 2023; Hawkins, 2020; Vu et al., 2022; Ma et al., 2023; Rivera et al., 2023) 34
Analyze Impact and Identify Root Causes of Faults Assess outcomes to measure impact and analyze root causes. Enhance resilience, identify weaknesses, and improve stability. (Bairyev, 2023; Siwach et al., 2022; Fogli et al., 2023; Zhang et al., 2019; Inc., 2023; Chen et al., 2024; Bhaskar, 2022; Al-Said Ahmad et al., 2024; Gill, 2022; Wickramasinghe, 2023; Yu et al., 2021; Ji et al., 2023; Gianchandani et al., 2022; Nikolaidis et al., 2021; Rivera et al., 2023; Konstantinou et al., 2021; Torkura et al., 2020) 17
Implement Safety Controls and Automated Rollbacks Set up safety mechanisms, automated rollbacks, and abort thresholds to prevent severe disruptions. Minimizes risk, ensures safe experimentation, and reduces downtime with quick recovery measures. (Wickramasinghe, 2023; Siwach et al., 2022; Bocetta, 2019; Gill, 2022; Fawcett, 2020; Manna, 2021; Gianchandani et al., 2022; Bairyev, 2023; Malik et al., 2023; Inc., 2023; Ikeuchi et al., 2020; Ma et al., 2023; Nikolaidis et al., 2023; Bhaskar, 2022; Simonsson et al., 2021; Gill, 2021; Krishnamurthy, 2021; Torkura et al., 2020; Katirtzis, 2022; Bedoya et al., 2023; Kadikar, 2023; Patel, 2022; Soldani and Brogi, 2021; Frank et al., 2021b; Rivera et al., 2023; Chen et al., 2024; Zhang et al., 2021a, 2019; Chen et al., 2022; Basiri et al., 2019; Bremmers, 2021; Jernberg et al., 2020; Kostic, 2024) 31
Ensure Regulatory Compliance Align chaos experiments with regulatory standards and monitor compliance. Reduce non-compliance risk and build trust by adhering to industry standards. (Wickramasinghe, 2023; Siwach et al., 2022; Bocetta, 2019; Gill, 2022; Fawcett, 2020; Manna, 2021; Gianchandani et al., 2022; Bairyev, 2023; Malik et al., 2023; Inc., 2023; Ikeuchi et al., 2020; Ma et al., 2023; Nikolaidis et al., 2023; Bhaskar, 2022; Simonsson et al., 2021; Gill, 2021; Jernberg et al., 2020; Kadikar, 2023; Kesim et al., 2020; Zhang et al., 2019; Krivas and Portela, 2020; Sbiai et al., 2023; Mooney, 2023; Dua, 2024; Torkura et al., 2021; Palani and Gupta, 2023; Pulcinelli et al., 2023; Kostic, 2024) 28
Integrate with CI/CD and Third-Party Tools Embed chaos experiments in CI/CD with monitoring integration. Automate resilience testing in the software lifecycle to boost efficiency. (Manna, 2021; Azure Readiness, 2023; Konstantinou et al., 2021; Palani and Gupta, 2023; Kostic, 2024; Torkura et al., 2021) 6
Automate Repetitive and Manual Tasks Automate setup and cleanup of chaos experiments using scripts and tools. Reduce time and errors by automating repetitive tasks for consistency and efficiency. (Camacho et al., 2022; Kadikar, 2023; Miles, 2019a; Green, 2023; Nikolaidis et al., 2021; Poltronieri et al., 2022; Fogli et al., 2023; Torkura et al., 2020; Gill, 2021; Cotroneo et al., 2022; Patel, 2022; Bairyev, 2023; Colyer, 2019; Manna, 2021; Zhang et al., 2019; Kostic, 2024; Fawcett, 2020; Jakkaraju, 2020; Zhang et al., 2021a; Al-Said Ahmad et al., 2024; Mooney, 2023; Kesim et al., 2020; Zhang et al., 2023; Jernberg et al., 2020; Bremmers, 2021; Torkura et al., 2021; Vinisky, 2024; Chen et al., 2022; Vu et al., 2022) 29
5. Chaos Engineering Framework: Taxonomy, Tools, Practices, and Evaluation (RQ3)
This section presents a taxonomy for chaos engineering and compares existing tools using the taxonomy dimensions. We also present the best practices for adopting chaos engineering in organizations.
5.1. Taxonomy of Chaos Engineering
Figure 9 shows the proposed taxonomy, which organizes the capabilities of a chaos engineering platform across several key dimensions: execution environment, automation strategy, automation mode, deployment type, and evaluation approach. This taxonomy provides organizations with a structured approach to adopting chaos engineering practices by offering clear guidelines for selecting appropriate execution environments, choosing automation strategies that integrate essential tools such as monitoring, logging, and observability, and determining the optimal deployment types (e.g., production or pre-production) based on their risk tolerance. It also helps define task modes (manual, semi-automated, or fully automated). It guides the selection of fault injection tools, ensuring that organizations have the necessary evaluation methods to measure the impact of experiments and derive actionable insights for system improvement.
5.1.1. Execution Environment
This refers to the underlying infrastructure and system components where chaos engineering experiments are performed. These components are categorized into endpoints, grouped by resource types and cloud environments, representing various deployment models. This section will discuss both of these in detail.
Endpoints
These are physical or virtual devices connected to a computer network to exchange data. These include virtual machines (VMs), containers, and serverless functions (Microsoft, 2024). Each type of endpoint presents unique challenges that chaos engineering addresses by simulating failure conditions to improve the robustness of the system and ensure continuous operation.
Virtual machines (VMs) offer a layer of abstraction over physical hardware, allowing multiple independent operating systems to share resources efficiently (Sharma et al., 2016; Smith and Nair, 2005). However, VMs can experience performance issues when resources are overutilized, and they are vulnerable to hypervisor failures and migration difficulties (Azure Readiness, 2023; Fogli et al., 2023). Chaos engineering addresses these challenges by testing VMs under stress conditions such as CPU or memory exhaustion to assess if VMs can recover from failures or seamlessly migrate workloads (Kostic, 2024; Chen et al., 2024; Green, 2023).
Containers, which encapsulate applications and their dependencies in lightweight environments, provide greater efficiency and portability compared to VMs (Sharma et al., 2016; Koskinen et al., 2019). However, containers face resource contention, network instability, and orchestration failures, especially in platforms such as Docker and Kubernetes (Watada et al., 2019). Chaos engineering tests on containers simulate abrupt stoppages and network disruptions, helping to evaluate containers’ ability to recover within container orchestration systems while maintaining application performance  (Simonsson et al., 2021; Camacho et al., 2022; Bremmers, 2021; Kostic, 2024; Ikeuchi et al., 2020).
Serverless functions allow applications to be triggered by specific events and run in isolated containers without requiring infrastructure management (Castro et al., 2017; Taibi et al., 2020). While serverless architectures are scalable and cost-effective, they face challenges such as resource limitations and the reliance on third-party cloud providers (Rajan, 2020; van Eyk and Iosup, 2018; Kadikar, 2023; Camacho et al., 2022). Chaos engineering helps ensure that serverless functions can handle these challenges by simulating cold starts and execution failures, verifying that functions recover quickly and maintain performance under varying loads (Starr et al., 2022; Jernberg et al., 2020; Gill, 2022; Bedoya et al., 2023).
Cloud
Cloud environments are integral to modern Information Technology (IT) infrastructures, providing flexible and scalable resources to meet dynamic business demands (Avram, 2014; Garrison and Nova, 2017). These environments are typically classified into three main models: public, private, and hybrid clouds (Balasubramanian and Aramudhan, 2012; Gorelik, 2013). Each cloud model has unique characteristics, risks, and opportunities, where chaos engineering can address potential failures.
Public cloud platforms, such as AWS, Microsoft Azure, and Google Cloud, offer computing resources over the Internet that are shared between multiple organizations (Saraswat and Tripathi, 2020). This model enables businesses to scale their infrastructure on demand, eliminating the need for significant investments in physical hardware. Due to its cost efficiency and scalability, the public cloud has become a popular choice for many enterprises (Dutta and Dutta, 2019). However, relying on shared infrastructure introduces specific vulnerabilities. For example, a regional outage or failure in an availability zone can have widespread impacts, simultaneously affecting many customers (Paavolainen, 2016). In addition, many organizations depend on third-party services, such as application programming interfaces (APIs) or managed databases, and failures in these services can trigger further disruptions in their systems (Zhou et al., 2010). To address these risks, chaos engineering can simulate large-scale failures, such as regional outages or availability zone failures, to test how well systems can failover to other regions without significant service interruptions (Fogli et al., 2023). Moreover, by injecting failures into critical third-party services, chaos engineering enables organizations to evaluate the effectiveness of their recovery strategies, ensuring that operations can continue even in the face of service disruptions (Ji et al., 2023; Yu et al., 2021).
Private cloud environments, which provide dedicated infrastructure exclusively for a single organization, can be hosted on-premises or managed by third-party providers, offering greater control over data privacy, security, and customization (Jadeja and Modi, 2012; Gastermann et al., 2015). This model suits organizations with strict regulatory requirements or managing sensitive information (Pearson, 2013). Despite their benefits, private clouds often face scalability and geographic redundancy challenges, as they typically lack the built-in redundancy of public clouds, making them more susceptible to service interruptions due to hardware failures or resource constraints (Kumar and Goudar, 2012). Without adequate failover mechanisms, private cloud environments may struggle to maintain operational continuity during hardware or network failures (Rimal et al., 2009; Tivey et al., 2019). Chaos engineering can be used in private clouds to simulate hardware failures, such as server crashes or network partitions, to evaluate the system’s ability to recover from localized disruptions (Patel, 2022). By simulating resource exhaustion, such as CPU or memory overloads, organizations can test whether their infrastructure can scale to meet unexpected surges in demand, ensuring continuous service availability under pressure (Gill, 2021).
Finally, hybrid cloud environments combine the strengths of public and private clouds, allowing organizations to manage sensitive workloads on private infrastructure while leveraging the scalability of public clouds for nonsensitive or high-demand tasks (Goyal, 2014). This approach provides flexibility but introduces complexity in maintaining smooth integration between the two environments (Lu et al., 2014). A key challenge of hybrid clouds is to ensure seamless communication between public and private components, as issues such as network latency, connectivity problems, and data synchronization delays can arise when workloads are distributed across both environments, potentially leading to service disruptions (Annapureddy, 2010). Chaos engineering can address this issue by allowing organizations to simulate network interruptions or increased latency between public and private clouds (Fogli et al., 2023). These tests help to evaluate whether workloads can continue to operate despite connectivity issues (Torkura et al., 2020). Furthermore, chaos experiments that simulate delays or failures in data replication help ensure that the system maintains data consistency and operational continuity, even when part of the infrastructure is compromised (Camacho et al., 2022).
On-Premise
This refers to the computing infrastructure hosted and managed within an organization’s physical location rather than relying on cloud providers (Willard and Hutson, 2024; Nikolaidis et al., 2021; Väisänen, 2023). These systems offer greater control over hardware, data, and network resources, essential for industries with strict regulatory requirements. However, they also face hardware failures and scaling limitations, affecting the reliability of the systems (Byström, 2022). For example, an organization managing critical on-premise data might simulate a hardware failure, such as a server or disk crash, to test how quickly backup systems respond. If synchronization delays are found, the replication process can be optimized to ensure smooth transitions and compliance with data availability standards (Fogli et al., 2023; Dedousis et al., 2023). Similarly, scaling limitations can be addressed by testing how an on-premise infrastructure responds to various demand surges, such as a sudden spike in user traffic during an e-commerce company’s high-demand event (Azure Readiness, 2023; Gianchandani et al., 2022). This helps identify bottlenecks, optimize resource allocation, and ensure that the system remains responsive under varying load conditions (Naqvi et al., 2022; Wu, 2021; Sbiai et al., 2023; Wickramasinghe, 2023).
5.1.2. Experimentation Modes
Chaos engineering platforms offer three modes of experimentation, namely manual, semi-automated, and fully automated, each tailored to different levels of control, complexity, and system scale.
Manual experimentation is most appropriate when engineers require high precision and control, as they take full responsibility for designing, configuring, and executing chaos tests (Kostic, 2024). This approach is suited for organizations with little or no experience in chaos engineering or when highly specific, customized experiments are necessary (Durai, 2022). For example, an engineer might manually trigger a database failure or introduce network latency to observe the system response in real time (Jernberg et al., 2020; Bremmers, 2021). While this mode allows for fine-tuned control, it can be time-consuming and prone to human error, making it ideal for small-scale, customized tests or critical systems where close monitoring is essential (Ikeuchi et al., 2020; Chen et al., 2024).
Semi-automated experimentation provides a middle ground between manual control and complete automation, where engineers define the experiment parameters (Sondhi et al., 2021; Jernberg et al., 2020). At the same time, the platform automates tasks such as triggering tests and collecting data, making this mode ideal for experiments that benefit from automated execution while still requiring moderate human involvement to reduce manual effort (Kesim et al., 2020; Nikolaidis et al., 2021; Gill, 2021). For example, engineers might set up an experiment to test resource exhaustion on specific microservices, with the platform automatically running the test at predefined intervals, collecting performance metrics, and allowing them to focus on adjusting parameters and interpreting results (Pierce et al., 2021; Dedousis et al., 2023).
Fully automated experimentation suits large-scale systems that require frequent and continuous testing (Zhang et al., 2021a; Simonsson et al., 2021; Basiri et al., 2019). After engineers configure the initial setup, the chaos engineering platform takes over, seamlessly integrating experiments into the system operations, typically through the CI/CD pipelines (Konstantinou et al., 2021; Pulcinelli et al., 2023; Manna, 2021; Colyer, 2019; Azure Readiness, 2023). For example, after a code deployment or infrastructure update, the platform can automatically simulate a service outage or introduce network disruptions, then autonomously gather performance metrics, identify system weaknesses, and generate detailed reports, requiring engineers to intervene only when necessary (Dedousis et al., 2023; Bairyev, 2023; Kesim et al., 2020; Cotroneo et al., 2022; Nikolaidis et al., 2021; Patel, 2022).
5.1.3. Chaos Automation Strategy
Chaos engineering platform employs an automation strategy to ensure consistent and repeatable experiments, reduce manual effort, and enable continuous testing (Kostic, 2024; Wickramasinghe, 2023; Inc., 2023).
Custom Code
This involves developing programs or scripts to extend the capabilities of existing tools and gain precise control over system configurations and experiments (Myers et al., 2000; Thüm et al., 2014; Khankhoje, 2022). It allows engineers to simulate unique failure conditions, ranging from simple component failures, such as restarting a single service, to complex multi-step sequences involving multiple systems, tailored to the specific challenges of their architecture that standard tools may not address (Gill, 2021; Inc., 2023; Torkura et al., 2020; Azure Readiness, 2023; Patel, 2022; Nikolaidis et al., 2023). For example, custom code can terminate random containers in distributed environments such as Kubernetes, simulating unexpected service failures and testing the system’s recovery capacity by automatically restarting services (Bremmers, 2021; Fawcett, 2020; Simonsson et al., 2021; Vu et al., 2022). In more complex scenarios, custom code can initiate a service outage followed by a database failure or a network partition to observe how these failures impact system recovery and resilience (Wickramasinghe, 2023; Gill, 2021; Simonsson et al., 2021). Additionally, custom code can be integrated into CI/CD pipelines to simulate high resource consumption, such as CPU throttling or memory exhaustion, after each deployment, allowing early detection of performance bottlenecks before they affect production (Azure Readiness, 2023; Konstantinou et al., 2021; Manna, 2021; Kadikar, 2023; Palani and Gupta, 2023). This custom code for chaos experiments can be written in various programming languages, such as Python, Bash, or Go, depending on the complexity and specific requirements of the experiment (Siwach et al., 2022; Malik et al., 2023; Sondhi et al., 2021; Simonsson et al., 2021; Zhang et al., 2021b).
Auto-Integration of Third-Party Tools
This process automatically integrates external systems, such as monitoring, log, observability, alerting, incident management, and security platforms, into chaos engineering experiments (Torkura et al., 2020; Simonsson et al., 2021). This automated integration ensures continuous data collection without manual intervention, enabling real-time analysis during system disruptions or failures. As a result, teams can monitor performance, identify issues quickly, and respond more efficiently to failures (Palani and Gupta, 2023; Kostic, 2024). This process is vital to integrating monitoring tools that track essential system metrics such as CPU usage, memory consumption, and network performance (Abdul, 2024; Wu, 2021; Gianchandani et al., 2022; Zhang et al., 2019). Monitoring tools can automatically detect performance deviations and provide immediate feedback. For example, if a service is overloaded due to an unexpected spike in user requests, monitoring systems measure the increase in CPU usage and network load, helping the team assess the impact and determine whether the system can handle the demand or if further intervention is needed (Wickramasinghe, 2023; Fogli et al., 2023; Cotroneo et al., 2022; Gill, 2021; Siwach et al., 2022; Malik et al., 2023). Logging tools record vital events such as service failures, error messages, and recovery attempts, documenting the system’s response in real-time. For example, if network communication between two microservices fails, the logging systems capture the exact moment the failure occurred and show the error messages generated by failed retries (Cotroneo et al., 2022; Ma et al., 2023). This data helps teams analyze how the failure affected communication and whether recovery mechanisms worked effectively (Naqvi et al., 2022; Malik et al., 2023; Kostic, 2024; Palani and Gupta, 2023). Building on this, observability tools provide a broader view by correlating logs and metrics to show how different components of the system interact during failure scenarios. These tools enable teams to identify how failures propagate across interconnected services (Simonsson et al., 2021; Manna, 2021; Green, 2023; Kostic, 2024; Krishnamurthy, 2021). For example, if latency is introduced between services in a chaos experiment, observability systems track the delays at each hop between services and how this latency affects downstream components such as database queries or API responses. This tracing allows the team to see if minor delays in one service cause issues elsewhere (Durai, 2022; Katirtzis, 2022; Kostic, 2024; Chen et al., 2024; Nikolaidis et al., 2023; Frank et al., 2023). Furthermore, alert notification systems can ensure that relevant teams are promptly notified when critical failures occur or performance thresholds are exceeded (Tivey et al., 2019; Gianchandani et al., 2022; Palani and Gupta, 2023; Bremmers, 2021; Starr et al., 2022). For example, if a database connection pool reaches its maximum capacity during a high-load chaos test, the alert system triggers a notification to the database operations team, allowing them to address the issue before it causes a system-wide failure. This early warning helps prevent more significant disturbances (Green, 2023; Malik et al., 2023). Finally, incident management tools streamline the resolution process in conjunction with alerting by automatically generating tickets for failures detected during chaos tests. These tickets are assigned to the appropriate teams and tracked until the issue is resolved (Inc., 2023; Patel, 2022; Green, 2023; Le, 2022). For example, if a chaos experiment causes a loss of connection to external APIs, the incident management system creates a ticket assigned to the appropriate API support team, ensuring that the problem is tracked and resolved efficiently (Rivera et al., 2023; Siwach et al., 2022).
Workflow
A chaos engineering workflow can automate the controlled execution of experiment steps, such as setting up the environment, executing failures, monitoring, and cleanup, ensuring that the experiments are consistent and repeatable  (Camilli et al., 2022). This reduces the risk of errors and guarantees the reliability in each run. For example, automating a database outage simulation ensures that critical steps such as downgrading the database, monitoring the system response, and restoring functionality are executed reliably every time  (Wickramasinghe, 2023). In addition, workflows simplify complex experiments that involve multiple systems or services. In such cases, they can coordinate the simultaneous failure of various microservices, allowing teams to test the system’s ability to maintain service availability despite several disruptions  (Nikolaidis et al., 2021; Manna, 2021). Furthermore, a workflow engine ensures that each step is executed correctly while capturing and analyzing the results, making the experimentation process both structured and efficient (Torkura et al., 2020; Bairyev, 2023).
Prioritization
This involves categorizing system components into critical and non-critical groups to focus testing efforts on the areas where failures would have the greatest impact (Dedousis et al., 2023; Anwar and Arunachalam, 2019). Critical components are essential for maintaining core operations, and their failure would cause significant disruption (Fogli et al., 2023; Kesim et al., 2020). For example, API gateways and distributed databases are critical because these failures can disrupt communication between services, cause widespread outages, or lead to data loss and downtime, directly affecting users (Siwach et al., 2022). In contrast, non-critical components have minimal immediate impact on the core functionality of the system. For instance, log aggregation systems are non-critical; their failure may affect internal monitoring but will not disrupt user-facing services or essential system functions (Dedousis et al., 2023). By prioritizing critical components, testing can focus on areas where failure would result in the most severe consequences. In contrast, the resilience issues of non-critical components can be addressed less urgently, since their failure poses a lower risk to overall system stability (Wickramasinghe, 2023).
5.1.4. Deployment
The type of deployment selected for chaos experiments is a critical decision in chaos engineering, as it defines the testing environment and significantly affects the performance and stability of the system (Meiklejohn et al., 2021; Abdul, 2024; Zhang et al., 2019). Choosing the appropriate deployment type ensures that the experiments yield meaningful insights while minimizing the risks to live operations (Siwach et al., 2022; Wickramasinghe, 2023). This section identifies two primary deployment types: production and pre-production.
Production.
This refers to running tests or experiments in a live environment where real users actively interact with the System Under Test (SUT) (Zhang et al., 2019; Bremmers, 2021; Vu et al., 2022; Nikolaidis et al., 2023; Cotroneo et al., 2022). For example, deploying a new feature on an e-commerce website during sales periods allows teams to observe user interactions and system performance under actual conditions (Jernberg et al., 2020). A significant benefit of production deployment is the ability to uncover vulnerabilities that might remain hidden in more isolated testing environments (Wickramasinghe, 2023; Nikolaidis et al., 2021). For example, testing a service during peak traffic can reveal weaknesses that would not otherwise emerge, making it crucial to gain insight into how the SUT handles real-world disruptions and help identify areas for further optimization (Fogli et al., 2023; Chen et al., 2024; Simonsson et al., 2021). Production deployment introduces risks, as failures during testing can directly affect users and disrupt business operations (Roa, 2022; Jernberg et al., 2020). Production experiments must be carefully planned to mitigate these risks, such as scheduling tests during low-traffic periods and implementing automated rollback mechanisms to restore normal operations if problems occur (Sondhi et al., 2021; Bremmers, 2021).
Pre-production.
This environment closely mirrors the production environment, allowing teams to validate system functionality and performance before going live (Naqvi et al., 2022; Vu et al., 2022; Bremmers, 2021). An example of a pre-production environment is a staging environment, where the SUT can be deployed with the same configurations as production, enabling comprehensive checks on how it will perform under expected loads (Jernberg et al., 2020; Manna, 2021). Pre-production deployment allows for complex tests that might be too risky to execute in live environments (Bhaskar, 2022). For example, simulating a network partition can reveal how well the SUT responds to such failures, helping teams resolve potential problems before they reach production (Green, 2023). This added layer of testing ensures major releases and helps reduce the risk of failure once the system is live  (Kostic, 2024). One limitation of pre-production testing is that it may not accurately reflect the complexities of real-world operations (Bhaskar, 2022; Naqvi et al., 2022). Consequently, these tests should be viewed as a complement, rather than a substitute, to production testing (Zhang et al., 2019; Torkura et al., 2020; Gianchandani et al., 2022).
5.1.5. Evaluation Approaches
Evaluation refers to determining the effectiveness of chaos engineering experiments by measuring their impact on system stability, the system’s ability to handle induced failures, and the growth and maturity of the organization’s chaos engineering practices. Evaluation approaches can be categorized into qualitative and quantitative methods.
Quantitative Evaluation
This involves systematically measuring numerical data to assess a system’s performance under failure conditions. It focuses on metrics that can be measured, such as response time, error rates, and system availability (Garg et al., 2013). By comparing the behavior of the system during chaos experiments with its regular operation, engineers can identify weaknesses and areas for improvement based on objective data. We identified eight quantitative evaluation approaches, which include the following:
  • Response Time Metrics measure the total time it takes for a system to process a request, from when it is received to when the result is returned, encompassing both processing and response delivery (Frank et al., 2021a; Abdul, 2024; Frank et al., 2021b; Dua, 2024; Cotroneo et al., 2022; Green, 2023; Kostic, 2024; Mooney, 2023; Siwach et al., 2022; Camacho et al., 2022; Sondhi et al., 2021; National Australia Bank, 2020; Al-Said Ahmad et al., 2024; Kesim et al., 2020). For example, a baseline response time of 18 milliseconds might increase to 27 milliseconds during chaos experiments such as ’KillNodes’, where parts of the system are intentionally shut down to test disruptions, indicating slower request processing and completion (Torkura et al., 2021).
  • Latency Metrics measure the initial delay or waiting time between when a request is made and when the system begins processing that request (Fogli et al., 2023; Wickramasinghe, 2023; Kostic, 2024; Poltronieri et al., 2022; Bremmers, 2021; Cotroneo et al., 2022; Azure Readiness, 2023; Colyer, 2019). It does not include the total processing time, only the delay before the system starts responding (Cotroneo et al., 2022; Durai, 2022; Fawcett, 2020; Gianchandani et al., 2022; Frank et al., 2021a; Siwach et al., 2022; Yu et al., 2021). For example, latency can increase from 44 to 70 milliseconds during chaos experiments, signaling that the system takes longer to react to requests (Kassab et al., 2022).
  • Resource Utilization Metrics track the consumption of system resources, such as CPU and memory (Azure Readiness, 2023; Fogli et al., 2023; Torkura et al., 2020; Simonsson et al., 2021; Zhang et al., 2021a; Camacho et al., 2022; Pulcinelli et al., 2023; Wu, 2021; Gianchandani et al., 2022; Roa, 2022; Kadikar, 2023; Green, 2023; Zhang et al., 2019). For example, CPU usage may spike from 50% to 80% during chaos experiments, indicating an increased strain on the system due to failures  (Simonsson et al., 2021).
  • Error Rate Metrics measure the frequency of errors encountered during system operation (Gill, 2021; Zhang et al., 2021a; Cotroneo et al., 2022; Basiri et al., 2019; Rivera et al., 2023; Roa, 2022; Kostic, 2024; Vinisky, 2024; Colyer, 2019; Durai, 2022). For instance, an increase in the error rate from 0.5% to 5% during chaos experiments suggests that the system is struggling to handle failures effectively (Zhang et al., 2023).
  • Availability Metrics refers to a system’s ability to remain operational, measured by uptime and downtime (Torkura et al., 2020; Gill, 2021; Camacho et al., 2022; Nikolaidis et al., 2021; Ma et al., 2023). Uptime indicates how long the system is fully functional, while downtime shows when the system is unavailable (Shah, 2021; Durai, 2022; Abdul, 2024; Gianchandani et al., 2022).
  • Throughput Metrics measure the volume of data or transactions the system can handle in a given time (Kostic, 2024; Vinisky, 2024; Al-Said Ahmad et al., 2024; Torkura et al., 2020). For example, throughput might drop from 1,000 transactions per second to 600 during chaos experiments, leading to delays in processing user requests or data transfers (Kassab et al., 2022; Abdul, 2024).
  • Mean Time to Recovery (MTTR) measures the average time required to restore the system after a failure (Durai, 2022; Abdul, 2024; Starr et al., 2022; Gianchandani et al., 2022). A higher MTTR, indicating slower recovery, reveals areas that need faster recovery processes (Starr et al., 2022; Frank et al., 2021b; Camacho et al., 2022; Bhaskar, 2022). For example, if a cloud-based e-commerce platform experiences a server crash during a flash sale and MTTR increases from 2 to 10 minutes, it signals the need for more efficient recovery mechanisms to prevent revenue loss and customer frustration (Siwach et al., 2022).
  • Mean Time Between Failures (MTBF) refers to the average time that a system operates without failure, reflecting the reliability of the system (Bhaskar, 2022; Camacho et al., 2022). A lower MTBF suggests frequent failures, highlighting the need for improved stability (Kadikar, 2023). For example, if a healthcare system that manages patient records experiences failures every two hours, it points to the need for a more reliable system architecture to ensure uninterrupted access to patient data during critical procedures.
Qualitative Evaluation
This gathers non-numerical insights to assess system performance during failures, focusing on user feedback, team experiences, and business impacts. By understanding these aspects, engineers gain deeper insight into the broader effects of disruptions. This complements quantitative methods by adding meaning to numerical data. We identified four qualitative evaluation approaches, which include:
  • User Experience Evaluation involves gathering qualitative feedback on how system incidents impact end users (Gill, 2021; Ji et al., 2023). This may include conducting surveys, interviews, or usability tests to assess how failures affect users’ ability to interact with the system (Camacho et al., 2022; Kostic, 2024). For example, users might report frustration or difficulty in completing tasks during a chaos experiment that causes slow page load times or feature outages (Fawcett, 2020; Sbiai et al., 2023; Gill, 2022). This evaluation is crucial because it reveals how technical failures affect user satisfaction and retention, providing insight beyond response time or availability figures (Wu, 2021; Le, 2022; Cotroneo et al., 2022).
  • Business Impact Evaluation assesses how chaos experiments affect key business metrics, including revenue, customer churn, and operational efficiency (Inc., 2023; Jernberg et al., 2020). This involves understanding specific financial losses, customer complaints, and service-level breaches resulting from experiments(National Australia Bank, 2020; Palani and Gupta, 2023). For instance, a chaos experiment that causes downtime during a product launch could result in significant lost sales and a spike in customer service tickets (Ji et al., 2023; Starr et al., 2022). This evaluation is critical to linking technical failures with direct business consequences, enabling more targeted investment in stability measures where it matters most (Manna, 2021; Moon, 2022).
  • System Behavior and Observability focuses on how effectively system components behave and are monitored during chaos experiments (Anwar and Arunachalam, 2019; Yu et al., 2021; Meiklejohn et al., 2021; Mooney, 2023). It involves reviewing logs/traces and monitoring gaps to identify areas where failures are not easily detected or diagnosed (Ikeuchi et al., 2020; Soldani and Brogi, 2021; Zhang et al., 2019; Green, 2023). For example, an experiment might expose missing alerts or inaccurate telemetry data for critical services, making it harder for engineers to respond (Colyer, 2019; Meiklejohn et al., 2021; Basiri et al., 2019). This evaluation helps to ensure that the system’s observability tools are robust enough to detect problems early and aid in faster diagnosis and recovery (Dedousis et al., 2023; Fawcett, 2020).
  • Team Feedback captures insights from engineering teams on their ability to respond to failures during chaos experiments (Kostic, 2024; Fogli et al., 2023; Torkura et al., 2020). It evaluates the effectiveness of incident response, communication, and decision-making processes under system stress (Zhang et al., 2019; Simonsson et al., 2021). For example, feedback might highlight confusion in handling specific failure scenarios or delays in coordinating recovery efforts (Inc., 2023; Krishnamurthy, 2021; Chen et al., 2022). This evaluation is essential to identify operational weaknesses in real-time incident management and to improve team workflows and readiness for actual outages (Kesim et al., 2020; Wu, 2021; Fawcett, 2020).
5.2. Chaos Engineering Tools
We used the proposed taxonomy to evaluate 10 popular chaos engineering tools (T1-T10 in Table 10). We selected these tools by identifying the most-cited tools from our pool of reviewed papers, then ranking them using GitHub metrics, precisely the ’Most stars’ measure, to confirm their popularity and community engagement (Borges et al., 2016). Collectively, these tools had over 93,000 stars as of November 2024. These tools have been mapped according to their applicability to different execution environments, automation modes, strategies, and deployment stages. This categorization enables practitioners to make well-informed choices by aligning tool selection with specific infrastructure needs, operational requirements, automation capabilities, and system risk tolerance. Consequently, selecting the appropriate tool in a software development project becomes a crucial decision (Kavaler et al., 2019). For example, T1 and T2 provide manual and semi-automated chaos testing capabilities, supporting Kubernetes, Docker, and virtual machine setups. While T1 supports performing chaos experiments in live production environments, T2 focuses on testing environments, particularly pre-production setups. T3 is a retired tool designed initially for AWS but lacks support for Kubernetes. Its functions have been migrated to newer Netflix tools such as Chaos Monkey, Swabbie, and Spinnaker, making it better suited to traditional infrastructures. T4 and T5 both support Kubernetes and virtual machine environments in hybrid clouds. These tools offer full automation options, with features like workflow automation and custom integrations, making them ideal for chaos testing in pre-production and production environments. T6 and T7 focus on Kubernetes and serverless functions, offering manual, semi-automated, and fully automated modes. While T6 is optimized for hybrid cloud setups with advanced workflow customization, T7 is better suited for production resilience testing in Kubernetes environments. T8 and T9 emphasize adaptability across diverse infrastructures. T8 supports Kubernetes, serverless functions, and on-premise setups, focusing on manual testing modes with customizable workflows. T9, on the other hand, operates across virtual machines and Kubernetes environments, providing semi-automated and fully automated capabilities with workflow automation and custom integration options, making it suitable for public cloud systems. Finally, T10 excels in hybrid cloud environments, supporting Kubernetes and virtual machines with advanced automation and integration capabilities for pre-production and production stages. This taxonomy ensures that chaos engineering tools are mapped to their ideal use cases, helping practitioners choose the best tools for their systems. Our online appendix provides the complete list of tools considered.
Figure 9. A Taxonomy of Chaos Engineering Platforms.
[…]
A taxonomy of chaos engineering platforms.
Table 10. Top-10 Most Used Chaos Engineering Platforms.
Tool Code Tool Name Execution Environment Automation Mode Automation Strategy Deployment Stages
Endpoint On- Premise Cloud Manual Semi- Automated Fully Automated Prioritization Workflow Automation Custom Code Auto- Integration Pre- production Production
Docker Kubernetes
Virtual
Machine
Serverless
Function
Private Public Hybrid
T1 Chaos Monkey
T2 Toxiproxy
T3 Simian Army
T4 Chaos Mesh
T5 Chaosblade
T6 Litmus
T7 Pumba
T8 Chaos Toolkit
T9 Powerful seal
T10 Chaoskube
5.3. Adoption Practices
Adopting chaos engineering requires a structured approach, balancing best practices with avoiding common mistakes. To effectively implement chaos experiments and reduce the risk of system failures or disruptions, organizations should follow well-established guidelines for planning, executing, and evaluating these experiments  (Siwach et al., 2022; Green, 2023). Table 11 summarizes essential best practices, such as starting with small, controlled experiments, setting clear objectives, automating tests, and closely monitoring system behavior. Additionally, organizations should limit the scope of experiments, encourage cross-team collaboration, and continuously refine processes based on experiment outcomes. While following best practices is crucial, organizations must also avoid common pitfalls, such as those outlined in Table 12, which include failing to tailor experiments to specific systems, neglecting proper planning and hypothesis formation, and lacking adequate safety mechanisms. By adhering to best practices and avoiding these common pitfalls, organizations can confidently integrate chaos engineering into their systems and achieve successful outcomes (Siwach et al., 2022).
Table 11. Best Practices
No Best Practices Description Articles Total
1 Start Small and Scale Up Start with manageable tests, then expand gradually. (Fogli et al., 2023; Torkura et al., 2020; Bremmers, 2021; Patel, 2022; Gill, 2021; Chen et al., 2022; Siwach et al., 2022; Jernberg et al., 2020; Anwar and Arunachalam, 2019; Palani and Gupta, 2023; Kadikar, 2023; Green, 2023) 12
2 Define Clear Objectives Set precise goals to guide each experiment’s purpose and outcomes. (Torkura et al., 2020; Dedousis et al., 2023; Inc., 2023; Patel, 2022; Zhang et al., 2021a; Al-Said Ahmad et al., 2024; Kesim et al., 2020; Soldani and Brogi, 2021; Siwach et al., 2022; Jernberg et al., 2020; Jakkaraju, 2020; Dua, 2024; Kadikar, 2023; Green, 2023) 14
3 Automate Experiments Use automation tools to run chaos experiments regularly, ensuring consistency and frequency. (Torkura et al., 2020; Zhang et al., 2019; Inc., 2023; Gill, 2021; Wickramasinghe, 2023; Camacho et al., 2022; Zhang et al., 2021a; Al-Said Ahmad et al., 2024; Vu et al., 2022; Kassab et al., 2022; Ikeuchi et al., 2020; Naqvi et al., 2022; Malik et al., 2023; Soldani and Brogi, 2021; Siwach et al., 2022; Jernberg et al., 2020; Bhaskar, 2022; Gill, 2022; Fawcett, 2020; Bairyev, 2023; Sbiai et al., 2023; Azure Readiness, 2023; Roa, 2022; Green, 2023) 24
4 Track and Measure Results Monitor the system’s performance during experiments and collect data to see the impact. (Fogli et al., 2023; Torkura et al., 2020; Dedousis et al., 2023; Inc., 2023; Bremmers, 2021; Patel, 2022; Gill, 2021; Camacho et al., 2022; Torkura et al., 2021; Zhang et al., 2021a; Vu et al., 2022; Kassab et al., 2022; Ikeuchi et al., 2020; Malik et al., 2023; Soldani and Brogi, 2021; Siwach et al., 2022; Konstantinou et al., 2021; Jernberg et al., 2020; Hawkins, 2020; Manna, 2021; Azure Readiness, 2023; Kadikar, 2023; Green, 2023) 23
5 Involve all Key Stakeholders Include team members to ensure everyone is informed and aligned. (Torkura et al., 2020; Inc., 2023; Bremmers, 2021; Patel, 2022; Gill, 2021; Frank et al., 2021b; Siwach et al., 2022; Dua, 2024; Kadikar, 2023; Green, 2023; Gill, 2022; Kostic, 2024) 12
6 Document and Improve Record experiments to continually improve the system. (Torkura et al., 2020; Zhang et al., 2019; Inc., 2023; Bremmers, 2021; Patel, 2022; Gill, 2021; Siwach et al., 2022; Palani and Gupta, 2023; Kadikar, 2023; Green, 2023; Kesim et al., 2020) 11
Table 12. Bad Practices
No Bad Practices Description Articles Total
1 Misaligned Chaos Chaos experiments that don’t match the system yield misleading results. (Fogli et al., 2023; Torkura et al., 2020; Zhang et al., 2019; Dedousis et al., 2023; Vu et al., 2022; Nikolaidis et al., 2023; Ma et al., 2023; Sondhi et al., 2021; Jernberg et al., 2020; Konstantinou et al., 2021; Park et al., 2019; Zhang et al., 2023; Rivera et al., 2023; Patel, 2022; Le, 2022; Lardo, 2019) 16
2 Inadequate Experimentation Without planning, clear goals, and baseline tracking, assessing chaos test impacts becomes difficult. (Inc., 2023; Wickramasinghe, 2023; Gill, 2021; Bremmers, 2021; Jakkaraju, 2020; Le, 2022; Hawkins, 2020; Anwar and Arunachalam, 2019; Dua, 2024; National Australia Bank, 2020; Mooney, 2023; Starr et al., 2022; Palani and Gupta, 2023; Gill, 2022; Fawcett, 2020; Gianchandani et al., 2022; Bairyev, 2023; Abdul, 2024; Azure Readiness, 2023; Kadikar, 2023; Green, 2023; Aharon, 2024; Krivas and Portela, 2020; Kostic, 2024; Frank et al., 2021b; Jernberg et al., 2020; Konstantinou et al., 2021; Ji et al., 2023; Soldani and Brogi, 2021; Park et al., 2019; Wachtel, 2023) 31
3 Safety Mechanism Oversight Not testing rollback and safety features risks uncontrolled failures. (Inc., 2023; Wickramasinghe, 2023; Gill, 2021; Bremmers, 2021; Patel, 2022; Jakkaraju, 2020; Bhaskar, 2022; Katirtzis, 2022; Dua, 2024; Starr et al., 2022; Tivey et al., 2019; Palani and Gupta, 2023; Bairyev, 2023; Sbiai et al., 2023; Kadikar, 2023; Green, 2023; Vinisky, 2024; Pierce et al., 2021; Al-Said Ahmad et al., 2024; Vu et al., 2022; Ikeuchi et al., 2020; Chen et al., 2022; Naqvi et al., 2022; Soldani and Brogi, 2021; Konstantinou et al., 2021) 25
4 Insufficient Tooling Lacking observability, automation, and integration tools makes monitoring and analyzing chaos experiments difficult. (Inc., 2023; Wickramasinghe, 2023; Gill, 2021; Hawkins, 2020; Starr et al., 2022; National Australia Bank, 2020; Tomka, 2024; Dua, 2024; Mooney, 2023; Lardo, 2019; Gianchandani et al., 2022; Abdul, 2024; Azure Readiness, 2023; Roa, 2022; Kadikar, 2023; Green, 2023; Krivas and Portela, 2020; Kostic, 2024; Simonsson et al., 2021; Nikolaidis et al., 2021; Soldani and Brogi, 2021; Al-Said Ahmad et al., 2024; Vu et al., 2022; Naqvi et al., 2022; Malik et al., 2023; Camilli et al., 2022; Bremmers, 2021; Patel, 2022; Le, 2022; Zhang et al., 2021a; Cotroneo et al., 2022; Ikeuchi et al., 2020; Bedoya et al., 2023; Basiri et al., 2019; Pulcinelli et al., 2023; Jernberg et al., 2020; Anwar and Arunachalam, 2019; Durai, 2022; Jakkaraju, 2020) 39
5 Poor Stakeholder Collaboration Excluding key team members causes misalignment and resistance. (Bremmers, 2021; Starr et al., 2022; Le, 2022; Krishnamurthy, 2021; Katirtzis, 2022; Durai, 2022; Tivey et al., 2019; Gill, 2022; Green, 2023; Poltronieri et al., 2022; Frank et al., 2021b; Al-Said Ahmad et al., 2024; Jernberg et al., 2020; Kesim et al., 2020) 14
6. Current State of Chaos Engineering (RQ4)
This section first presents the classification of research in chaos engineering, then examines the number of sources published per year and venue type, and finally reviews the venue rankings and key contributors in the field.
6.1. Number of Sources by Research Type
To better understand the current state of chaos engineering research, we classified the primary studies based on the research classification framework proposed by Wieringa et al. (Wieringa et al., 2006). This framework includes the following categories:
  • Evaluation research: These studies investigate problems in software engineering practice, often uncovering new knowledge about causal or logical relationships among different phenomena. This type of research is conducted through case or field studies, field experiments, and surveys.
  • Philosophical papers: These studies introduce new conceptual frameworks or ways of thinking.
  • Solution proposal papers: These studies suggest techniques and argue for their relevance without full-scale validation.
  • Validation research: These studies rigorously examine the properties of a proposed solution that has yet to be implemented in practice, employing methodologically sound approaches such as experiments and simulations.
  • Opinion papers: These studies present the author’s views on what is good or bad about a particular issue and how things should be approached.
  • Personal experience papers: These studies describe what was done based on the author’s own experience with one or more projects rather than explaining why it was done.
The first author did the classification, which was then reviewed by the second author, with all discrepancies resolved through discussion. The results of this classification, shown in Figure 12, indicate that solution proposal papers were the most frequent type of research between 2019 and 2024. In the academic literature, solution proposal papers dominated, with nine papers in 2021 and 2022 and slightly fewer (eight) in 2023. The grey literature, on the other hand, showed a marked rise in experience reports, growing from six reports in 2022 to eight in 2023. This classification highlights research activities’ evolving distribution and focus, emphasizing the dominance of solution proposals in academic literature and the increasing role of practical experience reports in grey literature.
6.2. Number of Sources by Year and Venue Type
We plot the distribution of chaos engineering studies from 2019 to 2024 to examine its adoption pattern (see Figure 10). In 2019, academic publications were minimal, limited to conferences and workshops, while grey literature focused on blogs, reflecting an early exploratory phase. By 2020, academic interest grew, focusing on conferences, though grey literature centered around blogs. In 2021, academic publications peaked across journals, conferences, and workshops, marking the “early adopter” phase, while grey literature remained stable, illustrating the challenges of Moore’s Chasm (see Figure 1). In 2022, academic output slightly declined. Still, grey literature grew, diversifying with blogs and white papers, suggesting chaos engineering was beginning to bridge Moore’s Chasm and gain broader acceptance. By 2023, academic publications remained steady, though workshops declined, while grey literature expanded further, especially with white papers, signaling continued industry engagement. The slight decline observed in 2024 for both academic and grey literature suggests a possible temporary slowdown, as the year is not yet complete, and more publications may emerge (Werbińska-Wojciechowska and Winiarska, 2023).
A precise distribution emerges between academic and grey literature publications when examining venue types. Of the 44 scholarly publications, 28 (63.64%) were conference papers, 11 (25%) were journal articles, and 5 (11.36%) were workshops, reflecting a strong preference for conferences as the primary channel for academic dissemination. Among the 44 grey literature publications, 41 (93.18%) were blogs, while 3 (6.82%) were white papers, underscoring blogs as the dominant platform for informal knowledge sharing. These trends indicate that chaos engineering, which began as an emerging research area, has evolved into a mature field of interest (Varela-Vaca and Quintero, 2021), attracting significant attention from academia and industry.
Figure 10. Distribution of Academic and Grey Literature Studies over Years and Venue Type.
[…]
Distribution of academic(N = 44) and grey(N = 44) literature studies over years and venue type.
6.3. Publication trends and key Contributors in Academic and Grey Literature
We analyzed publication trends in chaos engineering by examining venues based on their type and ranking and identifying key contributors within academic and grey literature. This analysis provides insights into the maturity of the domain (Varela-Vaca and Quintero, 2021).
6.3.1. Academic Literature
First, we analyzed venues based on their name, ranking, and type (journal or conference). As shown in Table 13, prominent venues include key publishers such as IEEE and ACM, which are widely recognized for hosting high-impact conferences and journals in chaos engineering. The venues were categorized using the Computing Research and Education Association of Australasia (CORE) ranking for conferences (CORE 2023) [^3] https://portal.core.edu.au/conf-ranks/ and the Scimago ranking for journals (Scimago 2023) [^4] https://www.scimagojr.com/journalrank.php. Figure 12 presents the distribution of sources in these venues. The data indicates that the largest share of studies is found in unranked venues (12 sources), with a significant number published in Q1 venues reflecting solid academic interest. High-ranking venues, such as A and A*, include several prominent IEEE and ACM conferences, emphasizing the quality of research in the field. Lower-ranked venues, including B, C, and workshops, also contribute to the research landscape, although fewer studies are published in Q2 venues. Table 13 highlights the top 10 primary studies in chaos engineering, ranked by citation count. It should be noted that only the first author of each study is listed, even if there are multiple authors  (Varela-Vaca and Quintero, 2021). Study A42 from Sun Yat-Sen University, China, leads the list, presenting MicroRank, a tool for end-to-end latency issue localization in microservice environments. Following closely are studies A02 and A09 from the Hasso Plattner Institute in Germany, which explore chaos engineering applications in cloud security and continuous auditing and threat detection across multi-cloud environments. KTH Royal Institute of Technology, Sweden, is represented by three influential works, A03, A05, and A06, investigating fault injection and resilience in containerized systems and Java Virtual Machine (JVM) environments. Netflix contributes with study A36, focusing on automating chaos experiments within production environments, while Carnegie Mellon University’s A22 emphasizes service-level fault injection testing via the Filibuster tool. Additional contributions include A26 from Lund University, which presents an implementation framework for chaos engineering, and A35 from the University of Stuttgart, which proposes an approach for prioritizing chaos experiments using established risk analysis techniques. This collection of studies from institutions in countries such as China, Germany, Sweden, and the United States reflects a global commitment to advancing chaos engineering with applications to enhance system resiliency, strengthen cloud infrastructure, and improve reliability in production environments worldwide.
6.3.2. Grey Literature
Unlike traditional academic literature, which follows a structured format and is ranked using systems such as h-index or Scimago, grey literature lacks standardized organization. It cannot be assessed using typical academic metrics. Instead, alt-metrics (alternative metrics used to measure online attention and engagement), such as comments, likes, shares, and social engagement, are used to gauge the quality and impact of grey literature. Table 14 highlights the top five articles in grey literature based on these alt-metrics. Grey literature on chaos engineering is often published by well-known IT companies (e.g., Microsoft, Gremlin, and Capital One) on company blogs or online publishing platforms (e.g., LinkedIn and Medium). These platforms are essential venues for sharing industry-driven insights in formats such as technical articles, blog posts, and white papers. The alt-metrics used, such as likes, comments, and shares, help measure the influence and engagement of the content within the industry, indicating its impact on chaos engineering. In cases where social engagement data was limited, additional quality assessments were considered, including the reputation of the venues, the authors’ expertise, and the clarity of the content (see Table  3).
As shown in Table 14, the top five grey articles are authored by industry professionals who provide valuable insights into chaos engineering. Key contributors come from prominent organizations, including Lyft, National Australia Bank (NAB), Expedia, Chaos Toolkit, and Eat Takeaway.com. For example, the study G12 from Lyft, published on Medium, received 348 likes and 2 comments for its introduction of a chaos experimentation framework. The study G11 from NAB, also published on Medium, received 156 likes for its focus on enhancing observability in microservices during chaos experiments. Additional notable contributions include the study G15 from Expedia, whose article on Medium earned 128 likes for detailing Kubernetes-integrated chaos engineering frameworks. The study G25, authored on LinkedIn by Chaos Toolkit, received 52 likes and 52 comments, emphasizing the importance of human interaction in chaos experiments. Lastly, the study G14 from Just Eat Takeaway.com, published on Medium, received 24 likes and 1 comment, highlighting the discovery of hidden system vulnerabilities through chaos engineering. These industry contributions have significantly influenced the development of chaos engineering practices, providing practical insights directly from leaders in the field. As discussed earlier, the same quality assessment criteria were applied, focusing on venue reputation, author expertise, and content clarity, when data on social engagement were insufficient (see Table  3).
Table 13. Top-10 of Papers Ordered by Citations in Academic Literature (SC - SCOPUS, GS - Google Scholar).
Rank ID Title Author Affiliation SC GS
1 A42 MicroRank: End-to-End Latency Issue Localization with Extended Spectrum Analysis in Microservice Environments Guangba Yu Sun Yat-Sen University, China 50 72
2 A02 CloudStrike: Chaos Engineering for Security and Resiliency in Cloud Infrastructure Kennedy A. Torkura Hasso Plattner Institute, Germany 39 70
3 A09 Continuous Auditing and Threat Detection in Multi-cloud Infrastructure Kennedy A. Torkura Hasso Plattner Institute, Germany 28 53
4 A03 A Chaos Engineering System for Live Analysis and Falsification of Exception-Handling in the JVM Long Zhang KTH Royal Institute of Technology, Sweden 15 44
5 A05 Observability and Chaos Engineering on System Calls for Containerized Applications in Docker Long Zhang KTH Royal Institute of Technology, Sweden 17 40
6 A36 Automating Chaos Experiments in Production Ali Basiri Netflix in Los Gatos, California. 15 40
7 A26 Getting Started with Chaos Engineering - Design of an Implementation Framework in Practice Hugo Jernberg Lund University and ICA Gruppen AB, Lund/Solna, Sweden 7 21
8 A22 Service-level Fault Injection Testing Christopher S. Meiklejohn Carnegie Mellon University, Pittsburgh, PA, United States 14 19
9 A06 Maximizing Error Injection Realism for Chaos Engineering With System Calls Long Zhang KTH Royal Institute of Technology, Sweden 6 18
10 A35 Identifying and Prioritizing Chaos Experiments by Using Established Risk Analysis Techniques Dominik Kesim University of Stuttgart, Germany 8 16
Table 14. Top-5 Articles Ordered by Altmetrics (Comments / Likes/Claps) in Grey Literature.
Rank ID Comments Likes/Claps Author Affiliation Venue
1 G12 2 348 Kathan Shah & Subhra Mondal Lyft Medium
2 G11 - 156 Prateek Sachan National Australia Bank (NAB) Medium
3 G15 - 128 Nikos Katirzis Expedia Group Medium
4 G25 - 52 Russ Miles Chaos Toolkit Medium
5 G14 1 24 Mathew Moon Just Eat Takeaway.com Medium
Figure 11. Research Type of Academic (AL) and Grey (GL) Literature Studies.
Figure 12. Venue Ranking of Academic Literature Studies.
7. Discussion
In this section, we summarize the key findings of our review and discuss their implications. We also outline open issues in chaos engineering research and discuss potential threats to our study’s validity.
7.1. Implications of Findings
The analysis highlights five key implications for researchers and practitioners in chaos engineering:
  • Need for a Unified Definition: The literature reveals inconsistencies in defining chaos engineering. Although tech giants like Google, Amazon, and Netflix provide definitions that emphasize resilience testing through controlled failure injection, there is variation in focus across other companies, with some prioritizing production-level experimentation, self-healing capabilities, or anti-fragility. A unified definition can align research efforts and support the systematic adoption of chaos engineering practices, strengthening reliability across complex Information Technology (IT) environments (Siswanto et al., 2024; Jayakody and Wijayanayake, 2024; Jones, 2019).
  • Complementary Role to Traditional Testing: Chaos engineering should be viewed as a complement to traditional testing rather than a replacement (Bocetta, 2019; Palani and Gupta, 2023). While conventional testing may miss subtle vulnerabilities within complex systems, chaos engineering exposes these issues through proactive experimentation. Recognizing chaos engineering as an additional layer to traditional methods emphasizes its role in enhancing system resilience, especially in high-stakes, mission-critical environments. Integrating chaos engineering into testing workflows is encouraged for those aiming to build more robust and reliable systems.
  • Growing Alignment of Research Theory and Practice: The prevalence of solution-oriented papers in academic literature alongside experience reports in grey literature reflects an increasing alignment between research and practice in chaos engineering. This synergy allows researchers to refine models based on real-world applications while practitioners benefit from research-backed frameworks, thus bridging the gap between theory and implementation. As suggested in previous studies (Jernberg et al., 2020; Rosenthal and Jones, 2020), fostering industry-academia collaborations could further integrate academic rigor with practical insights, advancing chaos engineering as a more resilient and impactful field.
  • Targeted Tool Selection: The study suggests that organizing chaos engineering tools by factors like execution environment and automation level might help practitioners make informed decisions when selecting tools. This taxonomy enables organizations to choose tools that match their specific infrastructure, operational needs, and risk tolerance, minimizing disruptions, managing costs, and enhancing resilience testing across development stages. This targeted approach allows tools to align with specific testing phases, from initial experimentation to production-level assessments, ensuring that chaos engineering efforts are both cost-efficient and impactful, ultimately supporting system reliability.
  • Standardized Practices: The findings highlight the importance of a standardized approach in chaos engineering to ensure consistent and practical application across diverse environments. Following best practices, which contribute to reliable and predictable outcomes (Kumara et al., 2021), enables organizations to build resilience and systematically manage risks. These practices, such as starting with small, controlled experiments, setting clear objectives, and automating processes where possible, support organizations in achieving reliable outcomes crucial for managing complex systems. Conversely, neglecting these practices, such as misaligned experiment design or insufficient safety mechanisms, can lead to inconsistent testing and increased system vulnerability. These findings underscore the need to balance adherence to best practices with avoiding common pitfalls, ensuring that chaos engineering aligns with broader operational goals to foster stability and reliability.
  • Insights through Quantitative and Qualitative Fusion: An essential implication of the findings is that combining quantitative and qualitative evaluation methods provides a comprehensive framework for assessing and improving system stability in chaos engineering. Quantitative metrics offer objective data on performance under stress, such as error rates, MTTR, and availability. At the same time, qualitative insights, including user experience and team feedback, reveal the broader impacts of system failures. This combined approach helps organizations identify technical vulnerabilities and understand their user and business impacts, enabling targeted investments in both stability and operational efficiency to align improvements with business goals and user expectations.
7.2. Open Research Issues and Practical Challenges in Chaos Engineering
Our review identifies several unresolved issues in chaos engineering, including research gaps and organizational challenges. We examine these issues from three perspectives: people, process, and technology.
7.2.1. People
In chaos engineering, people include engineers, managers, and other team members who collaborate to design and run experiments, ensuring that the systems are thoroughly tested (Moon, 2022; Nikolaidis et al., 2021; Graglia, 2024). First, some organizations struggle with poor communication and collaboration between teams, leading to incomplete or ineffective experiments  (Zhang et al., 2019). Encouraging more vital teamwork and shared ownership could improve results by fostering better department communication (Gill, 2021; Katirtzis, 2022; Kostic, 2024). Another challenge is the shortage of specialists, making it difficult for organizations to carry out chaos experiments. This gap could be addressed by creating comprehensive training programs and certifications that allow staff to develop the necessary expertise (Dedousis et al., 2023; Nikolaidis et al., 2021; Yu et al., 2021). Finally, cultural resistance is expected, as fear of failure and blame often hinder wider adoption  (Simonsson et al., 2021; Rosenthal and Jones, 2020). Implementing cultural change initiatives that promote learning from controlled failures and reduce fear of backlash can create a more open environment for experimentation (Green, 2023).
7.2.2. Process
The process aspect refers to the specific steps and methodologies, such as planning, testing, and reporting, that guide the implementation of chaos experiments (Bairyev, 2023; Kostic, 2024; Olmstead, 2023). First, the lack of standardized guidelines across industries makes consistent application difficult (Nikolaidis et al., 2021; Cotroneo et al., 2022; Kostic, 2024). Developing a set of best practices and frameworks would help organizations measure progress and improve the effectiveness of these practices (Shah, 2021; Bremmers, 2021). The research shows that many organizations struggle to prioritize chaos experiments alongside daily operations, often finding it difficult to allocate time and resources while maintaining operational stability. Establishing transparent processes for scheduling and running these experiments during non-critical periods or as part of routine maintenance can help mitigate the issue (Kesim et al., 2020).
7.2.3. Technology
Technology refers to the systems and infrastructure used to design, execute, and monitor failure scenarios, supporting and amplifying human efforts and processes (Graglia, 2024; Olmstead, 2023). First, chaos experiments are resource-intensive and require significant computational power and storage. This makes it difficult for organizations to allocate infrastructure without affecting critical operations  (Dedousis et al., 2023; Nikolaidis et al., 2021). To manage this, companies can schedule experiments during low-traffic periods or use scalable cloud resources to reduce strain  (Kostic, 2024; Gill, 2021). Second, deploying chaos tools requires specialized knowledge, making it challenging to scale experiments consistently (Zhang et al., 2019; Yu et al., 2021). This challenge can be alleviated with better documentation, training, and platforms that simplify deployment for teams with less expertise (Bairyev, 2023). Third, many chaos engineering tools lack automation, making setup and execution time-consuming and prone to errors. Automating processes like fault injection and monitoring can reduce the burden on engineers, minimize mistakes, and improve scalability. Organizations can address this through commercial tools or internal solutions (Bairyev, 2023; Cotroneo et al., 2022; Shah, 2021). Finally, while Artificial Intelligence (AI) holds promise for further streamlining experiment design and execution, research in this area is still limited. Using AI could reduce errors, improve efficiency, and help organizations scale their resilience efforts more effectively (Zhang et al., 2023).
7.3. Threats to Validity
We considered several potential threats to the external, construct, internal, and conclusion validity (Wohlin et al., 2012) that may impact our study.
7.3.1. Threats to External Validity
External validity refers to the extent to which the study results are applicable in a broader context (Wohlin et al., 2012). Since our multivocal review incorporated studies from various online sources, our findings may have limited applicability across the wider landscape of practices and general disciplines in chaos engineering, thus posing a threat to external validity. To enhance the external validity of our findings, we conducted multiple feedback sessions throughout the analysis of the selected studies. The discussions of each session were analyzed, providing qualitative insights that allowed us to refine our research methods and the applicability of our results. Furthermore, we compiled an online appendix with all artifacts generated during our analysis, including a complete list of sources, codes, and groups (see Section  2), making them publicly accessible for further examination. Another potential threat is the risk of missing relevant literature (grey and scientific) due to alternative terminology for related concepts that may not match our search terms. We explicitly included relevant associated synonyms and terms in our search strings to address this threat. We also leveraged search engine features that inherently consider variations of terms in a search string. The items identified were rigorously assessed based on multiple quality dimensions (Wohlin, 2014). Finally, while there are numerous chaos engineering tools, we limited our analysis to ten due to practicality. To mitigate this limitation, we selected the ten most relevant tools currently used by practitioners. A detailed comparison of these chosen tools may provide additional insight, which we propose for future research.
7.3.2. Threats to Construct and Internal Validity
Construct validity pertains to the generalizability of the study’s constructs, while internal validity relates to the credibility of the methods used for data analysis, including possible biases (Wohlin et al., 2012). To reduce potential threats to construct and internal validity, we implemented multiple triangulation rounds, inter-rater reliability assessments, and quality control measures (see Section 2). These methods helped minimize potential biases, such as observer and interpretive biases, thereby enhancing the validity of our data analysis. Additional studies and codes were incorporated, although no further inter-rater evaluation was conducted. In particular, some risk of observer bias remains inherent in this methodology (Kumara et al., 2021).
7.3.3. Threats to Conclusion Validity
Conclusion validity refers to the extent to which the conclusions of the study are credibly supported by the available data  (Wohlin et al., 2012). To address this threat, we used thematic coding and inter-rater reliability assessments to reduce observer and interpretive biases, aiming to analyze the data thoroughly. Furthermore, the study conclusions were independently derived by different authors and subsequently reviewed and validated against the primary studies selected during a feedback session.
8. Related Surveys
In this section, we consider key surveys on resilience, robustness, and fault injection techniques, exploring their relationship to the proactive focus of chaos engineering. For example, Natella et al. conducted an extensive survey on Software Fault Injection (SFI) techniques, categorizing methods by fault types, injection approaches, and application domains. These techniques are valuable for assessing software reliability in controlled settings. However, SFI techniques are generally applied in pre-production or isolated environments rather than in the live, iterative fault-testing scenarios central to chaos engineering (Natella et al., 2016). Another study, conducted by Colman-Meixner et al. (Colman-Meixner et al., 2016), explores resilience strategies across cloud infrastructure layers, such as physical servers, virtualization, and application components. Their work proposes a multi-layered framework with redundancy and failover mechanisms to ensure continuity during failures. While this approach aligns with the resilience goals of chaos engineering, the focus remains on structured recovery strategies within controlled environments rather than proactive fault injection in live settings. Laranjeiro et al. (Laranjeiro et al., 2021) reviewed robustness assessment techniques for systems such as operating systems and middleware and discussed methods such as fault injection and API testing to evaluate dependability under stress. These approaches aim to identify weaknesses but emphasize static testing, which does not include the proactive, live-system experiments that chaos engineering relies on. More recently, Rouholamini et al. (Rouholamini et al., 2020) presented a study on proactive self-healing techniques in cloud computing, categorizing strategies such as automated fault detection, anomaly recognition, and self-correction mechanisms. These techniques improve system resilience by addressing faults as they arise, but differ from the chaos engineering approach of using intentional fault injection to reveal latent vulnerabilities preemptively.
In general, there is a significant knowledge gap in the area of proactive and iterative resilience testing directly within live production environments. Moreover, there are no comprehensive reviews of the literature on chaos engineering. Therefore, this review paper systematically analyzes, assesses, and summarizes the existing literature to clarify the unique contributions of chaos engineering to the domain of system resilience.
9. Conclusion and Future Work
As modern distributed systems become increasingly complex, chaos engineering offers a proactive approach to ensure stability by introducing controlled failures that reveal system weaknesses, prevent outages, and support high performance under stress. This study bridges the gap between research insight and industry practices through a multivocal literature review (MLR) of 88 sources from academic and grey literature. To promote clarity across academia and industry, we proposed a unified definition of chaos engineering, reinforcing its role as a vital complement to traditional testing. We identified the main functionalities of a chaos engineering platform and compiled a set of principles that can guide systematically planning and conducting chaos experiments. In addition, we presented the key components of a chaos engineering platform and identified 12 quality requirements to guide the selection and evaluation of the platform. We also identified eight qualitative and four quantitative metrics essential for evaluating the impact of chaos experiments, ensuring that chaos engineering meaningfully enhances system robustness. Our review highlights four technical and three socio-technical challenges that chaos engineering seeks to resolve. To support organizations in structuring chaos engineering practices, we developed a taxonomy that organizes tools and techniques by environment, automation mode, automation strategy, and deployment stage while identifying the ten most commonly used tools within this framework. This categorization is valuable to practitioners and researchers when selecting tools that align with specific infrastructure needs and risk profiles. In addition, we identified six best practices and five common pitfalls in the implementation of chaos engineering. These practices and potential pitfalls guide organizations in optimizing their chaos engineering initiatives to foster safer and more robust systems. Our findings also reveal that solution-oriented research in academia and practical experience in industry converge, underscoring chaos engineering’s essential role in strengthening modern distributed systems. Finally, we outlined several research and practical challenges that must be addressed to realize chaos engineering’s vision entirely. We believe that these identified open issues can guide future research in the domain of chaos engineering.
In future work, we plan to use GitHub repository mining to identify various open-source projects that implement chaos engineering practices and utilize associated tools. This exploration will complement the ten tools analyzed in this literature review, offering broader insights into the prevalence and practical application of chaos engineering in diverse projects. Additionally, we aim to apply chaos engineering to improve the robustness of AI-enabled systems by evaluating how such systems handle disruptions in accuracy, adaptability, and consistency, ensuring that they continue to deliver reliable performance even under challenging conditions.

### v2

@@
 Abstract.
-Organizations, particularly medium and large enterprises, typically today rely heavily on complex, distributed systems to deliver critical services and products.
+Organizations, particularly medium and large enterprises, typically rely heavily on complex, distributed systems to deliver critical services and products.
 However, the growing complexity of these systems poses challenges in ensuring service availability, performance, and reliability.
-Traditional resilience testing methods often fail to capture modern systems’ intricate interactions and failure modes.
+Traditional resilience testing methods often fail to capture the intricate interactions and failure modes of modern systems.
 Chaos Engineering addresses these challenges by proactively testing how systems in production behave under turbulent conditions, allowing developers to uncover and resolve potential issues before they escalate into outages.
-Though chaos engineering has received growing attention from researchers and practitioners alike, we observed a lack of a comprehensive literature review.
-Hence, we performed a Multivocal Literature Review (MLR) on chaos engineering to fill this research gap by systematically analyzing 88 academic and grey literature sources published from January 2019 to April 2024.
-We first used the selected sources to derive a unified definition of chaos engineering and to identify key capabilities, components, and adoption drivers.
-We also developed a taxonomy for chaos engineering and compared the relevant tools using it.
-Finally, we analyzed the state of the current chaos engineering research and identified several open research issues.
+Though chaos engineering has received growing attention from researchers and practitioners alike, we observed a lack of reviews that synthesize insights from both academic and grey literature.
+Hence, we conducted a Multivocal Literature Review (MLR) on chaos engineering to address this research gap by systematically analyzing 96 academic and grey literature sources published between January 2016 and April 2024.
+We first used the chosen sources to derive a unified definition of chaos engineering and to identify key functionalities, components, and adoption drivers.
+We also developed a taxonomy for chaos engineering platforms and compared the relevant tools using it.
+Finally, we analyzed the current state of chaos engineering research and identified several open research issues.
 Chaos Engineering, Fault Injection, Distributed System, Resilience, Multivocal Literature Review, Microservice
-copyright: noneisbn: 978-1-4503-XXXX-X/18/06ccs: Software and its engineering Software testing and debugging
+copyright: acmlicensedisbn: 978-1-4503-XXXX-X/18/06ccs: Software and its engineering Software testing and debugging
 1. Introduction
@@
 The 2017 ITIC (Information Technology Intelligence Consulting) survey also reported that 98% of organizations experienced downtime with financial losses exceeding \$100,000 per hour (Inc., 2023).
-An effective approach to overcoming the challenges of ensuring reliable and resilient cloud-based systems is chaos engineering, which proactively tests system resilience by introducing controlled failures into production environments (Bhaskar, 2022; Fawcett, 2020; Palani and Gupta, 2023).
+An effective approach to overcoming the challenges of ensuring reliable and resilient cloud-based systems is chaos engineering, which proactively tests system resilience by introducing controlled failures into production environments (Bhaskar, 2022; Fawcett, 2020; Palani and Gupta, 2023a).
 This approach was pioneered by Netflix in 2008 to fortify its video streaming infrastructure.
 Since then, chaos engineering has grown and extended its influence, evolving from its original focus on cloud services and microservice architecture to broader applications across industries (Bairyev, 2023).
-Today, this approach is widely employed in critical sectors like fintech and healthcare, with major technology companies such as LinkedIn, Facebook, Google, Microsoft, Amazon, Slack, and others adopting it to enhance their system resilience, improve customer experience, and drive business growth.
+Today, this approach is widely employed in critical sectors, such as fintech and healthcare, with major technology companies like LinkedIn, Facebook, Google, Microsoft, Amazon, and Slack adopting it to enhance system resilience, improve customer experience, and drive business growth.
 (Inc., 2023; Patel, 2022; Roa, 2022; Fawcett, 2020; Al-Said Ahmad et al., 2024; Bremmers, 2021; Bairyev, 2023).
-Figure 1 illustrates chaos engineering’s adoption through Moore’s ’Crossing the Chasm’ model (Moore and McKenna, 1999).
-Adoption begins with Innovators (early risk-takers) and Early Adopters (those seeking competitive advantage), followed by the Early Majority (who adopt once proven), the Late Majority (who wait until widely supported), and finally, Laggards (the last to adopt, often only when essential).
-The critical hurdle, Moore’s Chasm, is where many approaches fail due to complexity and unclear benefits.
-Crossing this chasm leads to mainstream adoption and integration across the SDLC (Software Development Lifecycle) and industries (Mukkara, 2022).
-Figure 1. Chaos Engineering in the Context of Geoffrey Moore’s Technology Adoption Lifecycle Model from Crossing the Chasm  (Mukkara, 2022).
-[…]
-Chaos Engineering on the Technology Adoption Lifecycle
-We observed many researchers and practitioners have explored chaos engineering in various contexts, discussing its implementations and challenges through research papers and grey literature such as online blogs, technical papers, and forums.
+We observed that many researchers and practitioners have explored chaos engineering in various contexts, discussing its implementations and challenges through research papers and grey literature such as online blogs, technical papers, and forums.
 However, despite its increasing adoption, much of the existing knowledge remains dispersed across the literature.
@@
 In particular, as chaos engineering emerged from the industry and most tools are from the industry, the grey literature is a highly relevant and valuable source of information.
-In this MLR, we selected and analyzed 88 sources, spanning academic and grey literature sources.
+In this MLR, we selected and analyzed 96 sources, encompassing both academic and grey literature.
+We selected 2016 as the starting point, when chaos engineering principles were first formally articulated by Basiri et al (Basiri et al., 2016).
 Our findings provide a detailed overview of chaos engineering, including its core principles, quality requirements, and foundational components defined by researchers and industry experts.
 We also examine the motivations behind adopting chaos engineering, the specific resilience challenges it addresses, and the practices and evaluation methods that support its effective implementation.
-Additionally, we introduce a taxonomy that can help practitioners and decision-makers make informed decisions when comparing and selecting chaos engineering tools.
+Additionally, we introduce a taxonomy of chaos engineering platforms and their capabilities, designed to support informed decision-making around tool adoption and deployment strategies.
 By analyzing the benefits and shortcomings of current approaches in the chaos engineering domain, this MLR contributes to a thorough understanding of chaos engineering and provides guidelines for future studies to advance this critical field.
 The remainder of the paper is organized as follows.
-Section  2 presents the research methodology used in this study, while Section 3 provides a comprehensive definition of chaos engineering, along with its functionalities, quality requirements, and core components.
+Section 2 outlines the background and motivation for this review, including a discussion of related surveys, the distinction between chaos engineering and traditional testing, and the need to extend the existing body of knowledge on chaos engineering.
+Section 3 presents the research methodology used in this study, while Section 4 provides a comprehensive definition of chaos engineering, along with its functionalities, quality requirements, and core components.
 Section 5 focuses on the motivations for chaos engineering and the specific challenges it addresses.
@@
 Section 7 examines the current state of chaos engineering in research and industry.
-Section  7 discusses the implications of our findings, highlighting key insights, identifying open issues, and addressing potential threats to validity.
-Finally, Section 8 summarizes the related literature surveys, and Section  9 concludes the paper.
+Section  8 discusses the implications of our findings, highlighting key insights, identifying open issues, and addressing potential threats to validity.
+Finally, Section  9 concludes the paper.
+2. Background and Motivation
+This section outlines the motivations for conducting this review by examining the limitations of existing literature (Section2.1), explaining how chaos engineering diverges from traditional testing methods (Section 2.2), and establishing the novelty and timeliness of this multivocal literature review (Section 2.3).
+2.1. Related Work and Survey Gap
+In this section, we consider key surveys on resilience, robustness, and fault injection techniques, exploring their relationship to the proactive focus of chaos engineering.
+For instance, Natella et al.
+(Natella et al., 2016) surveyed software fault injection techniques, categorizing faults by origin such as software, hardware, and environment, and by injection approach such as code changes or fault effect injection including state and interface errors.
+However, their focus is on pre-deployment environments (e.g., simulators, testbeds), whereas chaos engineering involves live, iterative experiments in production-like systems.
+Similarly, Colman-Meixner et al.
+(Colman-Meixner et al., 2016) focus on cloud resilience through redundancy and failover, but explore reactive recovery rather than proactive experimentation, as in chaos engineering.
+Laranjeiro et al.
+(Laranjeiro et al., 2021) consider static robustness techniques such as fuzzing and API testing, which lack real-time, adaptive feedback loops.
+Ruholamini et al.
+(Rouholamini et al., 2020) surveyed self-healing in cloud-native systems, emphasizing failure detection and recovery rather than proactive failure exploration.
+More recently, Arsecularatne and Wickramarachchi (Arsecularatne and Wickramarachchi, 2023) conducted a systematic review and qualitative study on the adoptability of chaos engineering in DevOps.
+Their work highlights organizational challenges and potential benefits but does not provide a structured analysis of tools, architectures, or evaluation metrics.
+Similarly, Mhatre et al.
+(Mhatre et al., 2024) examined chaos engineering within DevSecOps and SRE (Site Reliability Engineering) contexts, emphasizing conceptual alignment and use cases, but without offering a comprehensive taxonomy or resilience assessment frameworks.
+In summary, existing reviews examine fault injection or discuss chaos engineering within DevOps contexts.
+However, they offer little or no clear insight into its unified definition, functionalities, architectural components, quality attributes, comprehensive tool taxonomy, or the technical and organizational challenges of implementing it in production-like environments.
+This review addresses these gaps by presenting the first multivocal literature review (MLR) on chaos engineering, integrating both academic and industry sources to inform future research and practical adoption.
+2.2. Chaos Engineering vs. Traditional Software Testing Techniques
+Traditional software testing techniques, including unit, integration, stress, and robustness testing, primarily focus on validating correctness and performance under expected conditions in pre-deployment environments (Basiri et al., 2019; Shekhar, 2024).
+These approaches typically operate in isolated, controlled settings and verify that system components behave correctly with known inputs (Ji et al., 2023; Shekhar, 2024).
+However, in modern distributed, cloud-native systems, such methods often fail to expose cascading or emergent failures caused by asynchronous interactions and dynamic runtime behaviors (Migirditch et al., 2022; Ji et al., 2023; Fogli et al., 2023).
+Chaos engineering addresses these limitations by introducing controlled faults into production or production-like systems to explore system resilience (Zhang et al., 2019; Torkura et al., 2020; Fogli et al., 2023).
+Rather than validating functionality using predefined test scripts, chaos experiments simulate failures, such as service crashes, latency spikes, or regional outages,to evaluate system behavior and recovery under adverse runtime conditions (Zhang et al., 2019; Migirditch et al., 2022; Shekhar, 2024).
+This methodology helps surface hidden vulnerabilities that traditional test environments often miss (Ji et al., 2023; Parsons, 2021).
+Importantly, chaos engineering is not a replacement for conventional testing, but a complement.
+While traditional methods confirm system correctness under expected conditions, chaos engineering validates system behavior under known faults and real-world failure scenarios  (Basiri et al., 2019; Fogli et al., 2023).
+This paradigm shift—from pre-deployment pass/fail validation to post-deployment resilience assessment—is what makes chaos engineering essential for today’s complex, distributed infrastructure (Zhang et al., 2019; Ji et al., 2023; Fogli et al., 2023).
+To clarify these distinctions, Table 1 highlights key differences between traditional testing and chaos engineering.
+The comparison focuses on three main areas: the testing environment (pre-production vs.
+production-like systems), the primary objective (functional verification vs.
+uncovering hidden system weaknesses), and the types of failures addressed (expected vs complex real-world faults).
+Table 1. Comparison Between Traditional Software Testing Techniques and Chaos Engineering
+Category Traditional Testing Chaos Engineering Articles Referenced
+Testing Environment Pre-production or simulated systems.
+Production or production-like systems.
+(Bairyev, 2023; Zhang et al., 2019; Simonsson et al., 2021; Zhang et al., 2021b; Gill, 2021; Bocetta, 2019; Bhaskar, 2022; Krivas and Portela, 2020; Wilms, 2018)
+Testing Objective Verifies specific functions or components against pre-defined requirements and standards.
+Identifies obscure system weaknesses that may only become apparent during turbulent failure conditions.
+(Torkura et al., 2020; Inc., 2023; Bairyev, 2023)
+Types of Failures Focuses on expected and known usage scenarios and errors to verify how the system behaves under normal or predictable conditions.
+Simulates real-world failures (e.g., network outages, high load, hardware issues) to see how the system reacts to these changes.
+(Torkura et al., 2020; Pierce et al., 2021; Ji et al., 2023; Konstantinou et al., 2021; Dua, 2024; Miles, 2019a; Bairyev, 2023)
+2.3. Research Contribution and Novelty
+Despite the growing popularity of chaos engineering (Monroe, 2019), there remains a lack of unified academic frameworks that synthesize its principles, classify its tool landscape, or integrate its practices across both research and industry domains.
+To the best of our knowledge, this paper presents the first Multivocal Literature Review (MLR) on chaos engineering, integrating both academic studies and grey literature.
+Our contributions are fourfold.
+First, we synthesize and unify definitions, core principles, functionalities, and architectural components of chaos engineering from academic and industry perspectives (RQ1).
+Second, we identify the key technical and socio-technical challenges driving its adoption (RQ2).
+Third, we propose a taxonomy of chaos engineering tools, classify evaluation methods, and highlight adoption practices (RQ3).
+Fourth, we analyze publication trends, venue types, and key contributors to assess the current state of research and practice (RQ4).
+Together, these contributions provide a foundational reference on chaos engineering that bridges research and practice, supports rigorous evaluation of chaos engineering tools, enables their systematic adoption in industry, and offers directions for future research.
 3. Research Methodology
@@
 (Garousi et al., 2019) to systematically collect and analyze multivocal literature on chaos engineering.
-We also drew methodological inspiration from other multivocal literature reviews, such as Buck et al.
+We also drew methodological inspiration from other multivocal literature reviews, such as those by Buck et al.
 (Buck et al., 2021) and Islam et al.
 (Islam et al., 2019).
-Figure  2 illustrates our MLR methodology, comprising several steps divided into three phases: planning and designing the review protocol, executing the review, and documenting the findings.
+Figure 1 illustrates our MLR methodology, comprising several steps divided into three phases: planning and designing the review protocol, executing the review, and documenting the findings.
 The remainder of this section provides a detailed discussion of each step.
 3.1. Research Identification
-We employed a search strategy to identify appropriate literature grounded in a series of research questions in Table  1.
+We employed a search strategy to identify appropriate literature grounded in a series of research questions in Table 2.
 The first three questions cover the various aspects of chaos engineering, including its definitions, key features, benefits, challenges, use cases, tools and techniques, and best and bad practices.
@@
 RQ1.1.
-Which core principles and primary functionalities are most commonly associated with chaos engineering?
+Which core activities and primary functionalities are most commonly associated with chaos engineering?
 RQ1.2.
@@
 What quality requirements are essential for an effective chaos engineering platform, and how are they quantified?
-This question clarifies the definition of chaos engineering, exploring its core principles, key components, and quality requirements.
+This question clarifies the definition of chaos engineering, exploring its core activities, key components, and quality requirements.
 These elements build a foundation for understanding, implementing, and evaluating chaos engineering solutions.
@@
 These libraries serve as primary sources for software engineering studies (Chen et al., 2010).
-We performed multi-keyword searches targeting titles, abstracts, and keywords for papers published between January 2019 and April 2024.
+Following the formal articulation of chaos engineering principles by Basiri et al.
+in 2016 (Basiri et al., 2016), we limited our search to publications from January 2016 to April 2024.
+We performed multi-keyword searches over titles, abstracts, and keywords within this range.
 Additionally, Google Scholar was used to pinpoint any relevant literature that was not included in the previous searches.
@@
 We adapted and applied the Google Search search string “(Chaos AND Engineering)” for grey literature to capture relevant non-peer-reviewed sources.
+Table 3. Search Strings and Results.
+Search Strings ACM DL IEEE Xplore Scopus Web of Science Total
+"chaos engineering" OR "chaos test" OR "chaos experiment" OR "chaos toolkit" OR "chaos platform" OR "chaos mesh" OR "chaos integration" OR "chaos architecture" OR "chaos infrastructure" OR "chaos injection" OR "chaos tool" 108 81 142 57 388
+"chaos engineering" AND ("trends" OR "developments" OR "progression") 30 0 20 0 50
+"chaos engineering" AND ("leader" OR "pioneer" OR "expert" OR "contributor") 39 3 4 0 46
+"chaos engineering" AND ("phases" OR "stages" OR "steps" OR "pipeline") 71 6 11 1 89
+"chaos engineering" AND ("challenges" OR "issues" OR "problems") 91 19 37 4 151
+"chaos engineering" AND ("applications" OR "benefits") 89 22 43 9 163
+"chaos engineering" AND ("solutions" OR "approaches" OR "methods") 84 22 60 4 170
+"chaos engineering" AND ("tools" OR "techniques") 92 30 42 10 174
+"chaos engineering" AND ("best practices" OR "bad practices" OR "implementation" OR "Culture") 77 9 17 5 108
+"chaos engineering" AND ("evaluation" OR "metrics") 85 25 28 4 142
+Total 766 217 404 94 1481
 3.3. Eligibility Criteria
-Following some existing MLR studies (Islam et al., 2019; Butijn et al., 2020; Garousi et al., 2019), we applied a specific set of inclusion and exclusion criteria, as detailed in Table  2
+Following some existing MLR studies (Islam et al., 2019; Butijn et al., 2020; Garousi et al., 2019), we applied a specific set of inclusion and exclusion criteria, as detailed in Table 4
 Table 4. Inclusion and Exclusion Criteria.
@@
 IC1: Articles must be in English and be readily accessible in full text. EC1: Short articles or brief papers (less than six pages).
-IC2: Articles must include robust validation methods, including working prototypes or tools with proper references, especially for grey sources.
-EC2: Articles that lack the relevant focus on chaos engineering.
+IC2: Articles must include clear validation or evaluation methods, with supporting references. EC2: Articles that lack the relevant focus on chaos engineering.
 IC3: Articles that report practices and challenges in implementing chaos engineering, including case studies, best practices, or guidelines.
@@
 The libraries’ advanced query modes were used to maximize search coverage.
-The search was conducted in April 2024, and the results were saved in BibTeX and CSV formats.
+The search was conducted in April 2024, and the search results were exported from each database in BibTeX and CSV formats for reference management and data processing.
 Duplicates were removed before proceeding with further analysis.
-Figure  3 illustrates the selection process for academic and grey literature, showing the databases searched and the number of selected papers at each stage.
+Figure  2 illustrates the selection process for academic and grey literature, showing the databases searched and the number of selected papers at each stage.
 Different approaches were followed to select grey and academic literature.
 3.4.1. Academic Literature Selection
-Searches in Scopus, IEEE Xplore, ACM DL, and Web of Science yielded 1,290 papers.
-After duplicate removal, 222 unique papers remained.
-Applying the inclusion and exclusion criteria, the selection was narrowed to 42 relevant papers.
-An additional search was conducted using Google Scholar, which added two more papers to the final count, bringing the total to 44 papers.
-Figure 3. MLR Study Selection Process.
+Searches in Scopus, IEEE Xplore, ACM DL, and Web of Science yielded 1,481 papers.
+After duplicate removal, 374 unique papers remained.
+Applying the inclusion and exclusion criteria, the selection was narrowed to 47 relevant papers.
+An additional search was conducted using Google Scholar, which identified two more articles, bringing the total to 49 papers.
+Figure 2.
+The MLR study selection process.
+Steps are shown sequentially for both academic and grey literature.
+After final screening, inter-rater reliability was assessed using Cohen’s Kappa (Landis and Koch, 1977), yielding $\kappa=0.826$ for academic and $\kappa=0.857$ for grey literature, indicating strong reviewer agreement.
 […]
@@
 Table 5. Criteria for Qualitative Assessment of Grey Literature.
-Criteria Questions
+Category Questions Criteria Score
 Authority (Q1). Is the organization publishing the work credible?
@@
 (Q3). How closely is the author connected to a reputable institution?
+2/3
 Methodology (Q4). Is the objective of the source evident and well-stated?
 (Q5). Can the author be considered an expert in this field?
-Date (Q6). Is the publication date of the source within the past five years?
-Table 4. Search Strings and Results.
-Search Strings ACM DL IEEE Xplore Scopus Web of Science Total
-"chaos engineering" OR "chaos test" OR "chaos experiment" OR "chaos toolkit" OR "chaos platform" OR "chaos mesh" OR "chaos integration" OR "chaos architecture" OR "chaos infrastructure" OR "chaos injection" OR "chaos tool" 95 63 121 47 326
-"chaos engineering" AND ("trends" OR "developments" OR "progression") 29 0 17 0 46
-"chaos engineering" AND ("leader" OR "pioneer" OR "expert" OR "contributor") 35 3 4 0 42
-"chaos engineering" AND ("phases" OR "stages" OR "steps" OR"pipeline") 62 5 11 0 78
-"chaos engineering" AND ("challenges" OR "issues" OR "problems") 79 16 34 4 133
-"chaos engineering" AND ("applications" OR "benefits") 77 20 37 8 142
-"chaos engineering" AND ("solutions" OR "approaches" OR "methods") 72 20 55 4 151
-"chaos engineering" AND ("tools" OR "techniques") 79 24 36 8 147
-"chaos engineering" AND ("best practices" OR "bad practices" OR "implementation" OR "Culture") 68 8 16 5 97
-"chaos engineering" AND ("evaluation" OR "metrics") 75 24 26 3 128
-Total 671 183 357 79 1290
+2/2
+Date (Q6). Is the publication date of the source within the past eight years? 1/1
 3.4.2. Grey Literature Selection
@@
 (Rahman et al., 2023).
-We incorporated additional quality assessment criteria beyond the inclusion and exclusion criteria, as outlined in Table  3.
-We picked six key criteria: the credibility of the publishing organization (Q1), the author’s prior publications in the field (Q2), the author’s institutional affiliation (Q3), clarity of purpose (Q4), the author’s expertise (Q5), and recency of the source (Q6).
-The responses to these criteria were classified as ‘Yes’ (1), ‘Partly’ (0.5), and ‘No’ (0), with scores assigned accordingly for each category.
-These criteria were adapted from prior studies on grey and multi-vocal literature reviews  (Soldani et al., 2018; Butijn et al., 2020; Kumara et al., 2021), ensuring a thorough evaluation of the sources.
-As a result, we selected 44 documents, contributing to 88 unique articles (Figure  3) for extracting and synthesizing data.
-Table  5 illustrates the selected sources’ distribution across various venues.
+We assessed the quality and relevance of the primary grey literature sources we obtained, as it cannot be assumed that their quality is guaranteed.
+The quality assessment criteria suggested by Garousi et al(Garousi and Mäntylä, 2016) have been used for this purpose (see Table 5).
+Those criteria comprise three quality categories, ranging from the authority of the producer to the publication date (see the first column of Table 5).
+These categories encompass six criteria (see the second column) that were assessed individually for each grey literature item.
+The evaluation was conducted independently by the first two authors of this study, who indicated whether each item (a) met or (b) failed to meet a criterion.
+In cases where a criterion could not be assessed, such as when author or affiliation details were missing, it was treated as not satisfied.
+The final column of Table 5 specifies how many of the criteria needed to be satisfied for inclusion.
+For example, an item had to meet at least 2 out of 3 criteria related to the authority of the source to be considered valid.
+To further ensure internal agreement and consistency in applying these quality checks, all disagreements between reviewers were resolved through discussion.
+As a result of this procedure, 47 grey literature sources were selected.
+These, together with 49 academic papers, resulted in a total of 96 sources used for data extraction and synthesis (Figure 2).
+The distribution of selected sources across venues is summarized in Table 6.
 Table 6. Selected Study for Data Synthesis and Qualitative Evaluation.
 Academic Literature Grey Literature
-Websites & Blogs Whitepaper
-(Fogli et al., 2023; Torkura et al., 2020; Zhang et al., 2019; Dedousis et al., 2023; Simonsson et al., 2021; Zhang et al., 2021a; Cotroneo et al., 2022; Camacho et al., 2022; Torkura et al., 2021; Poltronieri et al., 2022; Zhang et al., 2023; Nikolaidis et al., 2021; Frank et al., 2021b, a; Pierce et al., 2021; Soldani and Brogi, 2021; Al-Said Ahmad et al., 2024; Vu et al., 2022; Chen et al., 2024; Nikolaidis et al., 2023; Ma et al., 2023; Meiklejohn et al., 2021; Frank et al., 2023; Sondhi et al., 2021; Kassab et al., 2022; Jernberg et al., 2020; Konstantinou et al., 2021; Pulcinelli et al., 2023; Ji et al., 2023; Ikeuchi et al., 2020; Chen et al., 2022; Naqvi et al., 2022; Malik et al., 2023; Bedoya et al., 2023; Kesim et al., 2020; Basiri et al., 2019; Park et al., 2019; Klinaku et al., 2022; Rivera et al., 2023; Camilli et al., 2022; Yu et al., 2021; Siwach et al., 2022) = 44 (Inc., 2023; Wickramasinghe, 2023; Gill, 2021; Bremmers, 2021; Patel, 2022; Jakkaraju, 2020; Hawkins, 2020; Nombela, 2023; Le, 2022; National Australia Bank, 2020; Shah, 2021; Krishnamurthy, 2021; Moon, 2022; Katirtzis, 2022; Wachtel, 2023; Durai, 2022; Anwar and Arunachalam, 2019; Tomka, 2024; Dua, 2024; Bocetta, 2019; Mooney, 2023; Wu, 2021; Lardo, 2019; Miles, 2019a; Bhaskar, 2022; Azure Readiness, 2023; Gill, 2022; Fawcett, 2020; Manna, 2021; Gianchandani et al., 2022; Bairyev, 2023; Sbiai et al., 2023; Abdul, 2024; Roa, 2022; Kadikar, 2023; Green, 2023; Aharon, 2024; Krivas and Portela, 2020; Kostic, 2024; Vinisky, 2024; Colyer, 2019) = 41 (Palani and Gupta, 2023; Tivey et al., 2019; Starr et al., 2022) = 3
+Websites & Blogs White Papers
+(Fogli et al., 2023; Torkura et al., 2020; Zhang et al., 2019; Dedousis et al., 2023; Simonsson et al., 2021; Zhang et al., 2021a; Cotroneo et al., 2022; Camacho et al., 2022; Torkura et al., 2021; Poltronieri et al., 2022; Zhang et al., 2023; Nikolaidis et al., 2021; Frank et al., 2021b, a; Pierce et al., 2021; Soldani and Brogi, 2021; Al-Said Ahmad et al., 2024; Vu et al., 2022; Chen et al., 2024; Nikolaidis et al., 2023; Ma et al., 2023; Meiklejohn et al., 2021; Frank et al., 2023; Sondhi et al., 2021; Kassab et al., 2022; Jernberg et al., 2020; Konstantinou et al., 2021; Pulcinelli et al., 2023; Ji et al., 2023; Ikeuchi et al., 2020; Chen et al., 2022; Naqvi et al., 2022; Malik et al., 2023; Bedoya et al., 2023; Kesim et al., 2020; Basiri et al., 2019; Park et al., 2019; Klinaku et al., 2022; Rivera et al., 2023; Camilli et al., 2022; Yu et al., 2021; Siwach et al., 2022; Basiri et al., 2016; Tucker et al., 2018; Alvaro and Tymon, 2017; Sousa et al., 2018; Alvaro et al., 2016) = 49 (Inc., 2023; Wickramasinghe, 2023; Gill, 2021; Bremmers, 2021; Patel, 2022; Jakkaraju, 2020; Hawkins, 2020; Nombela, 2023; Le, 2022; National Australia Bank, 2020; Shah, 2021; Krishnamurthy, 2021; Moon, 2022; Katirtzis, 2022; Wachtel, 2023; Durai, 2022; Anwar and Arunachalam, 2019; Tomka, 2024; Dua, 2024; Bocetta, 2019; Mooney, 2023; Wu, 2021; Lardo, 2019; Miles, 2019a; Bhaskar, 2022; Azure Readiness, 2023; Gill, 2022; Fawcett, 2020; Manna, 2021; Gianchandani et al., 2022; Bairyev, 2023; Sbiai et al., 2023; Abdul, 2024; Roa, 2022; Kadikar, 2023; Green, 2023; Aharon, 2024; Krivas and Portela, 2020; Kostic, 2024; Vinisky, 2024; Colyer, 2019; Hochstein and Rosenthal, 2016; Butow, 2018; Wilms, 2018) = 44 (Palani and Gupta, 2023a; Tivey et al., 2019; Starr et al., 2022) = 3
 3.5. Data Extraction, Synthesis, and Analysis
@@
 We applied thematic analysis to synthesize insights across our research questions, ensuring that coding remained objective and independent of the data extraction structure or biases (Stewart, 2024).
-Full papers were subsequently imported into ATLAS.ti, a qualitative data analysis tool that allowed for direct text coding.
+The full papers were subsequently imported into ATLAS.ti  [^1] https://atlasti.com/, a qualitative data analysis tool that allowed for direct text coding.
 This process adhered to established thematic analysis steps (Braun and Clarke, 2006), providing consistency and depth in our approach.
@@
   • Familiarization with the Data: We began by reading the selected papers to understand the data comprehensively, enabling us to identify the key elements relevant to our research questions.
-  • Pilot Study: We randomly selected twenty sources for initial analysis, with the first two authors of the paper independently analyzing each to establish preliminary codes using structural and descriptive coding approaches (Saldaña, 2021).
-This phase involved creating codes by conceptualizing information related to the research questions.
-Through iterative review and refinement, researchers adjusted codes to clarify emerging theories and form thematic clusters.
-  • Inter-Rater Consensus Analysis for the Pilot Study: After finalizing the coding process, the first two authors of the paper conducted an inter-rater review, assessing each other’s codes to establish consensus on naming, types, and categorization.
-This collaborative effort helped to refine codes and maintain their consistency.
-  • Complete Dataset Coding: The remaining sources were subsequently coded based on the agreement reached during the pilot study.
-The first author of this paper conducted this step, and the second author reviewed the resulting codes.
-  • Inter-Rater Consensus Analysis for the Complete Study: After completing the individual coding, we performed a final inter-rater assessment to ensure the coding aligned with the established framework.
-Any discrepancies were discussed and resolved.
-  • Generating the report: The results of this thematic analysis are presented in Sections 3,  4,  5,  6 and  7.
+  • Pilot Study: We randomly selected twenty sources for initial analysis, with the first author applying structural and descriptive coding (Saldaña, 2021) to conceptualize data relevant to the research questions (Kumara et al., 2021).
+The process began by extracting quotations—direct statements from the primary studies—and identifying keywords that captured their core meaning.
+For example (see Figure 3), from quotations referring to the need to introduce failure scenarios and monitor experiment duration, the keywords Introduce failure scenarios and Experiment duration were extracted.
+These were then grouped to form the code Experiment Execution.
+Similarly, the keywords Monitoring sidecar and Observability for diagnosis were extracted and abstracted into the code Monitoring and Observability.
+These codes were clustered into the broader theme Functional Requirements.
+This phase yielded 536 initial codes and an initial set of 28 candidate themes, which served as the foundation for subsequent refinement.
+  • Inter-Rater Consensus Analysis for the Pilot Study: After completing the pilot coding, the second author reviewed the full set of codes independently, assessing the naming, scope, and categorization of each code.
+The discrepancies between authors were discussed and resolved collaboratively.
+This consensus process refined the initial 536 codes to a validated set of 421 codes, improving the clarity and consistency of the thematic structure.
+  • Complete Dataset Coding: The remaining sources were subsequently coded using the validated codebook.
+The first author conducted the full-dataset coding, and the second author reviewed the outputs.
+Discrepancies identified during this phase were resolved collaboratively.
+  • Inter-Rater Consensus Analysis for the Complete Study: After the complete coding process, a final inter-rater assessment was conducted to ensure consistent application of the codes across the dataset.
+The discrepancies in interpretation were discussed and resolved.
+During this phase, the initial 28 themes were critically reviewed for internal coherence and conceptual distinctiveness.
+Overlapped or weakly supported themes were consolidated or removed, resulting in a final set of 15 distinct, non-overlapping themes.
+At this stage, each finalized theme was explicitly mapped to the specific research question(s) it addressed (e.g., RQ1.1 – Functional Requirements, RQ3.2 – Best Practices), ensuring that the thematic framework remained purposeful and traceable throughout the analysis.
+  • Generating the report: The results of this thematic analysis are presented in Sections 4,  5,  6,  7 and  8.
+Figure 3.
+A representative output from our thematic analysis process, showing how quotations from academic and grey literature were distilled into keywords, abstracted into codes, and grouped under broader themes.
+( Refer to caption: grey literature; Refer to caption: academic literature )
+[…]
+Illustration of qualitative coding and thematic abstraction from quotes to themes.
 3.6. Replication Package
-For the validation and reproducibility of this work, we have provided an extensive replication package online [^1] https://drive.google.com/drive/folders/1wOu11lB9wsxK7cQKzPL63GAjD9UiiF0s?usp=sharing.
-This package includes the complete compilation of sources, search queries, and thematic analysis performed using Atlas.ti [^2] https://atlasti.com/ (such as codes, groupings, quotations, and analyses).
+For the validation and reproducibility of this work, we have provided an extensive replication package online  [^2] https://doi.org/10.5281/zenodo.15696868.
+This includes the complete compilation of sources, search queries, and thematic analysis performed using Atlas.ti (such as codes, groupings, quotations, and analyses), as well as the associated code tables.
 4. Chaos Engineering: Definitions, Functionaries, and Elements (RQ1)
-This section first examines practitioners’ definitions of chaos engineering.
-Next, it describes the phases, functional requirements, and non-functional requirements of chaos engineering.
-Finally, it identifies the main functional components of chaos engineering.
+This section first reviews how chaos engineering is defined across academic and industry sources.
+Next, it outlines its core activities and commonly associated functionalities (RQ1.1).
+Finally, it introduces the essential components (RQ1.2) and quality requirements (RQ1.3) that form the foundation for implementing effective chaos engineering platforms.
 4.1. Definitions
-Chaos engineering is a growing practice, but its definition remains inconsistent.
-Hence, we aim to establish a clear and shared definition by analyzing those proposed by different sources.
-Originally developed at Netflix, chaos engineering focuses on intentionally introducing faults within a system to assess and enhance its robustness  (Fogli et al., 2023; Torkura et al., 2020; Dedousis et al., 2023; Torkura et al., 2021).
-However, confusion often arises when distinguishing chaos engineering from traditional testing practices.
-While conventional testing remains vital, chaos engineering should be seen as a complement rather than a replacement (Green, 2023).
-Traditional testing and monitoring are crucial, yet they may not be sufficient to expose all possible failure modes in complex systems.
-Chaos engineering mitigates this issue by proactively identifying the weaknesses of a system that might go unnoticed during standard testing procedures but may surface as failures in production environments (Paul, 2019; Green, 2023).
-Table 6. Selected Definitions of Chaos Engineering
-Source Definition
-Netflix
-"the discipline of experimenting on a system to build confidence in its ability to withstand turbulent
-conditions in production" (Team, 2024).
-Gremlin
-"a disciplined approach to identifying failures before they become outages by intentionally
-’breaking things’ to build more resilient systems" (Inc., 2023).
-AWS
-"the process of stressing an application in testing or production environments by creating
-disruptive events, such as server outages or API throttling, observing how the system responds,
-and implementing improvements" (Starr et al., 2022; Kahil and Arenhage, 2021).
-Google "the practice of intentionally introducing failures into production systems under load" (Google Cloud, 2024).
-Microsoft "intentionally injecting faults into applications in a controlled manner to simulate real-world issues" (Microsoft, 2023).
-IBM
-"intentionally causing failures in production or pre-production environments
-to understand their impact and improve defense strategies" (IBM, 2024).
-Datadog
-"a discipline that encourages experimenting in production and injecting controlled failures
-into the system to understand how the system will react and improve reliability" (Pulido et al., 2021).
-Harness "the science of injecting failure to gauge system resiliency" (Schillerstrom, 2022).
-Twilio "intentionally injecting failures to test a system’s ability to self-heal" (Wong, 2017).
-Uber "the practice of injecting controlled failures directly into production environments to test resilience" (Jones, 2020).
-Splunk
-"the practice of assessing the resilience of production systems by intentionally introducing
-failures and running experiments" (Wickramasinghe, 2023).
-Figure 4. Word Cloud of Key Terms in Chaos Engineering Definitions
-Table 6 shows the excerpts from the definitions provided by a subset of the reviewed literature.
-We also created a word cloud to identify the key terms used in these definitions, as shown in Figure  4.
-The key themes emphasized in these definitions include system resilience, failure injection, testing in production environments, and controlled experiments.
+Chaos engineering has garnered significant attention (Monroe, 2019) in both academia and industry, yet its definition remains fragmented and inconsistent.
+To address this, we analyzed a set of definitions identified during our thematic analysis process.
+These include representative definitions from each, summarized in Tables 7 and 8.
+Across industry and academic sources, several key elements consistently emerge: system resilience, failure injection, controlled experiments, and testing in production or production-like environments.
 These recurring elements highlight chaos engineering’s focus on deliberately introducing failures in a controlled manner to assess how systems behave under stress and recover from disruptions.
-Based on these insights, we propose the following unified definition.
+Table 7. Practitioners’ Definitions of Chaos Engineering
+Definition Articles Referenced
+"The practice of deliberately breaking systems in controlled ways to identify potential failure points before they occur naturally." (Tomka, 2024; Palani and Gupta, 2023b; Abdul, 2024; Green, 2023; Krivas and Portela, 2020; Vinisky, 2024)
+"The methodology of creating disruptive events and stressing applications to observe system responses and implement targeted improvements." (Gill, 2021; Bhaskar, 2022; Bairyev, 2023; Kadikar, 2023)
+"The technique of gauging system resiliency by intentionally causing failures under load to test self-healing mechanisms." (Patel, 2022; Gill, 2022; Bairyev, 2023; Kadikar, 2023; Kostic, 2024)
+Table 8. Academic Definitions of Chaos Engineering
+Definition Articles Referenced
+"A disciplined practice of experimenting on a software or distributed system to build confidence in its ability to withstand turbulent or unexpected conditions in production." (Fogli et al., 2023; Torkura et al., 2020; Pierce et al., 2021; Al-Said Ahmad et al., 2024; Vu et al., 2022; Jernberg et al., 2020; Naqvi et al., 2022; Vinisky, 2024; Basiri et al., 2016; Sousa et al., 2018; Alvaro et al., 2016)
+"An empirical and scientific approach for learning how software systems behave under stress by conducting controlled experiments to evaluate and improve their resilience and fault-handling mechanisms." (Zhang et al., 2019; Sondhi et al., 2021; Ikeuchi et al., 2020; Malik et al., 2023)
+"A resilience-testing approach involving the intentional injection of failures or disruptions into a system—especially in production or production-like environments—to uncover weaknesses and improve fault tolerance." (Poltronieri et al., 2022; Ji et al., 2023; Kesim et al., 2020; Konstantinou et al., 2021)
+We followed the synthesis method proposed by Gong and Ribiere (Gong and Ribiere, 2021), drawing on conceptual evaluation criteria from Wacker (Wacker, 2004) and Suddaby (Suddaby, 2010) (see our online appendix Section 3.6 for details).
+A total of 48 definitions were analyzed, 27 from academic sources and 29 from grey literature.
+Based on this evaluation, 34 definitions were retained and 22 were excluded due to issues such as vagueness, conceptual overlap, or lack of definitional clarity.
+These inconsistencies highlight the difficulty of establishing a unified understanding of chaos engineering—a challenge observed in previous research, where fragmented definitions can hinder conceptual clarity and slow down the development of standardized practices (Gong and Ribiere, 2021; Degefa et al., 2021; Diaz-De-Arcaya et al., 2024; Awasthi et al., 2025).
+As demonstrated in those contexts, establishing a clear and shared definition can help eliminate conceptual ambiguity, enable the development of sound theory and valid constructs, and strengthen the connection between research and practice.
+Based on these insights, we propose the following unified definition:
 Chaos engineering is a resilience testing practice that intentionally injects controlled faults into software systems in production-like or actual production environments to simulate adverse real-world conditions.
@@
 It also enables the assessment and enhancement of systems’ capacity and operational readiness against such disruptions.
-In line with the definition, chaos engineering relies on several guiding principles to design and carry out experiments, ensuring that systems are rigorously tested under controlled conditions (Basiri et al., 2016; Aisosa, 2023).
-Table 7 summarizes these principles.
-Table 7. Principles and Measurement Metrics
-Principle Description Articles Referenced
+In line with the definition, chaos engineering relies on a set of core activities to guide the design and execution of experiments under controlled conditions (Basiri et al., 2016; Aisosa, 2023).
+Table 9 summarizes these activities.
+Table 9. Core Activities in the Chaos Engineering Lifecycle
+Activity Description Articles Referenced
 Establish Steady State Understand the system’s normal behavior and hypothesize how it will handle chaos (real-world disruptive incidents/failures).
-(Fogli et al., 2023; Torkura et al., 2020; Zhang et al., 2019; Dedousis et al., 2023; Simonsson et al., 2021; Inc., 2023; Bremmers, 2021; Gill, 2021; Wickramasinghe, 2023; Cotroneo et al., 2022; Nikolaidis et al., 2021; Frank et al., 2021b; Pierce et al., 2021; Frank et al., 2023; Konstantinou et al., 2021; Naqvi et al., 2022; Malik et al., 2023; Bedoya et al., 2023; Kesim et al., 2020; Basiri et al., 2019; Zhang et al., 2023; Siwach et al., 2022; Jernberg et al., 2020; Jakkaraju, 2020; Bocetta, 2019; Palani and Gupta, 2023; Gill, 2022; Fawcett, 2020; Manna, 2021; Bairyev, 2023; Sbiai et al., 2023; Green, 2023; Krivas and Portela, 2020; Kostic, 2024; Abdul, 2024; Katirtzis, 2022; Le, 2022; Colyer, 2019; Patel, 2022)
+(Fogli et al., 2023; Torkura et al., 2020; Zhang et al., 2019; Dedousis et al., 2023; Simonsson et al., 2021; Inc., 2023; Bremmers, 2021; Gill, 2021; Wickramasinghe, 2023; Cotroneo et al., 2022; Nikolaidis et al., 2021; Frank et al., 2021b; Pierce et al., 2021; Frank et al., 2023; Konstantinou et al., 2021; Naqvi et al., 2022; Malik et al., 2023; Bedoya et al., 2023; Kesim et al., 2020; Basiri et al., 2019; Zhang et al., 2023; Siwach et al., 2022; Jernberg et al., 2020; Jakkaraju, 2020; Bocetta, 2019; Palani and Gupta, 2023a; Gill, 2022; Fawcett, 2020; Manna, 2021; Bairyev, 2023; Sbiai et al., 2023; Green, 2023; Krivas and Portela, 2020; Kostic, 2024; Abdul, 2024; Katirtzis, 2022; Le, 2022; Colyer, 2019; Patel, 2022)
 Setup Monitoring Infrastructure Set up tools to monitor the system for deviations from its normal behavior during controlled experiments.
@@
 Conduct Chaos Experiments Design and execute experiments to simulate real-world failures.
-(Fogli et al., 2023; Torkura et al., 2020; Zhang et al., 2019; Dedousis et al., 2023; Simonsson et al., 2021; Inc., 2023; Bremmers, 2021; Gill, 2021; Pierce et al., 2021; Naqvi et al., 2022; Malik et al., 2023; Bedoya et al., 2023; Kesim et al., 2020; Jernberg et al., 2020; Jakkaraju, 2020; Palani and Gupta, 2023; Gill, 2022; Fawcett, 2020; Manna, 2021; Bairyev, 2023; Patel, 2022; Abdul, 2024; Roa, 2022)
+(Fogli et al., 2023; Torkura et al., 2020; Zhang et al., 2019; Dedousis et al., 2023; Simonsson et al., 2021; Inc., 2023; Bremmers, 2021; Gill, 2021; Pierce et al., 2021; Naqvi et al., 2022; Malik et al., 2023; Bedoya et al., 2023; Kesim et al., 2020; Jernberg et al., 2020; Jakkaraju, 2020; Palani and Gupta, 2023a; Gill, 2022; Fawcett, 2020; Manna, 2021; Bairyev, 2023; Patel, 2022; Abdul, 2024; Roa, 2022)
 Test and Refine Review results, apply improvements, and repeat to boost resilience.
@@
 Expand to Production Run experiments in production to build confidence in the system’s ability to endure adverse real-world events.
-(Nombela, 2023; Krishnamurthy, 2021; Moon, 2022; Simonsson et al., 2021; Pierce et al., 2021; Konstantinou et al., 2021; Park et al., 2019; Fawcett, 2020)
-3.2. Functionalities of Chaos Engineering
-Based on our review of academic and grey literature, we identified five key features that define an effective chaos engineering solution: experimentation hub, collaborative workflow management, observability suite, security and governance, and integration and automation (see Figure  5).
-The rest of this section discusses each in detail.
-3.2.1. Experimentation Hub
-Experimentation is a crucial aspect of a chaos engineering platform, allowing users to meticulously plan and seamlessly conduct chaos experiments (Katirtzis, 2022; Abdul, 2024; Krivas and Portela, 2020).
-In this phase, users first define the scope, targets, and parameters of chaos experiments with precision, ensuring a clear focus for testing (Kostic, 2024; Bremmers, 2021; Kadikar, 2023; Wu, 2021; Cotroneo et al., 2022; Katirtzis, 2022; Torkura et al., 2020).
-The next step is to establish the steady state of the system - baseline performance metrics and normal behavior - which serves as a reference point for evaluating the impact of disruptions (Frank et al., 2023; Konstantinou et al., 2021; Jernberg et al., 2020; Siwach et al., 2022).
-Based on this steady state, users form hypotheses about how the system will behave under stress, guiding both the design of experiments and the setting of resilience expectations  (Kostic, 2024; Zhang et al., 2019; Wickramasinghe, 2023; Fogli et al., 2023; Dedousis et al., 2023).
-Once the experiment is defined, the chaos engineering platform injects faults, such as network latency, packet loss, or server failures.
-This enables users to assess the weaknesses and operational readiness of a system by evaluating its fault handling capabilities (Kostic, 2024; Chen et al., 2024; Fogli et al., 2023; Zhang et al., 2023).
-As these faults are injected and system behavior is observed, careful planning and execution help chaos experiments effectively identify and address system vulnerabilities, ultimately strengthening overall resilience (Nikolaidis et al., 2021; Colyer, 2019).
-3.2.2. Collaborative Workflow Management
-Chaos engineering is a collaborative practice that brings together stakeholders such as developers, testers, operations teams, and security experts.
-These individuals must work closely to design, execute, and analyze experiments to identify system vulnerabilities.
-Through effective collaborative workflow management, teams can maintain clear communication, coordinate tasks efficiently, and ensure that all parties are informed throughout the experimentation process (Siwach et al., 2022; Dua, 2024; Green, 2023).
-Facilitating Team Communication and Coordination.
-Chaos engineering relies on continuous communication between relevant parties to ensure alignment with the objectives of the experiment and the expected results (Gill, 2021).
-By forming cross-functional working groups, participants can discuss scenarios, hypothesize outcomes, and maintain a shared understanding of the objectives of the experiment.
-This approach ensures that everyone stays coordinated and clear on their roles during the execution and analysis phases.
-For example, in a simulated database failure, developers can focus on diagnosing and resolving the issue while the operations group works to reroute traffic and maintain service availability.
-At the same time, security experts assess the potential vulnerabilities that the failure exposes, and testers verify the system response to failures (Patel, 2022).
-This collaborative process helps avoid information silos and supports collective learning, leading to better preparation for future incidents (Green, 2023; Kadikar, 2023).
-Notify Stakeholders about Experiment Progress.
-Keeping stakeholders informed about the status of the chaos engineering experiment is crucial to transparency and effective decision making.
-First, well-designed notification systems trigger alerts based on anomalies, such as unexpected error rates or system performance changes  (Bremmers, 2021).
-Then, these alerts are automatically sent to stakeholders through integrated notification features.
-These alerts enable quick issue communication and prompt responses to performance degradation or system instability.
-In addition, they provide real-time insight into system behaviors, facilitating adjustments during the experiment.
-Finally, this continuous feedback loop refines the monitoring systems and improves incident preparedness through improved alert accuracy (Shah, 2021).
-Maintain Documentation and Knowledge Repositories.
-Building a consistent knowledge base for future reference is essential for the long-term success and continuous improvement of chaos engineering practices (Vinisky, 2024; Kadikar, 2023).
-By maintaining detailed records of experiment scenarios, findings, and recommendations, organizations create a valuable foundation for troubleshooting and informed decision-making (Kadikar, 2023; Gill, 2022).
-This process should include standardized templates for experiments and comprehensive postmortem reports, ensuring that the documentation is consistent and thorough (Siwach et al., 2022).
-A well-organized knowledge base preserves essential information and fosters a deeper understanding of system behavior under stress  (Gill, 2021).
-These insights from previous experiments can directly improve system design and operational practices (Vinisky, 2024).
-Through consistent documentation and structured knowledge sharing, chaos engineering platforms ensure that lessons learned are preserved, actively shaping future experiments, and strengthening overall system performance (Torkura et al., 2020).
-3.2.3. Observability Suite
-Chaos engineering platforms use an observability suite to monitor and analyze systems during experiments, allowing teams to detect issues early, identify root causes, and take corrective actions (Green, 2023; National Australia Bank, 2020; Chen et al., 2022).
-Key features discussed in the following subsections include real-time dashboards and experiment impact analysis to identify faults’ root causes.
-Provide Real-Time Dashboards.
-Real-time observability depends on dashboards, metrics, traces, and logs for the visualization of experiments.
-(National Australia Bank, 2020; Gatev, 2021; Pulcinelli et al., 2023).
-First, dashboards, integrated with monitoring tools, can show critical metrics such as latency, throughput, error rates, and resource usage, allowing engineers to monitor system performance and quickly detect problems (Wickramasinghe, 2023; Manna, 2021; Fawcett, 2020; Green, 2023).
-Compared to performance metrics, traces can capture the flow of requests through the system, revealing interactions among components and identifying bottlenecks (Chen et al., 2024; Zhang et al., 2021a).
-Distributed tracing tools allow engineers to analyze component interactions and inefficiencies during chaos experiments (Konstantinou et al., 2021; Frank et al., 2021b).
-Finally, logs record system events, supporting diagnostic efforts, especially in security contexts (Klinaku et al., 2022; Chen et al., 2022; Green, 2023).
-Log analysis tools enable engineers to trace anomalies and maintain a complete audit trail  (Torkura et al., 2021; Soldani and Brogi, 2021; Vu et al., 2022).
-Analyze Impact and Identify Root Causes of Faults.
-After gathering and visualizing data through real-time dashboards, the next crucial step is to analyze the results of chaos experiments and identify the root causes of any faults, using each pillar of observability in the process  (Yu et al., 2021; Gianchandani et al., 2022) The metrics collected during chaos experiments are compared with baseline performance to detect deviations that may indicate vulnerabilities.
-These deviations help identify how well the system handled stress and where failures occurred (Ji et al., 2023).
-Logs and trace data provide deeper insight into how individual components interacted and responded to stress.
-This detailed data helps identify the source of failures and component interactions that contribute to performance problems (Green, 2023).
-These metrics, logs, and traces can be combined to facilitate root cause analysis and offer qualitative and quantitative insights.
-For example, this helps engineers determine whether failures were caused by bottlenecks, resource mismanagement, or faulty component interactions (Yu et al., 2021).
-By understanding the root causes of failures, teams can implement targeted improvements to enhance and prevent future failures (Gill, 2022; Nombela, 2023; Gianchandani et al., 2022).
-3.2.4. Security and Governance
-Chaos engineering platforms require a robust security and governance framework to ensure that experiments are conducted safely and comply with regulatory standards (Fogli et al., 2023; Kostic, 2024).
-This section outlines vital mechanisms that protect system integrity while adhering to industry regulations, focusing on implementing safety controls, automated rollbacks, and ensuring regulatory compliance.
-Implement Safety Controls and Automated Rollbacks.
-Our literature identifies several vital mechanisms, including abort conditions, automated rollbacks, and blast radius control, which ensure that experiments are conducted safely (Patel, 2022).
-  • Abort conditions: These ensure that chaos experiments are immediately stopped if predefined thresholds are exceeded (Abdul, 2024; Sbiai et al., 2023).
-For example, the experiment is automatically terminated if performance metrics, such as latency or availability, degrade beyond acceptable levels.
-This prevents further damage and maintains system stability by stopping experiments before they can cause widespread problems (Torkura et al., 2020).
-  • Automated rollbacks: These allow systems to revert to a known stable state after an experiment (Fawcett, 2020; Miles, 2019b).
-In cases where a simulated failure causes service disruption, automated rollback mechanisms quickly restore the system to its pre-experiment configuration.
-This reduces downtime and prevents prolonged instability in the system (Gill, 2022).
-  • Blast radius control: This mechanism limits the scope of chaos experiments to isolated parts of the system (Wickramasinghe, 2023; Manna, 2021).
-For example, teams might start by targeting a single microservice instead of testing the entire infrastructure.
-Gradually expanding the blast radius allows controlled testing, ensuring vulnerabilities can be identified without risking widespread system failures (Abdul, 2024).
-Additionally, Gremlin  (Inc., 2023) recommends using the Rumsfeld matrix, a framework to categorize system behaviors during chaos experiments.
-This approach enables teams to understand better and anticipate various types of system responses by classifying them as follows:
+(Nombela, 2023; Krishnamurthy, 2021; Moon, 2022; Simonsson et al., 2021; Pierce et al., 2021; Konstantinou et al., 2021; Park et al., 2019; Fawcett, 2020; Willard and Hutson, 2024; Tucker et al., 2018; Basiri et al., 2016; Alvaro et al., 2016)
+4.2. Functionalities of Chaos Engineering (RQ1.1)
+Chaos engineering comprises a structured set of core functionalities designed to reveal system weaknesses under real-world failure conditions.
+Based on our thematic analysis of academic and grey literature (see Section 3.5), we identified five key functionalities that characterize how chaos engineering is typically conducted: experiment design and planning, fault execution and control, monitoring and observability, post-experiment analysis, and automation and continuous integration (see Figure 4).
+For full transparency, the detailed mapping of raw codes to these functionalities and supporting themes is provided in the online appendix (see Section 3.6).
+The remainder of this section describes each functionality, emphasizing its purpose and role in validating system resilience.
+Figure 4. Key Functionalities Provided by Chaos Engineering.
+[…]
+Functional Requirement
+4.2.1. Experiment Design and Planning
+This phase establishes the conceptual and procedural foundation for chaos experimentation.
+It ensures that testing is deliberate, measurable, and risk-aware (Fogli et al., 2023; Simonsson et al., 2021; Patel, 2022; Vu et al., 2022).
+It involves three essential activities: identifying test targets and failure scenarios, defining steady-state behavior, and formulating testable hypotheses.
+Identify Test Targets and Failure Scenarios
+The first step is to identify components whose failure could significantly disrupt operations, such as microservices, APIs, databases, or message queues (Patel, 2022; Naqvi et al., 2022; Vu et al., 2022; Bremmers, 2021).
+Failure scenarios are then defined based on system architecture analysis (e.g., service dependencies, communication patterns), operational risks, or past incident reports (Fogli et al., 2023; Chen et al., 2024; Malik et al., 2023; Nikolaidis et al., 2023).
+Typical disruptions include network latency, packet loss, service crashes, or CPU throttling (Simonsson et al., 2021; Chen et al., 2024).
+These scenarios are prioritized according to their likelihood and potential business impact (Fogli et al., 2023; Naqvi et al., 2022; Siwach et al., 2022; Torkura et al., 2020).
+Define Steady-State Behavior
+Steady-state refers to the normal, healthy performance of the system.
+Teams define this using measurable indicators such as latency, error rates, throughput, or availability (Frank et al., 2023; Malik et al., 2023; Simonsson et al., 2021).
+These metrics act as baselines against which system behavior is compared during and after failure injection (Naqvi et al., 2022; Siwach et al., 2022; Gill, 2021).
+Establishing these baselines enables reliable detection of performance degradation and validation of resilience mechanisms (Naqvi et al., 2022; Inc., 2023).
+Formulate Testable Hypotheses
+Hypotheses describe the expected system response to injected faults.
+For instance, a hypothesis might state: “If service A fails, traffic will reroute to service B within 200ms without user impact” (Dedousis et al., 2023; Simonsson et al., 2021).
+These statements guide the design of experiments and help interpret the results.
+To manage uncertainty and uncover hidden risks, teams can use structured frameworks such as the Rumsfeld Matrix, which categorizes system knowledge as follows (Inc., 2023; Bocetta, 2019; Dedousis et al., 2023; Pierce et al., 2021; Malik et al., 2023; Palani and Gupta, 2023b).
   • Known Knowns: Aspects of the system that are fully understood and predictable.
@@
   • Unknown Unknowns: Completely unforeseen issues that may emerge during experimentation.
-Starting with known knowns, teams establish a baseline of predictable system behavior.
-They can then progressively move into more uncertain areas like Known Unknowns and Unknowns Knowns before finally addressing the most unpredictable Unknown Unknowns (Palani and Gupta, 2023).
-This step-by-step approach reduces system risks by building understanding before testing less predictable areas.
-Ensure Regulatory Compliance.
-Adhering to industry standards and legal frameworks is essential to ensure that systems meet established data protection and security requirements, including guidelines such as the General Data Protection Regulation (GDPR) and IEC 62443 (Helmke, 2019; Torkura et al., 2021).
-These frameworks provide organizations with critical guidance to safeguard sensitive information, maintain operational integrity, and protect vital infrastructures (Helmke, 2019; Kostic, 2024).
-To meet these compliance requirements, some organizations adopt Infrastructure as Code (IaC), which automates the provisioning and management of infrastructure (Torkura et al., 2020; Fogli et al., 2023).
-Tools such as Terraform and Ansible are commonly used in IaC to ensure consistent implementation of security policies, minimize human error, and maintain infrastructure in alignment with regulatory standards (Torkura et al., 2021; Vu et al., 2022).
-In addition to IaC, monitoring tools such as Datadog  (Abdul, 2024; Mooney, 2023), which detects compliance violations and configuration drift such as publicly exposed data, and Security Monkey, which monitors Secure Sockets Layer (SSL) certificates and other security settings to prevent vulnerabilities such as certificate expiration, are essential to maintain compliance (Dua, 2024; Tivey et al., 2019).
-3.2.5. Integration and Automation with Continuous Integration and Continuous Delivery (CI/CD)
-CI/CD pipelines are critical to optimizing the efficiency and consistency of chaos experiments (Rivera et al., 2023; Bairyev, 2023; Azure Readiness, 2023).
-When performed manually, tasks such as performing experiments and monitoring system behavior can be inefficient and prone to human error (Kostic, 2024; Jernberg et al., 2020; Bremmers, 2021; Chen et al., 2024).
-Automating these processes minimizes human intervention, ensuring tasks are performed reliably and consistently, thus reducing the likelihood of mistakes (Wickramasinghe, 2023; Camacho et al., 2022; Simonsson et al., 2021; Green, 2023).
-Integrating chaos experiments into CI/CD pipelines further enhances this reliability by automatically initiating tests as part of the routine development workflow (Roa, 2022).
-With each new deployment or configuration update, chaos tests run alongside other routine tests, enabling early detection of vulnerabilities.
-This proactive approach helps ensure issues are addressed in the development phase, resulting in a more comprehensive and practical testing process (Bremmers, 2021; Manna, 2021)
-Figure 5. Key Functionalities Provided by Chaos Engineering.
-[…]
-Functional Requirement
-3.3. Quality Requirements for Chaos Engineering Platform
-We have identified 12 quality requirements for a chaos engineering platform from the selected literature: performance, recoverability, resilience, timeliness, observability, scalability, effectiveness, robustness, accuracy, fault tolerance, cost, and adaptability.
-We provide these requirements and the corresponding metrics for measuring them in Table 8.
-Table 8. Quality Attributes and Measurement Metrics
-Quality Attribute Measurement Metrics Articles Referenced Total
-Performance Evaluate resource use and performance under attack, ensuring chaos experiments align with production metrics.
-(Fogli et al., 2023; Torkura et al., 2020; Zhang et al., 2019; Dedousis et al., 2023; Patel, 2022; Gill, 2021; Cotroneo et al., 2022; Nikolaidis et al., 2021; Vu et al., 2022; Klinaku et al., 2022) 10
-Recoverability Measure mean time to repair (MTTR), restoration speed, and detection response.
-(Fogli et al., 2023; Torkura et al., 2020; Dedousis et al., 2023; Patel, 2022; Gill, 2021; Cotroneo et al., 2022; Camacho et al., 2022; Zhang et al., 2021a; Nikolaidis et al., 2021; Pierce et al., 2021; Vu et al., 2022; Frank et al., 2021b; Malik et al., 2023; Zhang et al., 2023; Tomka, 2024; Kadikar, 2023) 16
-Resilience Tolerance for degraded performance under stress; success rate of automated restoration actions.
-(Fogli et al., 2023; Torkura et al., 2020; Zhang et al., 2019; Dedousis et al., 2023; Inc., 2023; Bremmers, 2021; Zhang et al., 2021a; Al-Said Ahmad et al., 2024; Vu et al., 2022; Nikolaidis et al., 2023; Frank et al., 2023; Kassab et al., 2022; Konstantinou et al., 2021; Nombela, 2023; Krishnamurthy, 2021; Bocetta, 2019; Manna, 2021; Kostic, 2024; Krivas and Portela, 2020; National Australia Bank, 2020; Azure Readiness, 2023) 20
-Timeliness Time to restore service and detect interruptions; impact on user experience and disruption.
-(Camacho et al., 2022; Pierce et al., 2021; Vu et al., 2022; Frank et al., 2021b; Malik et al., 2023; Cotroneo et al., 2022; Basiri et al., 2019; Gill, 2021; Zhang et al., 2019; Durai, 2022; Shah, 2021; Katirtzis, 2022; Dua, 2024; Wu, 2021; Palani and Gupta, 2023; Manna, 2021) 16
-Observability Assess logging completeness and monitoring uptime; measure alert latency and tracing success rate.
-(Fogli et al., 2023; Torkura et al., 2020; Dedousis et al., 2023; Inc., 2023; Bremmers, 2021; Patel, 2022; Gill, 2021; Camacho et al., 2022; Torkura et al., 2021; Zhang et al., 2021a; Vu et al., 2022; Frank et al., 2023; Malik et al., 2023; Kesim et al., 2020; Siwach et al., 2022; Konstantinou et al., 2021; Hawkins, 2020; Manna, 2021; Kostic, 2024; Kadikar, 2023; Azure Readiness, 2023) 20
-Scalability Manage increasing attack rates and scale experiment complexity.
-(Fogli et al., 2023; Torkura et al., 2020; Zhang et al., 2019; Patel, 2022; Gill, 2021; Vu et al., 2022; Sondhi et al., 2021; Malik et al., 2023; Klinaku et al., 2022; Camilli et al., 2022; Nikolaidis et al., 2023; Kassab et al., 2022; Rivera et al., 2023; Krishnamurthy, 2021; Durai, 2022; Wu, 2021; Palani and Gupta, 2023; Manna, 2021; Azure Readiness, 2023; Vinisky, 2024; Kostic, 2024) 20
-Effectiveness Measure achievement rates, end-user satisfaction, and time to achieve outcomes. (Sondhi et al., 2021; Chen et al., 2022; Malik et al., 2023; Wu, 2021) 4
-Robustness Evaluate stability at peak load and error tolerance. (Vu et al., 2022; Ma et al., 2023; Yu et al., 2021; Wu, 2021; Bedoya et al., 2023; Chen et al., 2022; Vinisky, 2024) 7
-Accuracy Evaluate system consistency through recovery accuracy and fault localization precision. (Green, 2023; Simonsson et al., 2021; Chen et al., 2022) 3
-Fault Tolerance Automated failover success rate during faults; immediate recovery time after detected faults.
-(Fogli et al., 2023; Torkura et al., 2020; Zhang et al., 2019; Dedousis et al., 2023; Gill, 2021; Wickramasinghe, 2023; Zhang et al., 2021a; Camacho et al., 2022; Poltronieri et al., 2022; Nikolaidis et al., 2021; Ma et al., 2023; Meiklejohn et al., 2021; Sondhi et al., 2021; Azure Readiness, 2023; Tomka, 2024; Manna, 2021) 15
-Cost Cost of chaos engineering and savings from preventing outages.
-(Gill, 2021; Patel, 2022; Kesim et al., 2020; Palani and Gupta, 2023; Klinaku et al., 2022; Inc., 2023; Bocetta, 2019; Basiri et al., 2019; National Australia Bank, 2020) 9
-Adaptability Change methods and platforms for new systems. (Vu et al., 2022; Ma et al., 2023; Yu et al., 2021; Wu, 2021) 4
-3.4. Key Components of Chaos Engineering
-Before adopting a chaos engineering framework, organizations should carefully evaluate its key components to ensure that they align with their operational requirements and objectives.
-We have identified several essential components for an effective chaos engineering platform.
-Most of the reviewed studies incorporate a combination of these components, which we have categorized into five central units: experimentation, collaboration and communication, observability and analysis, security and governance, and automation and integration with CI/CD.
-This classification, derived from the functionalities discussed in Section  3.2, is shown in Figure  6.
-The remainder of this section provides a detailed analysis of each component within this framework.
+.
+This framework supports a gradual testing strategy—starting with predictable behaviors and extending to more uncertain or unforeseen failure modes.
+Doing so helps reduce risk while expanding organizational understanding of system weaknesses (Palani and Gupta, 2023a; Inc., 2023; Chen et al., 2022).
+4.2.2. Fault Execution and Control
+This phase ensures that faults are introduced in a controlled, safe, and purposeful manner to validate the system’s resilience under adverse conditions.
+The goal is to simulate real-world failures while minimizing risk and maximizing learning (Zhang et al., 2021b; Green, 2023; Patel, 2022).
+Inject Faults Based on Defined Scenarios
+Chaos experiments begin by executing predefined fault scenarios aligned with the experiment plan.
+These include disruptions such as service crashes, network partitions, and CPU saturation (Zhang et al., 2021b; Cotroneo et al., 2022; Nikolaidis et al., 2023; Siwach et al., 2022).
+Faults are introduced progressively—starting with a minimal scope and gradually increasing impact—to monitor system behavior and avoid widespread disruption (Fogli et al., 2023; Wickramasinghe, 2023; Simonsson et al., 2021; Chen et al., 2024; Green, 2023).
+Execution readiness checks ensure that failure types are verified, system health is stable, and rollback plans are in place (Zhang et al., 2019; Vu et al., 2022; Chen et al., 2022; Kadikar, 2023; Bairyev, 2023).
+Apply Safety Mechanisms
+To protect against unintended system degradation, safety measures are implemented.
+These include automatic abort thresholds that halt experiments if system performance degrades beyond acceptable levels (Frank et al., 2023; Malik et al., 2023; Chen et al., 2024; Camacho et al., 2022).
+Rollback protocols ensure the system can return to a known steady state, while blast radius limits prevent faults from affecting critical components (Simonsson et al., 2021; Green, 2023; Aharon, 2024; Inc., 2023; Kadikar, 2023).
+Notify Stakeholders During Experiment Lifecycle
+Transparent communication is maintained throughout the experiment.
+Key stakeholders are informed at defined checkpoints (start, progress, and end of experiment) to facilitate oversight, enable timely intervention, and foster cross-functional trust in the experimentation process (Torkura et al., 2020; Siwach et al., 2022; Patel, 2022; Bairyev, 2023; Le, 2022; Green, 2023; Zhang et al., 2019).
+4.2.3. Monitoring and Observability
+This phase enables teams to understand how the system behaves under stress by continuously collecting and interpreting telemetry data during chaos experiments.
+Its purpose is to detect deviations, validate resilience assumptions, and guide response strategies (Zhang et al., 2019; Camacho et al., 2022; Inc., 2023; Nikolaidis et al., 2021).
+Capture Live System Telemetry
+The process begins with configuring telemetry tools to collect real-time data, such as metrics, logs, and traces (Simonsson et al., 2021; Pulcinelli et al., 2023; Nikolaidis et al., 2021).
+These data streams encompass both application-level indicators (e.g., service availability, request latency) and infrastructure metrics (e.g., CPU usage, memory consumption) (Gill, 2021; Chen et al., 2024).
+To ensure full system visibility, team members are assigned specific monitoring scopes—some focusing on business-critical services, others on backend components or network behavior (Vu et al., 2022; Nikolaidis et al., 2023).
+When unexpected behaviors occur, teams follow predefined diagnostic steps, such as analyzing error logs, correlating metrics, and reviewing recent deployment changes, to progressively isolate and identify the faulty components (Ma et al., 2023; Sondhi et al., 2021; Patel, 2022; Chen et al., 2022).
+Visualize Experiment Outcomes and Behaviors
+Collected data is visualized through dashboards that track key metrics over time and flag deviations from expected behavior (Inc., 2023; Gill, 2021; Nikolaidis et al., 2021; Chen et al., 2022; Bairyev, 2023; Poltronieri et al., 2022).
+These visual tools guide real-time decisions and help verify whether system behavior aligns with predefined hypotheses (Zhang et al., 2019; Butow, 2018; Bremmers, 2021; Nikolaidis et al., 2023).
+Observations are documented with relevant metadata, such as timestamps, environment context, and test conditions—to support future analysis and continuous learning (Naqvi et al., 2022; Sondhi et al., 2021; Dedousis et al., 2023; Green, 2023)
+4.2.4. Post-Experiment Analysis
+This phase focuses on deriving actionable insights from chaos experiments by assessing impact, diagnosing root causes, and preserving knowledge for future use (Chen et al., 2022; Bremmers, 2021; Frank et al., 2023; Patel, 2022; Al-Said Ahmad et al., 2024).
+The process involves three key areas of functionality.
+Analyze Impact and Diagnose Root Causes
+After fault injection, teams analyze telemetry, such as metrics, logs, and traces, to compare system performance against baseline expectations (Simonsson et al., 2021; Vu et al., 2022; Inc., 2023; Dedousis et al., 2023).
+Deviations are measured to understand the extent of impact on key indicators such as latency or availability (Frank et al., 2023; Poltronieri et al., 2022; Jernberg et al., 2020).
+This helps prioritize which failures matter most.
+Root cause analysis then investigates how the failure unfolded, using traces and service maps to locate the origin and propagation path of faults (Chen et al., 2024; Ma et al., 2023; Zhang et al., 2021a, 2019; Bremmers, 2021; Bairyev, 2023).
+Archive and Document Learnings
+All findings are documented in structured postmortem reports, which capture the fault scenario, affected services, violated hypotheses, and recovery behavior (Vu et al., 2022; Kadikar, 2023; Bairyev, 2023; Jernberg et al., 2020; Chen et al., 2022).
+These insights are stored in internal knowledge bases to support incident response and guide future experiments (Zhang et al., 2019; Siwach et al., 2022; Green, 2023).
+By archiving this information, teams enable iterative learning and systemic improvements in both technical and organizational resilience (Green, 2023; Dedousis et al., 2023).
+4.2.5. Automation and Integration
+This phase automates chaos experiments to minimize manual effort and then integrates them into delivery pipelines and platform workflows, ensuring that resilience testing becomes routine.
+Automate Chaos Testing Pipelines
+Automation reduces manual effort by enabling chaos experiments to run consistently across different stages of the software delivery process, such as development, staging, and production (Chen et al., 2022; Pierce et al., 2021; Bedoya et al., 2023).
+Teams can define test parameters and execution policies in advance, ensuring that faults are injected at predictable times or in response to deployment events (Manna, 2021; Bairyev, 2023).
+This supports repeatability, reduces errors, and allows for continuous resilience validation throughout the development lifecycle (Gill, 2021; Ikeuchi et al., 2020).
+Integrate with platform tools
+To embed chaos engineering into day-to-day engineering workflows, tools are integrated with existing platforms such as CI/CD (Continuous Integration and Continuous Delivery/Deployment) pipelines (e.g., Jenkins, GitHub Actions), cloud environments (e.g., Kubernetes, AWS), and observability frameworks (e.g., Prometheus, Grafana) (Simonsson et al., 2021; Manna, 2021; Camacho et al., 2022; Park et al., 2019).
+This integration enables the automated execution of chaos experiments during the build, test, and deployment phases, while also ensuring that experiment telemetry is collected and visualized in real-time.
+As a result, teams can detect weaknesses early in the lifecycle and improve response times during failure scenarios (Gill, 2021; Siwach et al., 2022; Bairyev, 2023; Nikolaidis et al., 2021).
+4.3. Key Components of Chaos Engineering (RQ1.2)
+Before adopting a chaos engineering platform, organizations should carefully evaluate its key components to ensure that they align with the operational requirements and objectives of the organization.
+Based on our systematic thematic analysis, we identified five primary components: the Experiment Design Unit, Fault Injection Unit, Observability Unit, Post-Experiment Analysis Unit, and Automation and Integration Unit (see Figure 5).
+This classification is based on the functionalities discussed in Section 4.2, mapping each component to the corresponding phase of the experimentation lifecycle.
+Each unit encapsulates modular services, such as hypothesis management, failure orchestration, observability pipelines, and automation engines, that together support scalable, safe, and repeatable experimentation.
+The remainder of this section details each component, its sub-modules, and their roles within the broader platform architecture.
+For details on coding, theme development, and mapping, see our online appendix (see Section 3.6).
 Figure 5. Key Components of Chaos Engineering.
@@
 Key Components of Chaos Engineering.
-3.5. Experimentation Unit
-Successful coordination of chaos experiments depends on a systematic approach to both planning and execution (Zhang et al., 2021a; Nikolaidis et al., 2021).
-To this end, the experimentation unit is built around two foundational components: the experiment designer and the failure injection hub, which will be explored further in this section.
-3.5.1. Experiment Designer
-This module is crucial for structuring chaos experiments, allowing teams to define which systems to target (Zhang et al., 2021a; Camacho et al., 2022).
-After identifying the systems, the teams can choose the fault types most relevant to their testing scenarios, such as network failures or system crashes (Zhang et al., 2021a).
-Once the faults are defined, the experiment designer helps organize the timing and order in which they are introduced, ensuring that they reflect real-world conditions (Camacho et al., 2022).
-By enabling these customized failure simulations, the experiment designer ensures that the tests are well-structured, providing consistent data for postmortem analysis (Zhang et al., 2021a; Konstantinou et al., 2021).
-This structure allows teams to track how the behavior of the system evolves under different failure conditions.
-For example, teams might simulate a server crash during peak traffic to test failover capabilities and measure the impact on user experience (Zhang et al., 2019; Simonsson et al., 2021).
-3.5.2. Fault Injection Hub
-Once the experiment is designed and the faults are organized, the next phase is to inject these controlled faults and workloads into the system, simulating real-world scenarios (Fawcett, 2020; Palani and Gupta, 2023).
-The fault injection process, as outlined by Hseh et al (Hsueh et al., 1997), in Figure  7, involves multiple components working together: a controller, fault injector, fault library, workload generator, workload library, and target system (Nikolaidis et al., 2021; Frank et al., 2021b).
-Each of these components plays a critical role in ensuring the accuracy and smooth execution of the chaos experiment.
-Central to this process is the controller, which manages the introduction of faults and dynamically adjusts based on real-time system feedback, ensuring that the experiment remains aligned with its objectives (Torkura et al., 2020).
-As part of the hub, the fault injector introduces specific disruptions, such as hardware or software failures, selected from a diverse fault library that offers teams a variety of failure scenarios to choose from (Soldani and Brogi, 2021; Frank et al., 2021b).
-In parallel, the workload generator creates synthetic workloads based on predefined patterns stored in the workload library.
-This allows teams to simulate normal or peak traffic conditions to observe how the system handles varying stress levels (Fogli et al., 2023; Kesim et al., 2020).
-These faults and workloads are then applied to the target system, be it a production or preproduction environment, providing teams with valuable information on the weaknesses of the system and the recovery capabilities (Wickramasinghe, 2023; Zhang et al., 2021a).
-Figure 7. Conceptual Schema of Fault Injection(Hsueh et al., 1997).
+4.3.1. Experiment Design Unit
+This architectural component transforms high-level test objectives into system-executable configurations (Soldani and Brogi, 2021; Zhang et al., 2021b).
+It offers formal interfaces to define hypotheses, select test targets, and validate readiness prior to chaos injection (Sbiai et al., 2023; Chen et al., 2024; Dedousis et al., 2023; Frank et al., 2021b; Azure Readiness, 2023).
+The experiment design unit comprises two integrated modules: the Hypothesis Manager and the Target Selector.
+Hypothesis Manager.
+This enables teams to define, store, and manage test hypotheses through structured templates that capture expected system behavior using time-series metrics (e.g., latency, availability) and other indicators such as error rates and log patterns (Gianchandani et al., 2022; Bhaskar, 2022; Kadikar, 2023; Green, 2023; Le, 2022).
+Each hypothesis is programmatically linked to safety controls, including abort thresholds and rollback triggers, via policy configurations (Chen et al., 2024; Simonsson et al., 2021; Poltronieri et al., 2022).
+This ensures automated enforcement of safety during fault injection and accurate post-experiment validation (Zhang et al., 2021b; Dedousis et al., 2023).
+Target Selector.
+This phase scopes where faults will be injected by allowing teams to define experiment boundaries using filters such as environment (e.g., staging, production), infrastructure layer (e.g., compute, network), and service tags or labels (Soldani and Brogi, 2021; Azure Readiness, 2023; Green, 2023).
+It ensures controlled impact by enforcing blast radius limits, so only approved targets are affected (Sbiai et al., 2023).To maintain operational safety, it performs validation checks on all configurations, verifying correctness, authorization, and compatibility before execution.
+This prevents unsafe or unintended experiments from proceeding (Zhang et al., 2021b; Cotroneo et al., 2022; Bairyev, 2023; Chen et al., 2024).
+4.3.2. Fault Injection Unit
+This is the core architectural component responsible for executing chaos experiments by introducing faults into the system under test.
+It is composed of three core modules: the Fault Execution Engine, the Safety Controller, and the Alert Manager, each supporting a distinct aspect of the fault injection lifecycle.
+Fault Execution Engine
+This component orchestrates the actual injection of predefined faults and workloads.
+Its design follows the foundational model by Hsueh et al.
+(Hsueh et al., 1997) (See Figure 6), and has been extended to suit modern, distributed environments (Soldani and Brogi, 2021; Frank et al., 2021b).
+It comprises several modules: (1) A Controller that sequences experiment steps and responds to telemetry, (2) A Fault Injector that applies disruptions from a configurable fault library, (3) A Workload Generator that simulates system activity using patterns from a workload library, (4) The Target System, which receives the injected faults and workloads and emits telemetry for analysis (Wickramasinghe, 2023).
+Other components from the original model (see Figure 6)—such as the Monitor—are represented in the subsequent architectural component through the Observability Unit, which is responsible for telemetry collection.
+Similarly, the Data Collector and Chaos Analyzer are addressed within the Post-Experiment Analysis Unit, which supports impact assessment and root cause diagnosis following fault injection.
+Safety Controller
+This enforces runtime protection by embedding safeguards directly into the fault injection lifecycle (Zhang et al., 2021b; Cotroneo et al., 2022; Bhaskar, 2022; Lenka et al., 2018).
+It comprises three subsystems:
+  • Blast Radius Control.
+This limits the impact of injected faults by using traffic routing and isolation rules.
+For instance, in Kubernetes, these rules restrict disruptions to specific services, namespaces, or pods within the system (Soldani and Brogi, 2021; Wu, 2021; Vu et al., 2022; Kesim et al., 2020).
+  • Abort Monitor. This continuously checks key metrics (e.g., error rate, response time) and halts the experiment if thresholds are breached (Kadikar, 2023; Green, 2023; Le, 2022).
+  • Rollback Handler.
+This restores the system to a steady state by triggering recovery workflows or replaying system snapshots (Cotroneo et al., 2022; Soldani and Brogi, 2021; Kadikar, 2023; Bhaskar, 2022).
+Alert Manager
+This component coordinates communication with stakeholders by triggering real-time notifications across channels (e.g., Slack, email) at critical points: start, escalation, and end (Gill, 2021; Green, 2023).
+It also enforces escalation policies when thresholds are breached to ensure rapid response (Patel, 2022; Azure Readiness, 2023; Camacho et al., 2022; Roa, 2022).
+Figure 6. Conceptual Schema of Fault Execution Engine(Hsueh et al., 1997).
 […]
 Conceptual schema of fault injection.
-3.6. Collaboration and Communication Unit
-Executing chaos experiments across teams requires strong communication and access to data from previous experiments (Kadikar, 2023; Patel, 2022; Siwach et al., 2022), which are supported by the alert and notification system and the knowledge base.
-3.6.1. Alert and Notification System
-The hub integrates chaos tools with existing alert systems to ensure smooth communication during chaos experiments.
-This integration ensures that teams are immediately notified of critical events such as experiment initiation, fault injection, and any unexpected behaviors during the process (Starr et al., 2022; Vinisky, 2024).
-These real-time notifications are distributed through multiple channels, including email and Slack, ensuring that all stakeholders receive timely updates.
-For example, if an issue arises, the system immediately alerts the relevant individuals, allowing them to respond quickly and address the problem effectively (Torkura et al., 2020; Shah, 2021).
-This prompt response mechanism is critical for maintaining system stability, as it helps mitigate potential risks before they escalate (Siwach et al., 2022).
-3.6.2. Knowledge Base
-Maintaining a well-organized archive of previous chaos experiments is crucial to continuously improve the resilience of the system and support real-time communication between users (Kadikar, 2023).
-This knowledge base is a central repository that stores comprehensive documentation that includes experiment designs, execution procedures, and results (Zhang et al., 2019; Torkura et al., 2020).
-Recalling past experiments, teams can leverage valuable insights from successful and unsuccessful tests, fostering a cycle of learning and enhancement.
-Detailed post-mortem reports, for example, allow teams to thoroughly analyze the results, identify improvement areas, and strengthen the resilience of the system (Inc., 2023; Palani and Gupta, 2023).
-Furthermore, using standardized templates ensures that this knowledge is systematically preserved, promoting consistency and reliability in future chaos experiments (Siwach et al., 2022; Zhang et al., 2019).
-3.7. Observability and Analysis Unit
-Monitoring system behavior during chaos experiments is crucial for understanding faults’ impact and identifying areas for improvement (Gianchandani et al., 2022; Nikolaidis et al., 2021; Yu et al., 2021).
-Essential components, such as data collector and data analyzer, play a vital role in this process, ensuring that the systems are evaluated and optimized effectively (Zhang et al., 2021a; Simonsson et al., 2021).
-3.7.1. Data Collector
-Effective data collection is essential during chaos experiments, capturing various types of information such as performance metrics, logs, and traces (Gatev, 2021).
-These data points clearly explain how the system responds under stress (Zhang et al., 2021a).
-The data collector gathers performance metrics to track key indicators such as response times, error rates, latency, and system throughput.
-These data are crucial to measuring the effectiveness of the system in detecting faults  (Cotroneo et al., 2022).
-Logs document significant events within the system, including errors and state transitions.
-These records help diagnose issues and improve the understanding of system behavior during failures  (Vu et al., 2022; Cotroneo et al., 2022).
-Tracing methods can monitor the flow of requests across distributed services, enabling the identification of bottlenecks and observing how traffic is managed under stress (Zhang et al., 2021a).
-All data collected are visualized through a centralized dashboard, providing a straightforward and organized interface that allows quick assessment of performance indicators and overall system health (Malik et al., 2023; Basiri et al., 2019; Camacho et al., 2022).
-3.7.2. Data Analyzer
-The information collected during chaos experiments is systematically analyzed to identify essential trends and patterns (Zhang et al., 2021a; Simonsson et al., 2021).
-This process involves impact analysis, which evaluates how failures influence system performance based on previously collected metrics (Pulcinelli et al., 2023).
-By analyzing these data, it is possible to determine how disruptions affect the system and identify its most vulnerable components  (Fawcett, 2020; Cotroneo et al., 2022).
-Additionally, root cause analysis leverages logs and traces to trace failures back to their origins, whether due to a configuration error, service malfunction, or hardware fault (Zhang et al., 2021a; Cotroneo et al., 2022).
-3.8. Security and Governance Unit
-Ensuring the secure execution of chaotic experiments in production environments is essential to prevent disruptions and unauthorized actions, achieved through adequate safety and access control mechanisms (Al-Said Ahmad et al., 2024; Green, 2023; Gianchandani et al., 2022).
-First, safety controls mitigate the risk of chaos experiments that cause severe outages by incorporating predefined thresholds, such as automated rollbacks, to restore systems to a steady state if an experiment causes prolonged disruptions (Zhang et al., 2021a; Gill, 2022).
-These controls allow the system to automatically abort experiments when severe impacts are detected, maintaining overall system stability (Kadikar, 2023).
-By limiting the blast radius to specific areas, safety controls ensure that experiments do not cause widespread failures (Simonsson et al., 2021; Gill, 2021), with fail-safe conditions ensuring immediate containment of issues (Bhaskar, 2022).
-Second, access control mechanisms further enhance security by restricting the ability to initiate, modify, or stop chaos experiments only to authorized personnel (Torkura et al., 2020; Sbiai et al., 2023).
-This ensures that only individuals with the appropriate permissions can execute or alter experiments, particularly in sensitive environments where unauthorized changes could lead to significant security and compliance risks  (Fogli et al., 2023; Tivey et al., 2019).
-3.9. Automation and Integration Unit
-This unit is vital in embedding chaos engineering into the software development lifecycle.
-It automates and integrates chaos experiments with existing systems to ensure consistent and efficient execution.
-First, the unit connects the chaos engineering platform with essential tools such as CI/CD pipelines, cloud platforms, and monitoring systems.
-This integration allows chaos experiments to be seamlessly incorporated into routine workflows such as deployments, updates, and system monitoring, enabling automatic execution as part of regular operations (Zhang et al., 2021a; Torkura et al., 2020).
-Second, the unit automates the execution of chaos experiments and includes a scheduling mechanism that allows tests to run at predefined intervals, such as during off-peak hours or after system updates (Cotroneo et al., 2022; Zhang et al., 2019).
-This ensures that chaos experiments are performed regularly and reliably, reducing the risk of human error and allowing organizations to test their systems without interrupting normal activities (Camacho et al., 2022; Simonsson et al., 2021).
+4.3.3. Observability Unit
+This component implements the infrastructure for capturing, processing, and presenting telemetry data during chaos experiments (Pulcinelli et al., 2023; Green, 2023).
+It is composed of two core subsystems: the Data Collector and the Dashboard.
+Data Collector
+This subsystem captures diverse telemetry from systems under test (Zhang et al., 2019, 2021b; Soldani and Brogi, 2021; Naqvi et al., 2022; Mooney, 2023).
+It contains adapters/connectors for various data sources, buffering mechanisms for handling high-volume streams, and preprocessing capabilities for initial data normalization, such as timestamp alignment, unit standardization (e.g., converting memory metrics to a common format), and removal of malformed or incomplete entries (Vu et al., 2022; Nikolaidis et al., 2023; Pulcinelli et al., 2023; Green, 2023; Camacho et al., 2022).
+Metrics Collector This component captures quantitative measurements, primarily in the form of time-series metrics, using instrumentation libraries and collection endpoints (Nikolaidis et al., 2023; Pulcinelli et al., 2023; Green, 2023; Azure Readiness, 2023).
+It stores data in specialized time-series databases with retention policies and aggregation capabilities (Nikolaidis et al., 2021; Pulcinelli et al., 2023).
+It supports both system-level metrics (CPU, memory, network) and application-level indicators (throughput, error rates) with statistical analysis functions (Zhang et al., 2021b; Soldani and Brogi, 2021; Pulcinelli et al., 2023; Fawcett, 2020; Bhaskar, 2022).
+Log Aggregator This component collects logs from all participating systems during a chaos experiment (Soldani and Brogi, 2021; Vu et al., 2022).
+It uses log shippers such as Fluentd   [^3] https://www.fluentd.org/ to forward entries to a central log system (Green, 2023; Azure Readiness, 2023).
+Logs are parsed into structured formats (e.g., JSON) for consistency (Chen et al., 2022; Sousa et al., 2018).
+The component typically indexes log data to enable fast retrieval and supports querying by time, service, or error type (Kadikar, 2023; Dedousis et al., 2023).
+Tracer This component tracks request flows across service boundaries in distributed systems (Zhang et al., 2021b; Nikolaidis et al., 2023; Naqvi et al., 2022).
+It utilizes lightweight instrumentation, context propagation protocols, and sampling algorithms for data volume management (Vu et al., 2022; Naqvi et al., 2022; Green, 2023).
+It can also visualize service relationships, analyze performance bottlenecks, and correlate traces with metrics and logs (Naqvi et al., 2022; Nikolaidis et al., 2021).
+Dashboard
+This component renders configurable monitoring interfaces from multiple data sources (Nikolaidis et al., 2021, 2023; Pulcinelli et al., 2023).
+It can include sub-components such as query interfaces, graphical rendering engines, and templating systems (Basiri et al., 2019; Gianchandani et al., 2022; Green, 2023).
+The common features include real-time data refresh, threshold alerting, time-range selection, and export capabilities for sharing observations (Bhaskar, 2022; Green, 2023; Pulcinelli et al., 2023).
+4.3.4. Post-Experiment Analysis Unit
+This unit implements the processes needed to evaluate chaos experiments, identify technical weaknesses, and retain organizational knowledge (Zhang et al., 2019; Simonsson et al., 2021; Vu et al., 2022; Bremmers, 2021; Kadikar, 2023).
+It includes two subsystems: the Chaos Analyzer and the Knowledge Base Manager.
+Chaos Analyzer
+This unit processes telemetry data to assess system behavior and diagnose failure mechanisms.
+Impact Analyzer This module compares actual telemetry (e.g., metrics, logs, traces) against steady-state thresholds to quantify performance degradation (Simonsson et al., 2021; Vu et al., 2022; Frank et al., 2023).
+It calculates impact scores based on key performance indicators (KPIs) such as latency and availability, and ranks vulnerabilities for follow-up (Bremmers, 2021).
+Root Cause Analyzer This module Correlates data across logs, traces, and metrics to trace fault propagation and uncover causal links (Kadikar, 2023; Kostic, 2024).
+It applies anomaly detection and pattern-matching to distinguish root issues from surface symptoms (Zhang et al., 2019; Simonsson et al., 2021).
+Knowledge Base Manager
+This module stores post-experiment findings, including violated hypotheses, impact severity, and root causes, in structured formats.
+These are indexed and archived in resilience knowledge bases, supporting institutional memory, cross-team learning, and onboarding of new engineers (Manna, 2021; Kadikar, 2023; Vu et al., 2022).
+4.3.5. Automation and Integration Unit
+This component operationalizes the process of continuous chaos validation.
+It allows teams to schedule, coordinate, and trigger fault injections across environments and platforms (Vu et al., 2022; Zhang et al., 2021b; Sbiai et al., 2023).
+The unit includes two major subsystems: Workflow Engine and Platform Connector.
+Workflow Engine
+This module orchestrates the automatic execution of chaos experiments based on scheduled intervals or event-driven triggers, such as deployments, pipeline stages, or config changes (Vu et al., 2022; Ma et al., 2023; Basiri et al., 2019; Fogli et al., 2023; Simonsson et al., 2021; Zhang et al., 2021b; Cotroneo et al., 2022; Camacho et al., 2022; Nikolaidis et al., 2021).
+It applies user-defined policies to determine timing, scope, and test conditions.
+The engine ensures fault injection is repeatable and synchronized with real system activity, minimizing manual intervention (Pierce et al., 2021; Nikolaidis et al., 2023; Meiklejohn et al., 2021; Chen et al., 2022; Gill, 2021; Bremmers, 2021; Gill, 2022; Bairyev, 2023; Sbiai et al., 2023; Torkura et al., 2020)
+Platform Connector
+This facilitates seamless integration with external environments.
+It interfaces with CI/CD systems (e.g., Jenkins, GitHub Actions), cloud platforms (e.g., AWS, Azure, GCP), and observability tools (e.g., Prometheus, Grafana) (Vu et al., 2022; Zhang et al., 2021b; Frank et al., 2021b; Nikolaidis et al., 2023; Meiklejohn et al., 2021; Malik et al., 2023; Basiri et al., 2019).
+Through these connections, it supports the triggering of chaos experiments based on pipeline events, deployment stages, or infrastructure changes.
+It also ensures synchronized data exchange across orchestration layers, telemetry systems, and alerting channels.
+(Fogli et al., 2023; Torkura et al., 2020; Zhang et al., 2019; Dedousis et al., 2023; Simonsson et al., 2021; Cotroneo et al., 2022)
+4.4. Quality Requirements for Chaos Engineering Platform (RQ1.3)
+We have identified 11 key quality requirements for a chaos engineering platform from the selected literature: performance, recoverability, resilience, timeliness, observability, scalability, effectiveness, robustness, accuracy, fault tolerance, and adaptability.
+Table 10 provides these requirements and the corresponding metrics for measuring them.
+Table 10. Quality Attributes, Measurement Metrics, and Referenced Articles.
+Quality Attribute Description Measurement Metrics Articles Referenced Total
+Performance Evaluate resource and performance under attack; align chaos tests with production metrics.
+Response time (ms), CPU/memory utilization (%), throughput (req/sec).
+(Fogli et al., 2023; Torkura et al., 2020; Zhang et al., 2019; Dedousis et al., 2023; Patel, 2022; Gill, 2021; Cotroneo et al., 2022; Nikolaidis et al., 2021; Vu et al., 2022; Klinaku et al., 2022) 10
+Recoverability Assess how quickly the system restores functionality after a failure.
+Mean Time to Repair (MTTR), rollback time, recovery time objective (RTO).
+(Fogli et al., 2023; Torkura et al., 2020; Dedousis et al., 2023; Patel, 2022; Gill, 2021; Cotroneo et al., 2022; Camacho et al., 2022; Zhang et al., 2021a; Nikolaidis et al., 2021; Pierce et al., 2021; Vu et al., 2022; Frank et al., 2021b; Malik et al., 2023; Zhang et al., 2023; Tomka, 2024; Kadikar, 2023) 16
+Resilience Evaluate the system’s ability to operate under failure without service disruption.
+Percentage of successful requests under fault, error rate, degradation threshold.
+(Fogli et al., 2023; Torkura et al., 2020; Dedousis et al., 2023; Bremmers, 2021; Zhang et al., 2021a; Al-Said Ahmad et al., 2024; Vu et al., 2022; Nikolaidis et al., 2023; Frank et al., 2023; Kassab et al., 2022; Konstantinou et al., 2021; Nombela, 2023; Krishnamurthy, 2021; Bocetta, 2019; Manna, 2021; Kostic, 2024; Krivas and Portela, 2020; National Australia Bank, 2020; Azure Readiness, 2023) 18
+Timeliness Measure how quickly faults are detected and addressed.
+Time to detect (TTD), alert response time.
+(Camacho et al., 2022; Pierce et al., 2021; Vu et al., 2022; Frank et al., 2021b; Malik et al., 2023; Cotroneo et al., 2022; Basiri et al., 2019; Gill, 2021; Zhang et al., 2019; Durai, 2022; Shah, 2021; Katirtzis, 2022; Dua, 2024; Wu, 2021; Palani and Gupta, 2023a; Manna, 2021) 16
+Observability Assess the visibility into system internals during chaos.
+Log completeness (%), alerting coverage, trace propagation rate, telemetry accuracy.
+(Fogli et al., 2023; Torkura et al., 2020; Dedousis et al., 2023; Inc., 2023; Bremmers, 2021; Patel, 2022; Gill, 2021; Camacho et al., 2022; Torkura et al., 2021; Zhang et al., 2021a; Vu et al., 2022; Frank et al., 2023; Malik et al., 2023; Kesim et al., 2020; Siwach et al., 2022; Konstantinou et al., 2021; Hawkins, 2020; Manna, 2021; Kostic, 2024; Kadikar, 2023; Azure Readiness, 2023) 20
+Scalability Gauge how chaos testing scales with system size and complexity.
+Number of nodes/services covered per experiment, test execution time at scale.
+(Fogli et al., 2023; Torkura et al., 2020; Zhang et al., 2019; Patel, 2022; Gill, 2021; Vu et al., 2022; Sondhi et al., 2021; Malik et al., 2023; Klinaku et al., 2022; Camilli et al., 2022; Nikolaidis et al., 2023; Kassab et al., 2022; Rivera et al., 2023; Krishnamurthy, 2021; Durai, 2022; Wu, 2021; Palani and Gupta, 2023a; Manna, 2021; Azure Readiness, 2023; Vinisky, 2024; Kostic, 2024) 20
+Effectiveness Determine whether chaos tests produce meaningful insights. Number of unknown faults revealed, test success rate. (Sondhi et al., 2021; Chen et al., 2022; Malik et al., 2023; Wu, 2021) 4
+Robustness Evaluate stability at peak load and error tolerance.
+System uptime under stress, error rate at max load.
+(Vu et al., 2022; Ma et al., 2023; Yu et al., 2021; Wu, 2021; Bedoya et al., 2023; Chen et al., 2022; Vinisky, 2024) 7
+Accuracy Assess how precisely the system recovers and localizes issues. Post-recovery state correctness (%), fault localization accuracy. (Green, 2023; Simonsson et al., 2021; Chen et al., 2022) 3
+Fault Tolerance Measure the success of self-healing and failover mechanisms.
+Failover completion time, success rate of automatic recovery actions, redundancy coverage.
+(Fogli et al., 2023; Torkura et al., 2020; Zhang et al., 2019; Dedousis et al., 2023; Gill, 2021; Wickramasinghe, 2023; Zhang et al., 2021a; Camacho et al., 2022; Poltronieri et al., 2022; Nikolaidis et al., 2021; Ma et al., 2023; Meiklejohn et al., 2021; Sondhi et al., 2021; Azure Readiness, 2023; Tomka, 2024; Manna, 2021) 15
+Adaptability Determine how easily the chaos system fits into varied environments.
+Number of supported platforms, integration effort (e.g., setup time), configuration reuse rate.
+(Vu et al., 2022; Ma et al., 2023; Yu et al., 2021; Wu, 2021) 4
 5. Motivation Behind Chaos Engineering (RQ2)
-This section delves into the challenges driving chaos engineering adoption in organizations. Figure 8 shows the technical and socio-technical challenges we identified from the reviewed literature.
+This section directly addresses RQ2 by delving into the challenges driving chaos engineering adoption in organizations.
+Figure 7 shows the technical and socio-technical challenges we identified from the reviewed literature.
+These challenges were identified through our coding and thematic analysis (see Section 3.5) and are documented in detail in our online appendix (see Section 3.6).
 Figure 7. Challenges that Drive Chaos Engineering Adoption.
@@
 5.1. Technical Challenges
-Chaos engineering aims to tackle various technical challenges that undermine system resilience. The following subsections explore these challenges in detail.
+Chaos engineering aims to tackle various technical challenges that undermine system resilience.For example (see Section 3.6), the theme ‘Unanticipated System Failures’ includes codes such as ‘network-partition’ and ‘cascade-failure’, capturing failures that are often missed by traditional software testing.
+Similarly, the theme ‘Inadequate Fault Tolerance Mechanisms’ incorporates codes like ‘multi-region outage’ and ‘resource exhaustion’, highlighting gaps that chaos experiments are designed to expose.
+The following subsections discuss these and related technical challenges in more detail.
 Unanticipated System Failures
@@
 In contrast, chaos engineering adopts a proactive approach by deliberately introducing failures, such as server crashes, network interruptions, and resource exhaustion, in environments that closely resemble production.
-As a result, engineers can observe how the system behaves under stress, thus revealing hidden weaknesses that conventional testing methods might overlook.
+As a result, engineers can observe how the system behaves under stress, thus revealing hidden weaknesses that traditional testing methods might overlook.
 By focusing on unpredictable interactions between distributed components, chaos engineering effectively uncovers vulnerabilities that arise from the complexity of the system (Bairyev, 2023).
@@
 Many distributed systems are built with fault tolerance mechanisms to handle routine issues, such as server failures or network delays (Inc., 2023).
-However, when larger-scale failures occur, e.g., multi-region outages or prolonged network disruptions, these mechanisms often fail to maintain system stability (Palani and Gupta, 2023; Silveira, 2023; Nallainathan, 2023).
+However, when larger-scale failures occur, e.g., multi-region outages or prolonged network disruptions, these mechanisms often fail to maintain system stability (Palani and Gupta, 2023a; Silveira, 2023; Nallainathan, 2023).
 This shortfall occurs because mechanisms designed to handle more minor issues may not be able to handle the broader impact of significant failures, leading to severe performance degradation or even total system outages (Basiri et al., 2019; Gill, 2021).
@@
 These controlled failures are monitored through observability tools that provide real-time visibility into metrics such as latency, error rates, and resource consumption (Wachtel, 2023).
-By analyzing these metrics, teams gain detailed insight into how services handle stress - uncovering issues such as delayed responses or inefficient resource utilization that would otherwise remain hidden during normal operations (Bremmers, 2021; Palani and Gupta, 2023).
+By analyzing these metrics, teams gain detailed insight into how services handle stress - uncovering issues such as delayed responses or inefficient resource utilization that would otherwise remain hidden during normal operations (Bremmers, 2021; Palani and Gupta, 2023a).
 5.2. Socio-Technical Challenges
 Chaos engineering can address various socio-technical challenges, expanding its impact beyond technical issues to encompass the complex interplay between people, processes, and technology (Miles, 2019b; Thomas et al., 2019; Patriarca et al., 2018).
+For instance, the theme ‘Inconsistent Experiment Design’ maps to codes such as ‘variable test timing’ and ‘lack of standardization’.
+Additional examples and detailed code-theme mappings can be found in the online appendix(see Section 3.6).
 Inconsistent Experiment Design and Execution
@@
 Chaos engineering solves these challenges by enabling controlled experimentation that tests system behavior under simulated and real-world failure conditions (Torkura et al., 2020; Kostic, 2024).
-In production environments, controlled blast radiuses ensure that failures are confined to isolated parts of the system, thereby preventing widespread disruption (Palani and Gupta, 2023; Siwach et al., 2022; Gianchandani et al., 2022; Mooney, 2023; Bocetta, 2019).
+In production environments, controlled blast radiuses ensure that failures are confined to isolated parts of the system, thereby preventing widespread disruption (Palani and Gupta, 2023a; Siwach et al., 2022; Gianchandani et al., 2022; Mooney, 2023; Bocetta, 2019).
 Predefined abort conditions and rollback procedures allow teams to stop experiments and recover quickly if the stability of the system is threatened (Abdul, 2024; Sbiai et al., 2023; Siwach et al., 2022).
@@
 For example, simulating a sudden network partition might show whether monitoring systems detect the issue promptly and if the system can reroute traffic as intended (Vinisky, 2024; Green, 2023; Gill, 2021).
-Similarly, inducing a Central Processing Unit (CPU) or memory exhaustion event could expose delays in triggering alerts or inefficiencies in the automated recovery process (Camacho et al., 2022; Kadikar, 2023; Palani and Gupta, 2023; Roa, 2022).
+Similarly, inducing a Central Processing Unit (CPU) or memory exhaustion event could expose delays in triggering alerts or inefficiencies in the automated recovery process (Camacho et al., 2022; Kadikar, 2023; Palani and Gupta, 2023a; Roa, 2022).
 By creating these controlled failure scenarios, chaos engineering helps teams identify where their incident response protocols fall short, whether in monitoring, alert escalation, or failover processes (Wu, 2021; Zhang et al., 2021a).
 This enables teams to proactively adjust their strategies, improving their ability to handle actual incidents faster, more accurately, and with minimal disruption (Dua, 2024; Rivera et al., 2023).
+Table 11. Mapping of Core Chaos Engineering Functionalities with Organizational Benefits.
+Core Functionality Activity Performed Benefits to Organization Articles Referenced Total
+Identify Test Targets and Failure Scenarios Select critical components (e.g., APIs, services) and likely fault types based on architecture and risk.
+Focuses efforts on the most vulnerable areas, improving preparedness and reducing downtime (Bremmers, 2021; Torkura et al., 2020; Zhang et al., 2019; Dedousis et al., 2023; Simonsson et al., 2021; Inc., 2023; Gill, 2021; Wickramasinghe, 2023; Zhang et al., 2021b; Cotroneo et al., 2022; Poltronieri et al., 2022; Zhang et al., 2021a; Frank et al., 2021a; Pierce et al., 2021; Soldani and Brogi, 2021; Al-Said Ahmad et al., 2024; Vu et al., 2022; Nikolaidis et al., 2023; Moon, 2022; Anwar and Arunachalam, 2019; Dua, 2024; Gill, 2022; Fawcett, 2020; Manna, 2021; Gianchandani et al., 2022; Abdul, 2024; Azure Readiness, 2023; Roa, 2022; Krivas and Portela, 2020; Kostic, 2024; Vinisky, 2024; Colyer, 2019; Green, 2023) 33
+Define Steady-State Behavior Establish baseline performance metrics (e.g., latency, throughput) Provides a reference for detecting anomalies and validating system health under failure.
+(Fogli et al., 2023; Torkura et al., 2020; Zhang et al., 2019; Dedousis et al., 2023; Simonsson et al., 2021; Inc., 2023; Bremmers, 2021; Patel, 2022; Gill, 2021; Cotroneo et al., 2022; Nikolaidis et al., 2021; Pierce et al., 2021; Frank et al., 2023; Konstantinou et al., 2021; Naqvi et al., 2022; Kesim et al., 2020; Park et al., 2019; Zhang et al., 2023; Siwach et al., 2022; Jernberg et al., 2020; Jakkaraju, 2020; Le, 2022; Moon, 2022; Bocetta, 2019; Palani and Gupta, 2023b; Gill, 2022; Fawcett, 2020; Bairyev, 2023; Abdul, 2024; Roa, 2022; Aharon, 2024; Krivas and Portela, 2020; Kostic, 2024) 33
+Formulate Testable Hypotheses Describe expected system responses to faults in measurable terms.
+Guides experiment design and enables structured interpretation of results (Fogli et al., 2023; Torkura et al., 2020; Zhang et al., 2019; Dedousis et al., 2023; Simonsson et al., 2021; Inc., 2023; Bremmers, 2021; Patel, 2022; Gill, 2021; Cotroneo et al., 2022; Frank et al., 2023; Konstantinou et al., 2021; Naqvi et al., 2022; Kesim et al., 2020; Park et al., 2019; Zhang et al., 2023; Siwach et al., 2022; Jernberg et al., 2020; Jakkaraju, 2020; Bocetta, 2019; Palani and Gupta, 2023b; Gill, 2022; Fawcett, 2020; Bairyev, 2023; Abdul, 2024) 25
+Inject Faults Based on Defined Scenarios Execute fault (e.g., CPU overload, network delay) on selected systems.
+Validates resilience strategies; exposes weak points under controlled stress (Zhang et al., 2019, 2021b; Cotroneo et al., 2022; Zhang et al., 2021a; Nikolaidis et al., 2021; Frank et al., 2021b; Pierce et al., 2021; Soldani and Brogi, 2021; Vu et al., 2022; Chen et al., 2024; Nikolaidis et al., 2023; Ma et al., 2023; Meiklejohn et al., 2021; Sondhi et al., 2021; Ikeuchi et al., 2020; Chen et al., 2022; Basiri et al., 2019; Gill, 2022; Gianchandani et al., 2022; Sbiai et al., 2023; Kadikar, 2023; Bhaskar, 2022) 22
+Apply Safety Mechanisms Enforce abort conditions, rollback procedures, and limit blast radius.
+Minimizes impact, safeguards critical systems, ensures quick recovery.
+(Bremmers, 2021; Patel, 2022; Gill, 2021; Wickramasinghe, 2023; Frank et al., 2021b; Pierce et al., 2021; Soldani and Brogi, 2021; Al-Said Ahmad et al., 2024; Vu et al., 2022; Konstantinou et al., 2021; Ikeuchi et al., 2020; Chen et al., 2022; Naqvi et al., 2022; Jakkaraju, 2020; Starr et al., 2022; Katirtzis, 2022; Dua, 2024; Tivey et al., 2019; Palani and Gupta, 2023b; Bairyev, 2023; Sbiai et al., 2023; Kadikar, 2023; Green, 2023; Vinisky, 2024) 24
+Notify Stakeholders During Experiment Lifecycle Send alerts and updates before, during, and after chaos experiments.
+Promotes transparency, readiness, and coordinated response among teams.
+(Torkura et al., 2020; Zhang et al., 2019; Simonsson et al., 2021; Patel, 2022; Gill, 2021; Wickramasinghe, 2023; Zhang et al., 2021b; Torkura et al., 2021; Nikolaidis et al., 2021; Pierce et al., 2021; Soldani and Brogi, 2021; Konstantinou et al., 2021; Siwach et al., 2022; Krishnamurthy, 2021; Durai, 2022; Tomka, 2024; Dua, 2024; Tivey et al., 2019; Manna, 2021; Abdul, 2024; Green, 2023; Kostic, 2024; Vinisky, 2024; Azure Readiness, 2023) 24
+Capture Live System Telemetry Monitor and record real-time metrics, logs, and traces during injection.
+Enables detailed analysis of system and supports rapid anomaly detection.
+(Zhang et al., 2019; Dedousis et al., 2023; Simonsson et al., 2021; Inc., 2023; Gill, 2021; Camacho et al., 2022; Poltronieri et al., 2022; Nikolaidis et al., 2021; Vu et al., 2022; Chen et al., 2024; Nikolaidis et al., 2023; Ma et al., 2023; Sondhi et al., 2021; Pulcinelli et al., 2023; Naqvi et al., 2022; Zhang et al., 2023; Siwach et al., 2022; Jernberg et al., 2020; Le, 2022; Butow, 2018) 20
+Visualize Experiment Outcomes and Behavior Display telemetry on dashboards to highlight key deviations and trends.
+Provides insights for decisions and facilitates shared understanding.
+(Fogli et al., 2023; Torkura et al., 2020; Zhang et al., 2019; Dedousis et al., 2023; Simonsson et al., 2021; Inc., 2023; Bremmers, 2021; Zhang et al., 2021b; Cotroneo et al., 2022; Nikolaidis et al., 2021; Green, 2023) 11
+Analyze Impact & Root Causes Compare metrics to baselines and trace root causes.
+Identifies weaknesses, guides improvements.
+(Fogli et al., 2023; Torkura et al., 2020; Zhang et al., 2019; Dedousis et al., 2023; Simonsson et al., 2021; Inc., 2023; Bremmers, 2021; Zhang et al., 2021b; Cotroneo et al., 2022; Poltronieri et al., 2022; Zhang et al., 2021a; Nikolaidis et al., 2021; Pierce et al., 2021; Soldani and Brogi, 2021; Bairyev, 2023; Green, 2023) 16
+Archive and Document Learnings Store findings, root causes, and lessons learned in accessible knowledge bases.
+Preserves knowledge and supports future analysis, audits, and training.
+(Fogli et al., 2023; Torkura et al., 2020; Zhang et al., 2019; Dedousis et al., 2023; Simonsson et al., 2021; Inc., 2023; Bremmers, 2021; Patel, 2022; Gill, 2021; Al-Said Ahmad et al., 2024; Vu et al., 2022; Frank et al., 2023; Chen et al., 2022; Kesim et al., 2020; Basiri et al., 2019; Camilli et al., 2022; Yu et al., 2021; Siwach et al., 2022; Konstantinou et al., 2021; Wu, 2021; Kadikar, 2023; Green, 2023) 22
+Automate Chaos Testing Pipelines Schedule and execute experiments with scripts or workflows.
+Enhances consistency, lowers overhead, supports continuous validation.
+(Torkura et al., 2020; Zhang et al., 2019; Simonsson et al., 2021; Gill, 2021; Nikolaidis et al., 2021; Pierce et al., 2021; Ikeuchi et al., 2020; Chen et al., 2022; Bedoya et al., 2023; Park et al., 2019; Siwach et al., 2022; Manna, 2021; Green, 2023) 13
+Integrate with Platform Tools Link chaos tools to CI/CD, monitoring, and cloud systems for orchestration.
+Embeds resilience testing into delivery pipelines and ensures synchronized operations.
+(Torkura et al., 2020; Dedousis et al., 2023; Bremmers, 2021; Cotroneo et al., 2022; Camacho et al., 2022; Poltronieri et al., 2022; Nikolaidis et al., 2021; Frank et al., 2021b; Soldani and Brogi, 2021) 9
 5.3. Benefits
 Table 11 outlines the organizational benefits derived from the functionalities described in Section 4.2, addressing the previously identified challenges.
-These functions include defining chaos experiments, facilitating team communication, notifying stakeholders, maintaining documentation, and using real-time monitoring dashboards.
+These functionalities were identified through our thematic analysis process.
+They include defining chaos experiments, facilitating team communication, notifying stakeholders, maintaining documentation, and using real-time monitoring dashboards.
 The corresponding activities, such as establishing experimental parameters, analyzing system performance metrics, and fostering cross-functional collaboration, contribute to several organizational advantages.
 These include improved system understanding, enhanced error-handling capabilities, vital team coordination during stress conditions, and proactive system tuning.
-Furthermore, automation of tasks and real-time system monitoring can minimize disruptions and optimize system performance and user experience.
-Table 9. Mapping of Key Activities Performed by Chaos Engineering with Organizational Benefits
-Key Functions Activity Performed Benefits of Organization Articles Total
-Define Chaos Experiments Set experiment scope, define normal conditions, hypothesize, and inject fault.
-Gain system insights, refine the design, establish a baseline, and test error handling.
-(Kostic, 2024; Gill, 2021; Manna, 2021; Inc., 2023; Wickramasinghe, 2023; Mooney, 2023; Fawcett, 2020; Vinisky, 2024; Al-Said Ahmad et al., 2024; Pierce et al., 2021; Torkura et al., 2020; Vu et al., 2022; Bedoya et al., 2023; Siwach et al., 2022; Zhang et al., 2019; Le, 2022; Palani and Gupta, 2023; Soldani and Brogi, 2021; Azure Readiness, 2023; Dedousis et al., 2023; Colyer, 2019; Patel, 2022; Shah, 2021; Green, 2023; Fogli et al., 2023; National Australia Bank, 2020; Cotroneo et al., 2022; Konstantinou et al., 2021; Zhang et al., 2023; Bairyev, 2023) 30
-Facilitate Team Coordination Create groups for team collaboration.
-Strengthens coordination and insight sharing.
-(Gill, 2021; Manna, 2021; Dua, 2024; Kostic, 2024; Katirtzis, 2022; Krishnamurthy, 2021; Torkura et al., 2020; Green, 2023; Frank et al., 2021b; Kesim et al., 2020; Bocetta, 2019; Bremmers, 2021) 12
-Notify Stakeholders Update stakeholders on progress and results.
-Promote transparency and prompt updates.
-(Gill, 2021; Manna, 2021; Dua, 2024; Kostic, 2024; Katirtzis, 2022; Krishnamurthy, 2021; Torkura et al., 2020; Green, 2023; Frank et al., 2021b; Kesim et al., 2020; Bocetta, 2019; Bremmers, 2021) 12
-Maintain Knowledge Repositories Share findings for organizational learning.
-Create a knowledge base to analyze system stress behavior.
-(Gill, 2021; Manna, 2021; Dua, 2024; Kostic, 2024; Katirtzis, 2022; Krishnamurthy, 2021; Torkura et al., 2020; Green, 2023; Frank et al., 2021b; Kesim et al., 2020; Bocetta, 2019; Starr et al., 2022; Vinisky, 2024; Anwar and Arunachalam, 2019; Kadikar, 2023; Patel, 2022; Vu et al., 2022; Zhang et al., 2019, 2023; Rivera et al., 2023; Bremmers, 2021) 21
-Provide Real-Time Dashboards for Visualization Track faults and visualize performance under stress.
-Assess system health for rapid anomaly detection and refinement.
-(Tivey et al., 2019; Kadikar, 2023; Pierce et al., 2021; Zhang et al., 2021a; Wickramasinghe, 2023; Gill, 2022; National Australia Bank, 2020; Simonsson et al., 2021; Al-Said Ahmad et al., 2024; Torkura et al., 2020; Gill, 2021; Kostic, 2024; Fogli et al., 2023; Bedoya et al., 2023; Manna, 2021; Krivas and Portela, 2020; Poltronieri et al., 2022; Camacho et al., 2022; Torkura et al., 2021; Pulcinelli et al., 2023; Chen et al., 2022; Yu et al., 2021; Gianchandani et al., 2022; Palani and Gupta, 2023; Anwar and Arunachalam, 2019; Krishnamurthy, 2021; Roa, 2022; Vinisky, 2024; Siwach et al., 2022; Green, 2023; Hawkins, 2020; Vu et al., 2022; Ma et al., 2023; Rivera et al., 2023) 34
-Analyze Impact and Identify Root Causes of Faults Assess outcomes to measure impact and analyze root causes.
-Enhance resilience, identify weaknesses, and improve stability.
-(Bairyev, 2023; Siwach et al., 2022; Fogli et al., 2023; Zhang et al., 2019; Inc., 2023; Chen et al., 2024; Bhaskar, 2022; Al-Said Ahmad et al., 2024; Gill, 2022; Wickramasinghe, 2023; Yu et al., 2021; Ji et al., 2023; Gianchandani et al., 2022; Nikolaidis et al., 2021; Rivera et al., 2023; Konstantinou et al., 2021; Torkura et al., 2020) 17
-Implement Safety Controls and Automated Rollbacks Set up safety mechanisms, automated rollbacks, and abort thresholds to prevent severe disruptions.
-Minimizes risk, ensures safe experimentation, and reduces downtime with quick recovery measures.
-(Wickramasinghe, 2023; Siwach et al., 2022; Bocetta, 2019; Gill, 2022; Fawcett, 2020; Manna, 2021; Gianchandani et al., 2022; Bairyev, 2023; Malik et al., 2023; Inc., 2023; Ikeuchi et al., 2020; Ma et al., 2023; Nikolaidis et al., 2023; Bhaskar, 2022; Simonsson et al., 2021; Gill, 2021; Krishnamurthy, 2021; Torkura et al., 2020; Katirtzis, 2022; Bedoya et al., 2023; Kadikar, 2023; Patel, 2022; Soldani and Brogi, 2021; Frank et al., 2021b; Rivera et al., 2023; Chen et al., 2024; Zhang et al., 2021a, 2019; Chen et al., 2022; Basiri et al., 2019; Bremmers, 2021; Jernberg et al., 2020; Kostic, 2024) 31
-Ensure Regulatory Compliance Align chaos experiments with regulatory standards and monitor compliance.
-Reduce non-compliance risk and build trust by adhering to industry standards.
-(Wickramasinghe, 2023; Siwach et al., 2022; Bocetta, 2019; Gill, 2022; Fawcett, 2020; Manna, 2021; Gianchandani et al., 2022; Bairyev, 2023; Malik et al., 2023; Inc., 2023; Ikeuchi et al., 2020; Ma et al., 2023; Nikolaidis et al., 2023; Bhaskar, 2022; Simonsson et al., 2021; Gill, 2021; Jernberg et al., 2020; Kadikar, 2023; Kesim et al., 2020; Zhang et al., 2019; Krivas and Portela, 2020; Sbiai et al., 2023; Mooney, 2023; Dua, 2024; Torkura et al., 2021; Palani and Gupta, 2023; Pulcinelli et al., 2023; Kostic, 2024) 28
-Integrate with CI/CD and Third-Party Tools Embed chaos experiments in CI/CD with monitoring integration.
-Automate resilience testing in the software lifecycle to boost efficiency.
-(Manna, 2021; Azure Readiness, 2023; Konstantinou et al., 2021; Palani and Gupta, 2023; Kostic, 2024; Torkura et al., 2021) 6
-Automate Repetitive and Manual Tasks Automate setup and cleanup of chaos experiments using scripts and tools.
-Reduce time and errors by automating repetitive tasks for consistency and efficiency.
-(Camacho et al., 2022; Kadikar, 2023; Miles, 2019a; Green, 2023; Nikolaidis et al., 2021; Poltronieri et al., 2022; Fogli et al., 2023; Torkura et al., 2020; Gill, 2021; Cotroneo et al., 2022; Patel, 2022; Bairyev, 2023; Colyer, 2019; Manna, 2021; Zhang et al., 2019; Kostic, 2024; Fawcett, 2020; Jakkaraju, 2020; Zhang et al., 2021a; Al-Said Ahmad et al., 2024; Mooney, 2023; Kesim et al., 2020; Zhang et al., 2023; Jernberg et al., 2020; Bremmers, 2021; Torkura et al., 2021; Vinisky, 2024; Chen et al., 2022; Vu et al., 2022) 29
+Furthermore, automating tasks and monitoring systems in real-time can minimize disruptions and optimize system performance and user experience.
 6. Chaos Engineering Framework: Taxonomy, Tools, Practices, and Evaluation (RQ3)
-This section presents a taxonomy for chaos engineering and compares existing tools using the taxonomy dimensions. We also present the best practices for adopting chaos engineering in organizations.
-5.1. Taxonomy of Chaos Engineering
-Figure 9 shows the proposed taxonomy, which organizes the capabilities of a chaos engineering platform across several key dimensions: execution environment, automation strategy, automation mode, deployment type, and evaluation approach.
+This section addresses (RQ3) by outlining proposed solutions in chaos engineering.
+It presents a taxonomy of tools and techniques (RQ3.1), highlights best and bad adoption practices (RQ3.2), and describes the methods used to evaluate chaos engineering approaches (RQ3.3).
+6.1. Taxonomy of Chaos Engineering (RQ3.1)
+To address RQ3.1, we propose a taxonomy (Figure 8) that organizes chaos engineering platforms across several key dimensions: execution environment, automation strategy, automation mode, deployment type, and evaluation approach.
+The dimensions of this taxonomy were developed from the core themes identified through our coding and thematic analysis.
+Table 12 presents an excerpt of the intermediate coding structure that grounds this taxonomy, mapping each conceptual category to its subcodes and supporting sources.
 This taxonomy provides organizations with a structured approach to adopting chaos engineering practices by offering clear guidelines for selecting appropriate execution environments, choosing automation strategies that integrate essential tools such as monitoring, logging, and observability, and determining the optimal deployment types (e.g., production or pre-production) based on their risk tolerance.
 It also helps define task modes (manual, semi-automated, or fully automated).
-It guides the selection of fault injection tools, ensuring that organizations have the necessary evaluation methods to measure the impact of experiments and derive actionable insights for system improvement.
+Additionally, it guides the selection of chaos engineering tools by aligning them with specific combinations of these dimensions—for example, tools differ in aspects such as the types of faults they inject and the level of automation they support, ensuring that organizations have the necessary evaluation methods to measure the impact of experiments and derive actionable insights for system improvement.
+Figure 8. A Taxonomy of Chaos Engineering Platforms and Their Capabilities.
+[…]
+A taxonomy of chaos engineering platforms.
+Table 12. Representative Codes and Subcodes for the Taxonomy of Chaos Engineering Platforms and Their Capabilities.
+Code Sub-code Sub-code
+Execution Environment Virtual Machines Variability introduced by different guest OS configurations.
+Resource contention issues examined under CPU/memory saturation scenarios.
+Automation Mode Manual Tends to be time and resource intensive, as each step must be performed and monitored manually.
+Experiments are custom designed for specific systems or hypotheses.
+Carries risk of human error, especially in timing, targeting, or recovery.
+Automation Mode Semi-Automated Engineer-Defined Parameters.
+Human-in-the-Loop Oversight.
+Automation Strategy Custom Code Used to simulate specific failure types (e.g., service crash, DB outage).
+Custom scripts extend existing chaos tools to support advanced scenarios.
+Allows precise control over failure timing, scope, and recovery validation.
+Automation Strategy Prioritization Critical Component Identification.
+Resource Allocation Strategy.
+Automation Strategy Workflow Automation Ensures repeatability and reduces manual error.
+Supports multi-step or multi-system coordination.
+Deployment Type Production Live Environment Testing.
+Real-User Interaction Observation.
+Deployment Type Pre-production Complement to Production Testing.
+Staging Environment Simulation.
 6.1.1. Execution Environment
@@
 Each type of endpoint presents unique challenges that chaos engineering addresses by simulating failure conditions to improve the robustness of the system and ensure continuous operation.
-Virtual machines (VMs) offer a layer of abstraction over physical hardware, allowing multiple independent operating systems to share resources efficiently (Sharma et al., 2016; Smith and Nair, 2005).
-However, VMs can experience performance issues when resources are overutilized, and they are vulnerable to hypervisor failures and migration difficulties (Azure Readiness, 2023; Fogli et al., 2023).
+Virtual machines (VMs) offer a layer of abstraction over physical hardware, allowing multiple independent operating systems to share resources efficiently (Sharma et al., 2016; Smith and Nair, 2005; Basiri et al., 2016).
+However, VMs can experience performance issues when resources are overutilized, and they are vulnerable to hypervisor failures and migration difficulties (Azure Readiness, 2023; Fogli et al., 2023; Sousa et al., 2018).
 Chaos engineering addresses these challenges by testing VMs under stress conditions such as CPU or memory exhaustion to assess if VMs can recover from failures or seamlessly migrate workloads (Kostic, 2024; Chen et al., 2024; Green, 2023).
-Containers, which encapsulate applications and their dependencies in lightweight environments, provide greater efficiency and portability compared to VMs (Sharma et al., 2016; Koskinen et al., 2019).
-However, containers face resource contention, network instability, and orchestration failures, especially in platforms such as Docker and Kubernetes (Watada et al., 2019).
-Chaos engineering tests on containers simulate abrupt stoppages and network disruptions, helping to evaluate containers’ ability to recover within container orchestration systems while maintaining application performance  (Simonsson et al., 2021; Camacho et al., 2022; Bremmers, 2021; Kostic, 2024; Ikeuchi et al., 2020).
-Serverless functions allow applications to be triggered by specific events and run in isolated containers without requiring infrastructure management (Castro et al., 2017; Taibi et al., 2020).
-While serverless architectures are scalable and cost-effective, they face challenges such as resource limitations and the reliance on third-party cloud providers (Rajan, 2020; van Eyk and Iosup, 2018; Kadikar, 2023; Camacho et al., 2022).
-Chaos engineering helps ensure that serverless functions can handle these challenges by simulating cold starts and execution failures, verifying that functions recover quickly and maintain performance under varying loads (Starr et al., 2022; Jernberg et al., 2020; Gill, 2022; Bedoya et al., 2023).
+Containers, which encapsulate applications and their dependencies in lightweight environments, provide greater efficiency and portability compared to VMs (Koskinen et al., 2019; Bedoya et al., 2023).
+However, containers face resource contention, network instability, and orchestration failures, especially in platforms such as Docker and Kubernetes (Simonsson et al., 2021).
+Chaos engineering tests on containers simulate abrupt stoppages and network disruptions, helping to evaluate containers’ ability to recover within container orchestration systems while maintaining application performance  (Simonsson et al., 2021; Camacho et al., 2022; Bremmers, 2021; Kostic, 2024; Ikeuchi et al., 2020; Bedoya et al., 2023).
+Serverless functions allow applications to be triggered by specific events and run in isolated containers without requiring infrastructure management (Castro et al., 2017; Taibi et al., 2020; Al-Said Ahmad et al., 2024; Aharon, 2024).
+While serverless architectures are scalable and cost-effective, they face challenges such as resource limitations and the reliance on third-party cloud providers (Rajan, 2020; van Eyk and Iosup, 2018; Nookala, 2023).
+Chaos engineering helps ensure that serverless functions can handle these challenges by simulating cold starts and execution failures, verifying that functions recover quickly and maintain performance under varying loads (Starr et al., 2022; Jernberg et al., 2020; Bedoya et al., 2023; Morichetta et al., 2023; Chatterjee, 2023).
 Cloud
@@
 Each cloud model has unique characteristics, risks, and opportunities, where chaos engineering can address potential failures.
-Public cloud platforms, such as AWS, Microsoft Azure, and Google Cloud, offer computing resources over the Internet that are shared between multiple organizations (Saraswat and Tripathi, 2020).
+Public cloud platforms, such as AWS, Microsoft Azure, and Google Cloud, offer computing resources over the Internet that are shared between multiple organizations (Saraswat and Tripathi, 2020; Torkura et al., 2020).
 This model enables businesses to scale their infrastructure on demand, eliminating the need for significant investments in physical hardware.
-Due to its cost efficiency and scalability, the public cloud has become a popular choice for many enterprises (Dutta and Dutta, 2019).
+Due to its cost efficiency and scalability, the public cloud has become a popular choice for many enterprises (Dutta and Dutta, 2019; Jakkaraju, 2020; Colyer, 2019).
 However, relying on shared infrastructure introduces specific vulnerabilities.
 For example, a regional outage or failure in an availability zone can have widespread impacts, simultaneously affecting many customers (Paavolainen, 2016).
-In addition, many organizations depend on third-party services, such as application programming interfaces (APIs) or managed databases, and failures in these services can trigger further disruptions in their systems (Zhou et al., 2010).
-To address these risks, chaos engineering can simulate large-scale failures, such as regional outages or availability zone failures, to test how well systems can failover to other regions without significant service interruptions (Fogli et al., 2023).
-Moreover, by injecting failures into critical third-party services, chaos engineering enables organizations to evaluate the effectiveness of their recovery strategies, ensuring that operations can continue even in the face of service disruptions (Ji et al., 2023; Yu et al., 2021).
+In addition, many organizations depend on third-party services, such as application programming interfaces (APIs) or managed databases, and failures in these services can trigger further disruptions in their systems (Zhou et al., 2010; Krishnamurthy, 2021).
+To address these risks, chaos engineering can simulate large-scale failures, such as regional outages or availability zone failures, to test how well systems can failover to other regions without significant service interruptions (Fogli et al., 2023; National Australia Bank, 2020).
+Moreover, by injecting failures into critical third-party services, chaos engineering enables organizations to evaluate the effectiveness of their recovery strategies, ensuring that operations can continue even in the face of service disruptions (Yu et al., 2021; Hawkins, 2020; Vinisky, 2024).
 Private cloud environments, which provide dedicated infrastructure exclusively for a single organization, can be hosted on-premises or managed by third-party providers, offering greater control over data privacy, security, and customization (Jadeja and Modi, 2012; Gastermann et al., 2015).
 This model suits organizations with strict regulatory requirements or managing sensitive information (Pearson, 2013).
-Despite their benefits, private clouds often face scalability and geographic redundancy challenges, as they typically lack the built-in redundancy of public clouds, making them more susceptible to service interruptions due to hardware failures or resource constraints (Kumar and Goudar, 2012).
+Despite their benefits, private clouds often face scalability and geographic redundancy challenges, as they typically lack the built-in redundancy of public clouds, making them more susceptible to service interruptions due to hardware failures or resource constraints (Kumar and Goudar, 2012; Kassab et al., 2022).
 Without adequate failover mechanisms, private cloud environments may struggle to maintain operational continuity during hardware or network failures (Rimal et al., 2009; Tivey et al., 2019).
-Chaos engineering can be used in private clouds to simulate hardware failures, such as server crashes or network partitions, to evaluate the system’s ability to recover from localized disruptions (Patel, 2022).
-By simulating resource exhaustion, such as CPU or memory overloads, organizations can test whether their infrastructure can scale to meet unexpected surges in demand, ensuring continuous service availability under pressure (Gill, 2021).
-Finally, hybrid cloud environments combine the strengths of public and private clouds, allowing organizations to manage sensitive workloads on private infrastructure while leveraging the scalability of public clouds for nonsensitive or high-demand tasks (Goyal, 2014).
+Chaos engineering can be used in private clouds to simulate hardware failures, such as server crashes or network partitions, to evaluate the system’s ability to recover from localized disruptions (Patel, 2022; Simonsson et al., 2021).
+By simulating resource exhaustion, such as CPU or memory overloads, organizations can test whether their infrastructure can scale to meet unexpected surges in demand, ensuring continuous service availability under pressure (Gill, 2021; Jernberg et al., 2020).
+Finally, hybrid cloud environments combine the strengths of public and private clouds, allowing organizations to manage sensitive workloads on private infrastructure while leveraging the scalability of public clouds for nonsensitive or high-demand tasks (Goyal, 2014; Poltronieri et al., 2022).
 This approach provides flexibility but introduces complexity in maintaining smooth integration between the two environments (Lu et al., 2014).
-A key challenge of hybrid clouds is to ensure seamless communication between public and private components, as issues such as network latency, connectivity problems, and data synchronization delays can arise when workloads are distributed across both environments, potentially leading to service disruptions (Annapureddy, 2010).
+A key challenge of hybrid clouds is to ensure seamless communication between public and private components, as issues such as network latency, connectivity problems, and data synchronization delays can arise when workloads are distributed across both environments, potentially leading to service disruptions (Annapureddy, 2010; Siwach et al., 2022).
 Chaos engineering can address this issue by allowing organizations to simulate network interruptions or increased latency between public and private clouds (Fogli et al., 2023).
-These tests help to evaluate whether workloads can continue to operate despite connectivity issues (Torkura et al., 2020).
+These tests help to evaluate whether workloads can continue to operate despite connectivity issues (Torkura et al., 2020; Dedousis et al., 2023).
 Furthermore, chaos experiments that simulate delays or failures in data replication help ensure that the system maintains data consistency and operational continuity, even when part of the infrastructure is compromised (Camacho et al., 2022).
@@
 Similarly, scaling limitations can be addressed by testing how an on-premise infrastructure responds to various demand surges, such as a sudden spike in user traffic during an e-commerce company’s high-demand event (Azure Readiness, 2023; Gianchandani et al., 2022).
-This helps identify bottlenecks, optimize resource allocation, and ensure that the system remains responsive under varying load conditions (Naqvi et al., 2022; Wu, 2021; Sbiai et al., 2023; Wickramasinghe, 2023).
+This helps identify bottlenecks, optimize resource allocation, and ensure that the system remains responsive under varying load conditions (Naqvi et al., 2022; Wickramasinghe, 2023; Camacho et al., 2022; Nikolaidis et al., 2023, 2021).
 6.1.2. Experimentation Modes
 Chaos engineering platforms offer three modes of experimentation, namely manual, semi-automated, and fully automated, each tailored to different levels of control, complexity, and system scale.
-Manual experimentation is most appropriate when engineers require high precision and control, as they take full responsibility for designing, configuring, and executing chaos tests (Kostic, 2024).
-This approach is suited for organizations with little or no experience in chaos engineering or when highly specific, customized experiments are necessary (Durai, 2022).
+Manual experimentation is most appropriate when engineers require high precision and control, as they take full responsibility for designing, configuring, and executing chaos tests (Kostic, 2024; Torkura et al., 2020).
+This approach is suited for organizations with little or no experience in chaos engineering or when highly specific, customized experiments are necessary (Durai, 2022; Zhang et al., 2019).
 For example, an engineer might manually trigger a database failure or introduce network latency to observe the system response in real time (Jernberg et al., 2020; Bremmers, 2021).
 While this mode allows for fine-tuned control, it can be time-consuming and prone to human error, making it ideal for small-scale, customized tests or critical systems where close monitoring is essential (Ikeuchi et al., 2020; Chen et al., 2024).
-Semi-automated experimentation provides a middle ground between manual control and complete automation, where engineers define the experiment parameters (Sondhi et al., 2021; Jernberg et al., 2020).
+Semi-automated experimentation provides a middle ground between manual control and complete automation, where engineers define the experiment parameters (Jernberg et al., 2020; Pierce et al., 2021; Frank et al., 2021b).
 At the same time, the platform automates tasks such as triggering tests and collecting data, making this mode ideal for experiments that benefit from automated execution while still requiring moderate human involvement to reduce manual effort (Kesim et al., 2020; Nikolaidis et al., 2021; Gill, 2021).
 For example, engineers might set up an experiment to test resource exhaustion on specific microservices, with the platform automatically running the test at predefined intervals, collecting performance metrics, and allowing them to focus on adjusting parameters and interpreting results (Pierce et al., 2021; Dedousis et al., 2023).
-Fully automated experimentation suits large-scale systems that require frequent and continuous testing (Zhang et al., 2021a; Simonsson et al., 2021; Basiri et al., 2019).
-After engineers configure the initial setup, the chaos engineering platform takes over, seamlessly integrating experiments into the system operations, typically through the CI/CD pipelines (Konstantinou et al., 2021; Pulcinelli et al., 2023; Manna, 2021; Colyer, 2019; Azure Readiness, 2023).
-For example, after a code deployment or infrastructure update, the platform can automatically simulate a service outage or introduce network disruptions, then autonomously gather performance metrics, identify system weaknesses, and generate detailed reports, requiring engineers to intervene only when necessary (Dedousis et al., 2023; Bairyev, 2023; Kesim et al., 2020; Cotroneo et al., 2022; Nikolaidis et al., 2021; Patel, 2022).
+Fully automated experimentation suits large-scale systems that require frequent and continuous testing (Simonsson et al., 2021; Basiri et al., 2019; Patel, 2022).
+After engineers configure the initial setup, the chaos engineering platform takes over, seamlessly integrating experiments into the system operations, typically through the CI/CD pipelines (Konstantinou et al., 2021; Pulcinelli et al., 2023; Manna, 2021; Colyer, 2019).
+For example, after a code deployment or infrastructure update, the platform can automatically simulate a service outage or introduce network disruptions, then autonomously gather performance metrics, identify system weaknesses, and generate detailed reports, requiring engineers to intervene only when necessary (Dedousis et al., 2023; Bairyev, 2023; Kesim et al., 2020; Cotroneo et al., 2022; Nikolaidis et al., 2021).
 6.1.3. Chaos Automation Strategy
@@
 Custom Code
-This involves developing programs or scripts to extend the capabilities of existing tools and gain precise control over system configurations and experiments (Myers et al., 2000; Thüm et al., 2014; Khankhoje, 2022).
+This involves developing programs or scripts to extend the capabilities of existing tools and gain precise control over system configurations and experiments (Myers et al., 2000; Thüm et al., 2014; Khankhoje, 2022; Cotroneo et al., 2022).
 It allows engineers to simulate unique failure conditions, ranging from simple component failures, such as restarting a single service, to complex multi-step sequences involving multiple systems, tailored to the specific challenges of their architecture that standard tools may not address (Gill, 2021; Inc., 2023; Torkura et al., 2020; Azure Readiness, 2023; Patel, 2022; Nikolaidis et al., 2023).
@@
 In more complex scenarios, custom code can initiate a service outage followed by a database failure or a network partition to observe how these failures impact system recovery and resilience (Wickramasinghe, 2023; Gill, 2021; Simonsson et al., 2021).
-Additionally, custom code can be integrated into CI/CD pipelines to simulate high resource consumption, such as CPU throttling or memory exhaustion, after each deployment, allowing early detection of performance bottlenecks before they affect production (Azure Readiness, 2023; Konstantinou et al., 2021; Manna, 2021; Kadikar, 2023; Palani and Gupta, 2023).
+Additionally, custom code can be integrated into CI/CD pipelines to simulate high resource consumption, such as CPU throttling or memory exhaustion, after each deployment, allowing early detection of performance bottlenecks before they affect production (Azure Readiness, 2023; Konstantinou et al., 2021; Manna, 2021; Kadikar, 2023; Palani and Gupta, 2023a).
 This custom code for chaos experiments can be written in various programming languages, such as Python, Bash, or Go, depending on the complexity and specific requirements of the experiment (Siwach et al., 2022; Malik et al., 2023; Sondhi et al., 2021; Simonsson et al., 2021; Zhang et al., 2021b).
@@
 This automated integration ensures continuous data collection without manual intervention, enabling real-time analysis during system disruptions or failures.
-As a result, teams can monitor performance, identify issues quickly, and respond more efficiently to failures (Palani and Gupta, 2023; Kostic, 2024).
-This process is vital to integrating monitoring tools that track essential system metrics such as CPU usage, memory consumption, and network performance (Abdul, 2024; Wu, 2021; Gianchandani et al., 2022; Zhang et al., 2019).
+As a result, teams can monitor performance, identify issues quickly, and respond more efficiently to failures (Palani and Gupta, 2023a; Kostic, 2024).
+This process is vital to integrating monitoring tools that track essential system metrics such as CPU usage, memory consumption, and network performance (Wu, 2021; Gianchandani et al., 2022; Zhang et al., 2019).
 Monitoring tools can automatically detect performance deviations and provide immediate feedback.
@@
 For example, if network communication between two microservices fails, the logging systems capture the exact moment the failure occurred and show the error messages generated by failed retries (Cotroneo et al., 2022; Ma et al., 2023).
-This data helps teams analyze how the failure affected communication and whether recovery mechanisms worked effectively (Naqvi et al., 2022; Malik et al., 2023; Kostic, 2024; Palani and Gupta, 2023).
+This data helps teams analyze how the failure affected communication and whether recovery mechanisms worked effectively (Naqvi et al., 2022; Malik et al., 2023; Kostic, 2024; Palani and Gupta, 2023a).
 Building on this, observability tools provide a broader view by correlating logs and metrics to show how different components of the system interact during failure scenarios.
@@
 This tracing allows the team to see if minor delays in one service cause issues elsewhere (Durai, 2022; Katirtzis, 2022; Kostic, 2024; Chen et al., 2024; Nikolaidis et al., 2023; Frank et al., 2023).
-Furthermore, alert notification systems can ensure that relevant teams are promptly notified when critical failures occur or performance thresholds are exceeded (Tivey et al., 2019; Gianchandani et al., 2022; Palani and Gupta, 2023; Bremmers, 2021; Starr et al., 2022).
+Furthermore, alert notification systems can ensure that relevant teams are promptly notified when critical failures occur or performance thresholds are exceeded (Tivey et al., 2019; Gianchandani et al., 2022; Palani and Gupta, 2023a; Bremmers, 2021; Starr et al., 2022).
 For example, if a database connection pool reaches its maximum capacity during a high-load chaos test, the alert system triggers a notification to the database operations team, allowing them to address the issue before it causes a system-wide failure.
@@
 Prioritization
-This involves categorizing system components into critical and non-critical groups to focus testing efforts on the areas where failures would have the greatest impact (Dedousis et al., 2023; Anwar and Arunachalam, 2019).
+This involves categorizing system components into critical and non-critical groups to focus testing efforts on the areas where failures would have the greatest impact (Dedousis et al., 2023; Anwar and Arunachalam, 2019; Chen et al., 2024).
 Critical components are essential for maintaining core operations, and their failure would cause significant disruption (Fogli et al., 2023; Kesim et al., 2020).
@@
 By prioritizing critical components, testing can focus on areas where failure would result in the most severe consequences.
-In contrast, the resilience issues of non-critical components can be addressed less urgently, since their failure poses a lower risk to overall system stability (Wickramasinghe, 2023).
+In contrast, the resilience issues of non-critical components can be addressed less urgently, since their failure poses a lower risk to overall system stability (Wickramasinghe, 2023; Inc., 2023).
 6.1.4. Deployment
@@
 Consequently, these tests should be viewed as a complement, rather than a substitute, to production testing (Zhang et al., 2019; Torkura et al., 2020; Gianchandani et al., 2022).
-5.1.5. Evaluation Approaches
+6.2. Chaos Engineering Tools
+To further address RQ3.1, we applied the taxonomy developed in this study to evaluate ten widely used chaos engineering tools (T1–T10 in Table 13).
+We identified 71 tools mentioned across academic and grey sources during our thematic analysis.
+We filtered this list by inspecting their GitHub repositories.
+We excluded tools that had been archived (read-only), were no longer maintained, or lacked sufficient documentation.
+This screening process reduced the set to 41 active tools with publicly available repositories.
+From this refined pool of 41 tools, we selected the ten most starred repositories on GitHub as of November 2024.
+The ‘most stars’ metric was used as a proxy for community interest and real-world relevance  (Borges et al., 2016).
+Each selected tool was evaluated using our proposed taxonomy, which classifies tools based on four key dimensions: execution environments, automation modes, failure injection strategies, and deployment stages.
+This mapping is presented in Table 13 and is intended to help engineering teams align tool selection with the architecture of the systems under test(e.g., microservices, containerized environments) and operational goals (e.g.,reducing latency, enhancing scalability) (Kavaler et al., 2019).
+For example, T1 and T2 support Kubernetes environments.
+T1 enables manual testing in both pre-production and production stages, while T2 extends automation with custom code and integration, targeting pre-production use.
+T3 to T5 offer a broader platform support, including Kubernetes, Docker, and VMs, and provide full automation features such as workflow integration, prioritization, and custom code, applicable across all deployment stages.
+T6 and T7 focus on Kubernetes and containers.
+Both support manual, semi-automated, and fully automated modes; T6 emphasizes hybrid setups with prioritization, while T7 supports Docker and targets production-grade testing.
+T8 and T9 operate across Kubernetes and virtual machines.
+T8 supports complete automation and is usable in both staging and live systems, while T9 focuses on pre-production with advanced integration options.
+Finally, T10 is tailored for hybrid cloud environments, supporting Kubernetes and providing full automation across both pre-production and production environments.
+Table 13. Top-10 Most Used Chaos Engineering Platforms.
+Tool Code Tool Name Execution Environment Automation Mode Automation Strategy Deployment Stages
+Endpoint On- Premise Cloud Manual Semi- Automated Fully Automated Prioritization Workflow Automation Custom Code Auto- Integration Pre- production Production
+Docker Kubernetes
+Virtual
+Machine
+Serverless
+Function
+Private Public Hybrid
+T1 Chaos Monkey
+T2 Toxiproxy
+T3 Chaos Mesh
+T4 ChaosBlade
+T5 LitmusChaos
+T6 Kube-monkey
+T7 Pumba
+T8 Chaos Toolkit
+T9 Powerfulseal
+T10 Chaoskube
+6.3. Adoption Practices (RQ3.2)
+To address RQ3.2, we highlight best practices and common pitfalls in adopting chaos engineering derived from our literature sources.
+Successfully adopting chaos engineering requires a structured approach that balances best practices with avoiding common mistakes.
+To minimize risk during experimentation, organizations should follow well-defined guidelines for planning, executing, and evaluating chaos experiments  (Siwach et al., 2022; Green, 2023).
+Table 14 summarizes essential best practices, such as starting with small, controlled experiments, setting clear objectives, automating tests, and closely monitoring system behavior.
+Additionally, organizations should limit the scope of experiments, encourage cross-team collaboration, and continuously refine processes based on experiment outcomes.
+While following best practices is crucial, organizations must also avoid common pitfalls, such as those outlined in Table 15, which include failing to tailor experiments to specific systems, neglecting proper planning and hypothesis formation, and lacking adequate safety mechanisms.
+By adhering to best practices and avoiding these common pitfalls, organizations can confidently integrate chaos engineering into their systems and achieve successful outcomes (Siwach et al., 2022).
+For full details, see the code-to-theme mapping in the online appendix (Section 3.6).
+Table 14. Best Practices.
+No Best Practices Description Articles Total
+1 Start Small and Scale Up Start with manageable tests, then expand gradually.
+(Fogli et al., 2023; Torkura et al., 2020; Bremmers, 2021; Patel, 2022; Gill, 2021; Chen et al., 2022; Siwach et al., 2022; Jernberg et al., 2020; Anwar and Arunachalam, 2019; Palani and Gupta, 2023a; Kadikar, 2023; Green, 2023) 12
+2 Define Clear Objectives Set precise goals to guide each experiment’s purpose and outcomes.
+(Torkura et al., 2020; Dedousis et al., 2023; Inc., 2023; Patel, 2022; Zhang et al., 2021a; Al-Said Ahmad et al., 2024; Kesim et al., 2020; Soldani and Brogi, 2021; Siwach et al., 2022; Jernberg et al., 2020; Jakkaraju, 2020; Dua, 2024; Kadikar, 2023; Green, 2023) 14
+3 Automate Experiments Use automation tools to run chaos experiments regularly, ensuring consistency and frequency.
+(Torkura et al., 2020; Zhang et al., 2019; Inc., 2023; Gill, 2021; Wickramasinghe, 2023; Camacho et al., 2022; Zhang et al., 2021a; Al-Said Ahmad et al., 2024; Vu et al., 2022; Kassab et al., 2022; Ikeuchi et al., 2020; Naqvi et al., 2022; Malik et al., 2023; Soldani and Brogi, 2021; Siwach et al., 2022; Jernberg et al., 2020; Bhaskar, 2022; Gill, 2022; Fawcett, 2020; Bairyev, 2023; Sbiai et al., 2023; Azure Readiness, 2023; Roa, 2022; Green, 2023) 24
+4 Track and Measure Results Monitor the system’s performance during experiments and collect data to see the impact.
+(Fogli et al., 2023; Torkura et al., 2020; Dedousis et al., 2023; Inc., 2023; Bremmers, 2021; Patel, 2022; Gill, 2021; Camacho et al., 2022; Torkura et al., 2021; Zhang et al., 2021a; Vu et al., 2022; Kassab et al., 2022; Ikeuchi et al., 2020; Malik et al., 2023; Soldani and Brogi, 2021; Siwach et al., 2022; Konstantinou et al., 2021; Jernberg et al., 2020; Hawkins, 2020; Manna, 2021; Azure Readiness, 2023; Kadikar, 2023; Green, 2023) 23
+5 Involve all Key Stakeholders Include team members to ensure everyone is informed and aligned.
+(Torkura et al., 2020; Inc., 2023; Bremmers, 2021; Patel, 2022; Gill, 2021; Frank et al., 2021b; Siwach et al., 2022; Dua, 2024; Kadikar, 2023; Green, 2023; Gill, 2022; Kostic, 2024) 12
+6 Document and Improve Record experiments to continually improve the system.
+(Torkura et al., 2020; Zhang et al., 2019; Inc., 2023; Bremmers, 2021; Patel, 2022; Gill, 2021; Siwach et al., 2022; Palani and Gupta, 2023a; Kadikar, 2023; Green, 2023; Kesim et al., 2020) 11
+Table 15. Bad Practices.
+No Bad Practices Description Articles Total
+1 Misaligned Chaos Chaos experiments that don’t match the system yield misleading results.
+(Fogli et al., 2023; Torkura et al., 2020; Zhang et al., 2019; Dedousis et al., 2023; Vu et al., 2022; Nikolaidis et al., 2023; Ma et al., 2023; Sondhi et al., 2021; Jernberg et al., 2020; Konstantinou et al., 2021; Park et al., 2019; Zhang et al., 2023; Rivera et al., 2023; Patel, 2022; Le, 2022; Lardo, 2019; Hochstein and Rosenthal, 2016; Basiri et al., 2016) 18
+2 Inadequate Experimentation Without planning, clear goals, and baseline tracking, assessing chaos test impacts becomes difficult.
+(Inc., 2023; Wickramasinghe, 2023; Gill, 2021; Bremmers, 2021; Jakkaraju, 2020; Le, 2022; Hawkins, 2020; Anwar and Arunachalam, 2019; Dua, 2024; National Australia Bank, 2020; Mooney, 2023; Starr et al., 2022; Palani and Gupta, 2023a; Gill, 2022; Fawcett, 2020; Gianchandani et al., 2022; Bairyev, 2023; Abdul, 2024; Azure Readiness, 2023; Kadikar, 2023; Green, 2023; Aharon, 2024; Krivas and Portela, 2020; Kostic, 2024; Frank et al., 2021b; Jernberg et al., 2020; Konstantinou et al., 2021; Ji et al., 2023; Soldani and Brogi, 2021; Park et al., 2019; Wachtel, 2023; Alvaro and Tymon, 2017; Sousa et al., 2018) 33
+3 Safety Mechanism Oversight Not testing rollback and safety features risks uncontrolled failures.
+(Inc., 2023; Wickramasinghe, 2023; Gill, 2021; Bremmers, 2021; Patel, 2022; Jakkaraju, 2020; Bhaskar, 2022; Katirtzis, 2022; Dua, 2024; Starr et al., 2022; Tivey et al., 2019; Palani and Gupta, 2023a; Bairyev, 2023; Sbiai et al., 2023; Kadikar, 2023; Green, 2023; Vinisky, 2024; Pierce et al., 2021; Al-Said Ahmad et al., 2024; Vu et al., 2022; Ikeuchi et al., 2020; Chen et al., 2022; Naqvi et al., 2022; Soldani and Brogi, 2021; Konstantinou et al., 2021; Alvaro and Tymon, 2017; Sousa et al., 2018) 27
+4 Insufficient Tooling Lacking observability, automation, and integration tools makes monitoring and analyzing chaos experiments difficult.
+(Inc., 2023; Wickramasinghe, 2023; Gill, 2021; Hawkins, 2020; Starr et al., 2022; National Australia Bank, 2020; Tomka, 2024; Dua, 2024; Mooney, 2023; Lardo, 2019; Gianchandani et al., 2022; Abdul, 2024; Azure Readiness, 2023; Roa, 2022; Kadikar, 2023; Green, 2023; Krivas and Portela, 2020; Kostic, 2024; Simonsson et al., 2021; Nikolaidis et al., 2021; Soldani and Brogi, 2021; Al-Said Ahmad et al., 2024; Vu et al., 2022; Naqvi et al., 2022; Malik et al., 2023; Camilli et al., 2022; Bremmers, 2021; Patel, 2022; Le, 2022; Zhang et al., 2021a; Cotroneo et al., 2022; Ikeuchi et al., 2020; Bedoya et al., 2023; Basiri et al., 2019; Pulcinelli et al., 2023; Jernberg et al., 2020; Anwar and Arunachalam, 2019; Durai, 2022; Jakkaraju, 2020) 39
+5 Poor Stakeholder Collaboration Excluding key team members causes misalignment and resistance.
+(Bremmers, 2021; Starr et al., 2022; Le, 2022; Krishnamurthy, 2021; Katirtzis, 2022; Durai, 2022; Tivey et al., 2019; Gill, 2022; Green, 2023; Poltronieri et al., 2022; Frank et al., 2021b; Al-Said Ahmad et al., 2024; Jernberg et al., 2020; Kesim et al., 2020) 14
+6.4. Evaluation Approaches (RQ3.3)
 Evaluation refers to determining the effectiveness of chaos engineering experiments by measuring their impact on system stability, the system’s ability to handle induced failures, and the growth and maturity of the organization’s chaos engineering practices.
-Evaluation approaches can be categorized into qualitative and quantitative methods.
+To address RQ3.3, we categorized evaluation approaches into quantitative and qualitative methods based on themes derived from our systematic coding and thematic analysis (see Section 3.5 and detailed mappings in the online appendix 3.6).
 Quantitative Evaluation
@@
   • Business Impact Evaluation assesses how chaos experiments affect key business metrics, including revenue, customer churn, and operational efficiency (Inc., 2023; Jernberg et al., 2020).
-This involves understanding specific financial losses, customer complaints, and service-level breaches resulting from experiments(National Australia Bank, 2020; Palani and Gupta, 2023).
+This involves understanding specific financial losses, customer complaints, and service-level breaches resulting from experiments(National Australia Bank, 2020; Palani and Gupta, 2023a).
 For instance, a chaos experiment that causes downtime during a product launch could result in significant lost sales and a spike in customer service tickets (Ji et al., 2023; Starr et al., 2022).
@@
 This evaluation is essential to identify operational weaknesses in real-time incident management and to improve team workflows and readiness for actual outages (Kesim et al., 2020; Wu, 2021; Fawcett, 2020).
-5.2. Chaos Engineering Tools
-We used the proposed taxonomy to evaluate 10 popular chaos engineering tools (T1-T10 in Table 10).
-We selected these tools by identifying the most-cited tools from our pool of reviewed papers, then ranking them using GitHub metrics, precisely the ’Most stars’ measure, to confirm their popularity and community engagement (Borges et al., 2016).
-Collectively, these tools had over 93,000 stars as of November 2024.
-These tools have been mapped according to their applicability to different execution environments, automation modes, strategies, and deployment stages.
-This categorization enables practitioners to make well-informed choices by aligning tool selection with specific infrastructure needs, operational requirements, automation capabilities, and system risk tolerance.
-Consequently, selecting the appropriate tool in a software development project becomes a crucial decision (Kavaler et al., 2019).
-For example, T1 and T2 provide manual and semi-automated chaos testing capabilities, supporting Kubernetes, Docker, and virtual machine setups.
-While T1 supports performing chaos experiments in live production environments, T2 focuses on testing environments, particularly pre-production setups.
-T3 is a retired tool designed initially for AWS but lacks support for Kubernetes.
-Its functions have been migrated to newer Netflix tools such as Chaos Monkey, Swabbie, and Spinnaker, making it better suited to traditional infrastructures.
-T4 and T5 both support Kubernetes and virtual machine environments in hybrid clouds.
-These tools offer full automation options, with features like workflow automation and custom integrations, making them ideal for chaos testing in pre-production and production environments.
-T6 and T7 focus on Kubernetes and serverless functions, offering manual, semi-automated, and fully automated modes.
-While T6 is optimized for hybrid cloud setups with advanced workflow customization, T7 is better suited for production resilience testing in Kubernetes environments.
-T8 and T9 emphasize adaptability across diverse infrastructures.
-T8 supports Kubernetes, serverless functions, and on-premise setups, focusing on manual testing modes with customizable workflows.
-T9, on the other hand, operates across virtual machines and Kubernetes environments, providing semi-automated and fully automated capabilities with workflow automation and custom integration options, making it suitable for public cloud systems.
-Finally, T10 excels in hybrid cloud environments, supporting Kubernetes and virtual machines with advanced automation and integration capabilities for pre-production and production stages.
-This taxonomy ensures that chaos engineering tools are mapped to their ideal use cases, helping practitioners choose the best tools for their systems.
-Our online appendix provides the complete list of tools considered.
-Figure 9. A Taxonomy of Chaos Engineering Platforms.
-[…]
-A taxonomy of chaos engineering platforms.
-Table 10. Top-10 Most Used Chaos Engineering Platforms.
-Tool Code Tool Name Execution Environment Automation Mode Automation Strategy Deployment Stages
-Endpoint On- Premise Cloud Manual Semi- Automated Fully Automated Prioritization Workflow Automation Custom Code Auto- Integration Pre- production Production
-Docker Kubernetes
-Virtual
-Machine
-Serverless
-Function
-Private Public Hybrid
-T1 Chaos Monkey
-T2 Toxiproxy
-T3 Simian Army
-T4 Chaos Mesh
-T5 Chaosblade
-T6 Litmus
-T7 Pumba
-T8 Chaos Toolkit
-T9 Powerful seal
-T10 Chaoskube
-5.3. Adoption Practices
-Adopting chaos engineering requires a structured approach, balancing best practices with avoiding common mistakes.
-To effectively implement chaos experiments and reduce the risk of system failures or disruptions, organizations should follow well-established guidelines for planning, executing, and evaluating these experiments  (Siwach et al., 2022; Green, 2023).
-Table 11 summarizes essential best practices, such as starting with small, controlled experiments, setting clear objectives, automating tests, and closely monitoring system behavior.
-Additionally, organizations should limit the scope of experiments, encourage cross-team collaboration, and continuously refine processes based on experiment outcomes.
-While following best practices is crucial, organizations must also avoid common pitfalls, such as those outlined in Table 12, which include failing to tailor experiments to specific systems, neglecting proper planning and hypothesis formation, and lacking adequate safety mechanisms.
-By adhering to best practices and avoiding these common pitfalls, organizations can confidently integrate chaos engineering into their systems and achieve successful outcomes (Siwach et al., 2022).
-Table 11. Best Practices
-No Best Practices Description Articles Total
-1 Start Small and Scale Up Start with manageable tests, then expand gradually.
-(Fogli et al., 2023; Torkura et al., 2020; Bremmers, 2021; Patel, 2022; Gill, 2021; Chen et al., 2022; Siwach et al., 2022; Jernberg et al., 2020; Anwar and Arunachalam, 2019; Palani and Gupta, 2023; Kadikar, 2023; Green, 2023) 12
-2 Define Clear Objectives Set precise goals to guide each experiment’s purpose and outcomes.
-(Torkura et al., 2020; Dedousis et al., 2023; Inc., 2023; Patel, 2022; Zhang et al., 2021a; Al-Said Ahmad et al., 2024; Kesim et al., 2020; Soldani and Brogi, 2021; Siwach et al., 2022; Jernberg et al., 2020; Jakkaraju, 2020; Dua, 2024; Kadikar, 2023; Green, 2023) 14
-3 Automate Experiments Use automation tools to run chaos experiments regularly, ensuring consistency and frequency.
-(Torkura et al., 2020; Zhang et al., 2019; Inc., 2023; Gill, 2021; Wickramasinghe, 2023; Camacho et al., 2022; Zhang et al., 2021a; Al-Said Ahmad et al., 2024; Vu et al., 2022; Kassab et al., 2022; Ikeuchi et al., 2020; Naqvi et al., 2022; Malik et al., 2023; Soldani and Brogi, 2021; Siwach et al., 2022; Jernberg et al., 2020; Bhaskar, 2022; Gill, 2022; Fawcett, 2020; Bairyev, 2023; Sbiai et al., 2023; Azure Readiness, 2023; Roa, 2022; Green, 2023) 24
-4 Track and Measure Results Monitor the system’s performance during experiments and collect data to see the impact.
-(Fogli et al., 2023; Torkura et al., 2020; Dedousis et al., 2023; Inc., 2023; Bremmers, 2021; Patel, 2022; Gill, 2021; Camacho et al., 2022; Torkura et al., 2021; Zhang et al., 2021a; Vu et al., 2022; Kassab et al., 2022; Ikeuchi et al., 2020; Malik et al., 2023; Soldani and Brogi, 2021; Siwach et al., 2022; Konstantinou et al., 2021; Jernberg et al., 2020; Hawkins, 2020; Manna, 2021; Azure Readiness, 2023; Kadikar, 2023; Green, 2023) 23
-5 Involve all Key Stakeholders Include team members to ensure everyone is informed and aligned.
-(Torkura et al., 2020; Inc., 2023; Bremmers, 2021; Patel, 2022; Gill, 2021; Frank et al., 2021b; Siwach et al., 2022; Dua, 2024; Kadikar, 2023; Green, 2023; Gill, 2022; Kostic, 2024) 12
-6 Document and Improve Record experiments to continually improve the system.
-(Torkura et al., 2020; Zhang et al., 2019; Inc., 2023; Bremmers, 2021; Patel, 2022; Gill, 2021; Siwach et al., 2022; Palani and Gupta, 2023; Kadikar, 2023; Green, 2023; Kesim et al., 2020) 11
-Table 12. Bad Practices
-No Bad Practices Description Articles Total
-1 Misaligned Chaos Chaos experiments that don’t match the system yield misleading results.
-(Fogli et al., 2023; Torkura et al., 2020; Zhang et al., 2019; Dedousis et al., 2023; Vu et al., 2022; Nikolaidis et al., 2023; Ma et al., 2023; Sondhi et al., 2021; Jernberg et al., 2020; Konstantinou et al., 2021; Park et al., 2019; Zhang et al., 2023; Rivera et al., 2023; Patel, 2022; Le, 2022; Lardo, 2019) 16
-2 Inadequate Experimentation Without planning, clear goals, and baseline tracking, assessing chaos test impacts becomes difficult.
-(Inc., 2023; Wickramasinghe, 2023; Gill, 2021; Bremmers, 2021; Jakkaraju, 2020; Le, 2022; Hawkins, 2020; Anwar and Arunachalam, 2019; Dua, 2024; National Australia Bank, 2020; Mooney, 2023; Starr et al., 2022; Palani and Gupta, 2023; Gill, 2022; Fawcett, 2020; Gianchandani et al., 2022; Bairyev, 2023; Abdul, 2024; Azure Readiness, 2023; Kadikar, 2023; Green, 2023; Aharon, 2024; Krivas and Portela, 2020; Kostic, 2024; Frank et al., 2021b; Jernberg et al., 2020; Konstantinou et al., 2021; Ji et al., 2023; Soldani and Brogi, 2021; Park et al., 2019; Wachtel, 2023) 31
-3 Safety Mechanism Oversight Not testing rollback and safety features risks uncontrolled failures.
-(Inc., 2023; Wickramasinghe, 2023; Gill, 2021; Bremmers, 2021; Patel, 2022; Jakkaraju, 2020; Bhaskar, 2022; Katirtzis, 2022; Dua, 2024; Starr et al., 2022; Tivey et al., 2019; Palani and Gupta, 2023; Bairyev, 2023; Sbiai et al., 2023; Kadikar, 2023; Green, 2023; Vinisky, 2024; Pierce et al., 2021; Al-Said Ahmad et al., 2024; Vu et al., 2022; Ikeuchi et al., 2020; Chen et al., 2022; Naqvi et al., 2022; Soldani and Brogi, 2021; Konstantinou et al., 2021) 25
-4 Insufficient Tooling Lacking observability, automation, and integration tools makes monitoring and analyzing chaos experiments difficult.
-(Inc., 2023; Wickramasinghe, 2023; Gill, 2021; Hawkins, 2020; Starr et al., 2022; National Australia Bank, 2020; Tomka, 2024; Dua, 2024; Mooney, 2023; Lardo, 2019; Gianchandani et al., 2022; Abdul, 2024; Azure Readiness, 2023; Roa, 2022; Kadikar, 2023; Green, 2023; Krivas and Portela, 2020; Kostic, 2024; Simonsson et al., 2021; Nikolaidis et al., 2021; Soldani and Brogi, 2021; Al-Said Ahmad et al., 2024; Vu et al., 2022; Naqvi et al., 2022; Malik et al., 2023; Camilli et al., 2022; Bremmers, 2021; Patel, 2022; Le, 2022; Zhang et al., 2021a; Cotroneo et al., 2022; Ikeuchi et al., 2020; Bedoya et al., 2023; Basiri et al., 2019; Pulcinelli et al., 2023; Jernberg et al., 2020; Anwar and Arunachalam, 2019; Durai, 2022; Jakkaraju, 2020) 39
-5 Poor Stakeholder Collaboration Excluding key team members causes misalignment and resistance.
-(Bremmers, 2021; Starr et al., 2022; Le, 2022; Krishnamurthy, 2021; Katirtzis, 2022; Durai, 2022; Tivey et al., 2019; Gill, 2022; Green, 2023; Poltronieri et al., 2022; Frank et al., 2021b; Al-Said Ahmad et al., 2024; Jernberg et al., 2020; Kesim et al., 2020) 14
 7. Current State of Chaos Engineering (RQ4)
-This section first presents the classification of research in chaos engineering, then examines the number of sources published per year and venue type, and finally reviews the venue rankings and key contributors in the field.
-6.1. Number of Sources by Research Type
-To better understand the current state of chaos engineering research, we classified the primary studies based on the research classification framework proposed by Wieringa et al.
+This section addresses RQ4 by exploring how chaos engineering has evolved in both academic and industry contexts.
+Specifically, it first presents the classification of research in chaos engineering, then examines the number of sources published per year and venue type (RQ4.1), and finally reviews the venue rankings and key contributors in the field (RQ4.2).
+7.1. Number of Sources by Research Type (RQ4.1)
+To better understand the current state of chaos engineering research (addressing RQ4.1), we classified the primary studies based on the research classification framework proposed by Wieringa et al.
 (Wieringa et al., 2006).
 This framework includes the following categories:
-  • Evaluation research: These studies investigate problems in software engineering practice, often uncovering new knowledge about causal or logical relationships among different phenomena.
-This type of research is conducted through case or field studies, field experiments, and surveys.
-  • Philosophical papers: These studies introduce new conceptual frameworks or ways of thinking.
-  • Solution proposal papers: These studies suggest techniques and argue for their relevance without full-scale validation.
-  • Validation research: These studies rigorously examine the properties of a proposed solution that has yet to be implemented in practice, employing methodologically sound approaches such as experiments and simulations.
-  • Opinion papers: These studies present the author’s views on what is good or bad about a particular issue and how things should be approached.
-  • Personal experience papers: These studies describe what was done based on the author’s own experience with one or more projects rather than explaining why it was done.
+  • Evaluation research: Investigates real-world problems through empirical methods such as case studies and surveys.
+  • Philosophical papers: Propose new conceptual models, frameworks, or perspectives.
+  • Solution proposals: Present novel techniques or tools, typically without full empirical validation.
+  • Validation research: Rigorously assess proposed solutions in controlled environments (e.g., simulations, experiments).
+  • Opinion papers: Express the author’s viewpoint on a topic, often arguing for or against certain practices.
+  • Experience reports: Describe applied practices and lessons from real-world projects.
 The first author did the classification, which was then reviewed by the second author, with all discrepancies resolved through discussion.
-The results of this classification, shown in Figure 12, indicate that solution proposal papers were the most frequent type of research between 2019 and 2024.
-In the academic literature, solution proposal papers dominated, with nine papers in 2021 and 2022 and slightly fewer (eight) in 2023.
-The grey literature, on the other hand, showed a marked rise in experience reports, growing from six reports in 2022 to eight in 2023.
-This classification highlights research activities’ evolving distribution and focus, emphasizing the dominance of solution proposals in academic literature and the increasing role of practical experience reports in grey literature.
-6.2. Number of Sources by Year and Venue Type
-We plot the distribution of chaos engineering studies from 2019 to 2024 to examine its adoption pattern (see Figure 10).
-In 2019, academic publications were minimal, limited to conferences and workshops, while grey literature focused on blogs, reflecting an early exploratory phase.
-By 2020, academic interest grew, focusing on conferences, though grey literature centered around blogs.
-In 2021, academic publications peaked across journals, conferences, and workshops, marking the “early adopter” phase, while grey literature remained stable, illustrating the challenges of Moore’s Chasm (see Figure 1).
-In 2022, academic output slightly declined.
-Still, grey literature grew, diversifying with blogs and white papers, suggesting chaos engineering was beginning to bridge Moore’s Chasm and gain broader acceptance.
-By 2023, academic publications remained steady, though workshops declined, while grey literature expanded further, especially with white papers, signaling continued industry engagement.
-The slight decline observed in 2024 for both academic and grey literature suggests a possible temporary slowdown, as the year is not yet complete, and more publications may emerge (Werbińska-Wojciechowska and Winiarska, 2023).
-A precise distribution emerges between academic and grey literature publications when examining venue types.
-Of the 44 scholarly publications, 28 (63.64%) were conference papers, 11 (25%) were journal articles, and 5 (11.36%) were workshops, reflecting a strong preference for conferences as the primary channel for academic dissemination.
-Among the 44 grey literature publications, 41 (93.18%) were blogs, while 3 (6.82%) were white papers, underscoring blogs as the dominant platform for informal knowledge sharing.
-These trends indicate that chaos engineering, which began as an emerging research area, has evolved into a mature field of interest (Varela-Vaca and Quintero, 2021), attracting significant attention from academia and industry.
+The results of this classification, shown in Figure 11, indicate that solution proposal papers were the most frequent type of research between 2016 and 2024.
+In the academic literature, solution proposal papers were the most prevalent, with nine papers in 2021 and 2022, and slightly fewer (eight) in 2023.
+On the other hand, the grey literature showed a slight upward trend in experience reports, rising from six in 2022 to eight in 2023.
+This classification highlights the evolving distribution and focus of research activities, emphasizing the dominance of solution proposals in academic literature and the increasing role of practical experience reports in the grey literature.
 Figure 9. Distribution of Academic and Grey Literature Studies over Years and Venue Type.
 […]
-Distribution of academic(N = 44) and grey(N = 44) literature studies over years and venue type.
-6.3. Publication trends and key Contributors in Academic and Grey Literature
+Distribution of academic(N = 49) and grey(N = 47) literature studies over years and venue type.
+7.2. Number of Sources by Year and Venue Type
+To further address RQ4.1, we analyzed the yearly distribution of academic and grey literature on chaos engineering from 2016 to early 2024 (see Figure 9).
+Academic output was limited from 2016 to 2018 but grew significantly from 2019, peaking in 2021 and 2022 with 13 and 12 studies, respectively, mainly in the form of conference papers.
+Grey literature remained minimal until 2019, rising steadily to a peak of 12 sources in 2023, dominated by blog posts and a few white papers.
+A slight drop in 2024 is likely due to the incomplete year (Werbińska-Wojciechowska and Winiarska, 2023).
+Overall, these trends underscore the growing scholarly and industrial interest in chaos engineering.
+Further breakdowns of publication types, venues, and source classifications are available in our online appendix (Section 3.6).
+7.3. Publication trends and key Contributors in Academic and Grey Literature (RQ4.2)
 We analyzed publication trends in chaos engineering by examining venues based on their type and ranking and identifying key contributors within academic and grey literature.
@@
 7.3.1. Academic Literature
-First, we analyzed venues based on their name, ranking, and type (journal or conference).
-As shown in Table 13, prominent venues include key publishers such as IEEE and ACM, which are widely recognized for hosting high-impact conferences and journals in chaos engineering.
+First, we analyzed publication venues based on their names, types (journal or conference), and rankings.
 The venues were categorized using the Computing Research and Education Association of Australasia (CORE) ranking for conferences (CORE 2023) [^4] https://portal.core.edu.au/conf-ranks/ and the Scimago ranking for journals (Scimago 2023) [^5] https://www.scimagojr.com/journalrank.php.
-Figure 12 presents the distribution of sources in these venues.
-The data indicates that the largest share of studies is found in unranked venues (12 sources), with a significant number published in Q1 venues reflecting solid academic interest.
-High-ranking venues, such as A and A*, include several prominent IEEE and ACM conferences, emphasizing the quality of research in the field.
-Lower-ranked venues, including B, C, and workshops, also contribute to the research landscape, although fewer studies are published in Q2 venues.
-Table 13 highlights the top 10 primary studies in chaos engineering, ranked by citation count.
-It should be noted that only the first author of each study is listed, even if there are multiple authors  (Varela-Vaca and Quintero, 2021).
-Study A42 from Sun Yat-Sen University, China, leads the list, presenting MicroRank, a tool for end-to-end latency issue localization in microservice environments.
-Following closely are studies A02 and A09 from the Hasso Plattner Institute in Germany, which explore chaos engineering applications in cloud security and continuous auditing and threat detection across multi-cloud environments.
-KTH Royal Institute of Technology, Sweden, is represented by three influential works, A03, A05, and A06, investigating fault injection and resilience in containerized systems and Java Virtual Machine (JVM) environments.
-Netflix contributes with study A36, focusing on automating chaos experiments within production environments, while Carnegie Mellon University’s A22 emphasizes service-level fault injection testing via the Filibuster tool.
-Additional contributions include A26 from Lund University, which presents an implementation framework for chaos engineering, and A35 from the University of Stuttgart, which proposes an approach for prioritizing chaos experiments using established risk analysis techniques.
-This collection of studies from institutions in countries such as China, Germany, Sweden, and the United States reflects a global commitment to advancing chaos engineering with applications to enhance system resiliency, strengthen cloud infrastructure, and improve reliability in production environments worldwide.
+As shown in Figure 11, the largest share of studies appeared in unranked venues (12), though many were published in top-tier sources—particularly CORE A/A* conferences and Scimago Q1 journals, often from IEEE and ACM.
+This suggests increasing scholarly validation of chaos engineering research.
+Lower-ranked venues (B, C) and workshops contributed modestly, with few studies published in Q2 venues.
+To highlight influential contributions, we identified the top 10 academic studies based on citation counts from Scopus and Google Scholar (first author only).
+The leading study, authored by Netflix (Basiri et al., 2016), has become foundational to the field, followed by diverse works from institutions in China, Germany, Sweden, and the U.S.
+that span topics such as end-to-end latency localization (e.g., MicroRank) (Yu et al., 2021), chaos-based security testing (Torkura et al., 2020), and exception handling in JVM environments (Zhang et al., 2019).
+These publications reflect a global effort toward improving system resiliency and reliability through chaos engineering.
+Full details are provided in our online appendix (Section 3.6).
 7.3.2. Grey Literature
-Unlike traditional academic literature, which follows a structured format and is ranked using systems such as h-index or Scimago, grey literature lacks standardized organization.
-It cannot be assessed using typical academic metrics.
-Instead, alt-metrics (alternative metrics used to measure online attention and engagement), such as comments, likes, shares, and social engagement, are used to gauge the quality and impact of grey literature.
-Table 14 highlights the top five articles in grey literature based on these alt-metrics.
-Grey literature on chaos engineering is often published by well-known IT companies (e.g., Microsoft, Gremlin, and Capital One) on company blogs or online publishing platforms (e.g., LinkedIn and Medium).
-These platforms are essential venues for sharing industry-driven insights in formats such as technical articles, blog posts, and white papers.
-The alt-metrics used, such as likes, comments, and shares, help measure the influence and engagement of the content within the industry, indicating its impact on chaos engineering.
-In cases where social engagement data was limited, additional quality assessments were considered, including the reputation of the venues, the authors’ expertise, and the clarity of the content (see Table  3).
-As shown in Table 14, the top five grey articles are authored by industry professionals who provide valuable insights into chaos engineering.
-Key contributors come from prominent organizations, including Lyft, National Australia Bank (NAB), Expedia, Chaos Toolkit, and Eat Takeaway.com.
-For example, the study G12 from Lyft, published on Medium, received 348 likes and 2 comments for its introduction of a chaos experimentation framework.
-The study G11 from NAB, also published on Medium, received 156 likes for its focus on enhancing observability in microservices during chaos experiments.
-Additional notable contributions include the study G15 from Expedia, whose article on Medium earned 128 likes for detailing Kubernetes-integrated chaos engineering frameworks.
-The study G25, authored on LinkedIn by Chaos Toolkit, received 52 likes and 52 comments, emphasizing the importance of human interaction in chaos experiments.
-Lastly, the study G14 from Just Eat Takeaway.com, published on Medium, received 24 likes and 1 comment, highlighting the discovery of hidden system vulnerabilities through chaos engineering.
-These industry contributions have significantly influenced the development of chaos engineering practices, providing practical insights directly from leaders in the field.
-As discussed earlier, the same quality assessment criteria were applied, focusing on venue reputation, author expertise, and content clarity, when data on social engagement were insufficient (see Table  3).
-Table 13. Top-10 of Papers Ordered by Citations in Academic Literature (SC - SCOPUS, GS - Google Scholar).
-Rank ID Title Author Affiliation SC GS
-1 A42 MicroRank: End-to-End Latency Issue Localization with Extended Spectrum Analysis in Microservice Environments Guangba Yu Sun Yat-Sen University, China 50 72
-2 A02 CloudStrike: Chaos Engineering for Security and Resiliency in Cloud Infrastructure Kennedy A. Torkura Hasso Plattner Institute, Germany 39 70
-3 A09 Continuous Auditing and Threat Detection in Multi-cloud Infrastructure Kennedy A. Torkura Hasso Plattner Institute, Germany 28 53
-4 A03 A Chaos Engineering System for Live Analysis and Falsification of Exception-Handling in the JVM Long Zhang KTH Royal Institute of Technology, Sweden 15 44
-5 A05 Observability and Chaos Engineering on System Calls for Containerized Applications in Docker Long Zhang KTH Royal Institute of Technology, Sweden 17 40
-6 A36 Automating Chaos Experiments in Production Ali Basiri Netflix in Los Gatos, California. 15 40
-7 A26 Getting Started with Chaos Engineering - Design of an Implementation Framework in Practice Hugo Jernberg Lund University and ICA Gruppen AB, Lund/Solna, Sweden 7 21
-8 A22 Service-level Fault Injection Testing Christopher S. Meiklejohn Carnegie Mellon University, Pittsburgh, PA, United States 14 19
-9 A06 Maximizing Error Injection Realism for Chaos Engineering With System Calls Long Zhang KTH Royal Institute of Technology, Sweden 6 18
-10 A35 Identifying and Prioritizing Chaos Experiments by Using Established Risk Analysis Techniques Dominik Kesim University of Stuttgart, Germany 8 16
-Table 14. Top-5 Articles Ordered by Altmetrics (Comments / Likes/Claps) in Grey Literature.
-Rank ID Comments Likes/Claps Author Affiliation Venue
-1 G12 2 348 Kathan Shah & Subhra Mondal Lyft Medium
-2 G11 - 156 Prateek Sachan National Australia Bank (NAB) Medium
-3 G15 - 128 Nikos Katirzis Expedia Group Medium
-4 G25 - 52 Russ Miles Chaos Toolkit Medium
-5 G14 1 24 Mathew Moon Just Eat Takeaway.com Medium
+Unlike academic publications, grey literature lacks formal rankings and peer review.
+To evaluate its quality and impact, we utilized alternative metrics (e.g., likes, comments, shares) and, when social engagement data were limited, applied qualitative criteria such as venue reputation, author expertise, and content clarity (see Table 5).
+The most engaged grey literature source was “Chaos Monkey Upgraded” by Netflix  (Hochstein and Rosenthal, 2016), which received 443 likes and 2 comments.
+Other widely read posts include Lyft’s framework introduction (348 likes) (Shah, 2021), NAB’s work on observability (156 likes) (Mondal and Sachan, 2020), and technical contributions from Expedia (Katirtzis, 2022) and the Chaos Toolkit community (Miles, 2019a).
+These practitioner-authored works offer valuable, experience-based perspectives that complement academic research.
+Further details, including top-ranked articles and alt-metric data, are available in the online appendix (Section 3.6).
+[…]
+Research Type
 Figure 10. Research Type of Academic (AL) and Grey (GL) Literature Studies.
+[…]
+Venue Diagram
 Figure 11. Venue Ranking of Academic Literature Studies.
 8. Discussion
-In this section, we summarize the key findings of our review and discuss their implications.
-We also outline open issues in chaos engineering research and discuss potential threats to our study’s validity.
+In this section, we summarize the key findings of our review and discuss their implications across RQ1 to RQ4.
+We also outline open issues in chaos engineering research and discuss potential threats to the validity of our study.
 8.1. Implications of Findings
 The analysis highlights five key implications for researchers and practitioners in chaos engineering:
-  • Need for a Unified Definition: The literature reveals inconsistencies in defining chaos engineering.
-Although tech giants like Google, Amazon, and Netflix provide definitions that emphasize resilience testing through controlled failure injection, there is variation in focus across other companies, with some prioritizing production-level experimentation, self-healing capabilities, or anti-fragility.
-A unified definition can align research efforts and support the systematic adoption of chaos engineering practices, strengthening reliability across complex Information Technology (IT) environments (Siswanto et al., 2024; Jayakody and Wijayanayake, 2024; Jones, 2019).
-  • Complementary Role to Traditional Testing: Chaos engineering should be viewed as a complement to traditional testing rather than a replacement (Bocetta, 2019; Palani and Gupta, 2023).
-While conventional testing may miss subtle vulnerabilities within complex systems, chaos engineering exposes these issues through proactive experimentation.
-Recognizing chaos engineering as an additional layer to traditional methods emphasizes its role in enhancing system resilience, especially in high-stakes, mission-critical environments.
-Integrating chaos engineering into testing workflows is encouraged for those aiming to build more robust and reliable systems.
-  • Growing Alignment of Research Theory and Practice: The prevalence of solution-oriented papers in academic literature alongside experience reports in grey literature reflects an increasing alignment between research and practice in chaos engineering.
-This synergy allows researchers to refine models based on real-world applications while practitioners benefit from research-backed frameworks, thus bridging the gap between theory and implementation.
-As suggested in previous studies (Jernberg et al., 2020; Rosenthal and Jones, 2020), fostering industry-academia collaborations could further integrate academic rigor with practical insights, advancing chaos engineering as a more resilient and impactful field.
-  • Targeted Tool Selection: The study suggests that organizing chaos engineering tools by factors like execution environment and automation level might help practitioners make informed decisions when selecting tools.
-This taxonomy enables organizations to choose tools that match their specific infrastructure, operational needs, and risk tolerance, minimizing disruptions, managing costs, and enhancing resilience testing across development stages.
-This targeted approach allows tools to align with specific testing phases, from initial experimentation to production-level assessments, ensuring that chaos engineering efforts are both cost-efficient and impactful, ultimately supporting system reliability.
-  • Standardized Practices: The findings highlight the importance of a standardized approach in chaos engineering to ensure consistent and practical application across diverse environments.
-Following best practices, which contribute to reliable and predictable outcomes (Kumara et al., 2021), enables organizations to build resilience and systematically manage risks.
-These practices, such as starting with small, controlled experiments, setting clear objectives, and automating processes where possible, support organizations in achieving reliable outcomes crucial for managing complex systems.
-Conversely, neglecting these practices, such as misaligned experiment design or insufficient safety mechanisms, can lead to inconsistent testing and increased system vulnerability.
+  • Extending Reliability Testing through Chaos Engineering: Chaos engineering should be understood not as a replacement for established reliability or software testing techniques, but as an extension that focuses on system behavior under failure and uncertainty.
+While traditional methods, such as unit tests, integration tests, or stress tests, validate expected behaviors in controlled environments, chaos engineering introduces carefully scoped disruptions to examine how systems respond and recover from unexpected faults.
+This approach contributes a complementary dimension to reliability assessment, particularly in distributed or production-like environments where emergent failures often go undetected.
+Rather than displacing existing practices, chaos engineering enhances them by validating resilience under real-world operational stress.
+  • Targeted Tool Selection (RQ3.1): This study introduces a taxonomy that helps practitioners select chaos engineering tools by considering factors such as execution environment, automation mode, and deployment stage (see Table 13).
+The taxonomy is intended to support informed decision-making by aligning tool capabilities with the specific infrastructure and operational needs of an organization.
+In practice, a practitioner would begin by assessing the architectural characteristics of their system, such as whether it runs on Kubernetes, serverless platforms, or hybrid cloud, and then identify tools from the taxonomy that match these conditions and the intended stage of testing, from early experimentation to full production deployment.
+Applying the taxonomy effectively also requires careful consideration of how well a tool integrates with existing workflows and team expertise.
+Practitioners must ensure that selected tools support their resilience objectives without introducing unnecessary complexity or risk.
+For example, using a highly automated tool in an immature testing environment could result in poorly scoped experiments or overlooked failure modes.
+Likewise, choosing tools that lack compatibility with the target infrastructure may lead to inconclusive or misleading outcomes.
+By mapping tools to their most appropriate use cases, the taxonomy can help mitigate such risks and promote more efficient and targeted adoption of chaos engineering practices across diverse technical landscapes.
+  • Standardized Practices (RQ3.2): The findings highlight the importance of a standardized approach in chaos engineering to ensure consistent and practical application across diverse environments.
+Following best practices, which contribute to reliable and predictable resilience outcomes, such as faster recovery times and minimal service disruption (Kumara et al., 2021), enables organizations to build resilience and systematically manage risks.
+These practices, such as starting with small, controlled experiments, setting clear objectives, and automating processes where possible, support organizations in achieving reliable outcomes crucial for managing complex systems (see Table 14).
+Conversely, neglecting these practices, such as misaligned experiment design or insufficient safety mechanisms (see Table 15), can lead to inconsistent testing and increased system vulnerability.
 These findings underscore the need to balance adherence to best practices with avoiding common pitfalls, ensuring that chaos engineering aligns with broader operational goals to foster stability and reliability.
-  • Insights through Quantitative and Qualitative Fusion: An essential implication of the findings is that combining quantitative and qualitative evaluation methods provides a comprehensive framework for assessing and improving system stability in chaos engineering.
-Quantitative metrics offer objective data on performance under stress, such as error rates, MTTR, and availability.
+  • Insights through Quantitative and Qualitative Fusion (RQ3.3): Another key implication of our findings is that combining quantitative and qualitative evaluation methods enhances the assessment and improvement of system stability in chaos engineering.
+Quantitative metrics provide objective data on performance under stress, including error rates, mean time to resolution, and availability.
 At the same time, qualitative insights, including user experience and team feedback, reveal the broader impacts of system failures.
-This combined approach helps organizations identify technical vulnerabilities and understand their user and business impacts, enabling targeted investments in both stability and operational efficiency to align improvements with business goals and user expectations.
+This combined approach enables organizations to identify technical vulnerabilities and understand their user and business impacts, allowing for targeted investments in both system stability and operational efficiency.
+This alignment of improvements with business goals and user expectations facilitates the development of solutions that effectively enhance system resilience, reduce failure impact, and improve user experience.
+  • Growing Alignment of Research Theory and Practice (RQ4): Our review empirically demonstrates the growing alignment between academic research and industry practice in chaos engineering.
+By classifying 96 primary sources, including both peer-reviewed papers and grey literature, we observed that solution proposal papers dominated academic publications between 2016 and 2023, while grey literature during the same period showed a slight increase in experience reports (see Figure 11).
+This dual trend reveals a convergence between theoretical contributions and practitioner-led experimentation, further reflected in the types of publication venues: top-ranked academic venues published novel frameworks and techniques, while widely cited blog posts detailed real-world implementation case studies.
+By capturing both perspectives, our study illustrates how academic research and industry practices are increasingly aligned, with researchers drawing insights from real-world applications and practitioners adopting concepts that resonate with emerging scholarly work to guide system resilience strategies.
+This evolving interplay signals an opportunity for deeper collaboration across academia and industry to accelerate innovation in chaos experiment design, such as advancements in automated fault orchestration, address implementation gaps, such as the lack of robust rollback or abort mechanisms in several research prototypes, and broaden the real-world adoption of chaos engineering.
 8.2. Open Research Issues and Practical Challenges in Chaos Engineering
-Our review identifies several unresolved issues in chaos engineering, including research gaps and organizational challenges.
-We examine these issues from three perspectives: people, process, and technology.
-7.2.1. People
-In chaos engineering, people include engineers, managers, and other team members who collaborate to design and run experiments, ensuring that the systems are thoroughly tested (Moon, 2022; Nikolaidis et al., 2021; Graglia, 2024).
-First, some organizations struggle with poor communication and collaboration between teams, leading to incomplete or ineffective experiments  (Zhang et al., 2019).
-Encouraging more vital teamwork and shared ownership could improve results by fostering better department communication (Gill, 2021; Katirtzis, 2022; Kostic, 2024).
-Another challenge is the shortage of specialists, making it difficult for organizations to carry out chaos experiments.
-This gap could be addressed by creating comprehensive training programs and certifications that allow staff to develop the necessary expertise (Dedousis et al., 2023; Nikolaidis et al., 2021; Yu et al., 2021).
-Finally, cultural resistance is expected, as fear of failure and blame often hinder wider adoption  (Simonsson et al., 2021; Rosenthal and Jones, 2020).
-Implementing cultural change initiatives that promote learning from controlled failures and reduce fear of backlash can create a more open environment for experimentation (Green, 2023).
-7.2.2. Process
-The process aspect refers to the specific steps and methodologies, such as planning, testing, and reporting, that guide the implementation of chaos experiments (Bairyev, 2023; Kostic, 2024; Olmstead, 2023).
-First, the lack of standardized guidelines across industries makes consistent application difficult (Nikolaidis et al., 2021; Cotroneo et al., 2022; Kostic, 2024).
-Developing a set of best practices and frameworks would help organizations measure progress and improve the effectiveness of these practices (Shah, 2021; Bremmers, 2021).
-The research shows that many organizations struggle to prioritize chaos experiments alongside daily operations, often finding it difficult to allocate time and resources while maintaining operational stability.
-Establishing transparent processes for scheduling and running these experiments during non-critical periods or as part of routine maintenance can help mitigate the issue (Kesim et al., 2020).
-7.2.3. Technology
-Technology refers to the systems and infrastructure used to design, execute, and monitor failure scenarios, supporting and amplifying human efforts and processes (Graglia, 2024; Olmstead, 2023).
-First, chaos experiments are resource-intensive and require significant computational power and storage.
-This makes it difficult for organizations to allocate infrastructure without affecting critical operations  (Dedousis et al., 2023; Nikolaidis et al., 2021).
-To manage this, companies can schedule experiments during low-traffic periods or use scalable cloud resources to reduce strain  (Kostic, 2024; Gill, 2021).
-Second, deploying chaos tools requires specialized knowledge, making it challenging to scale experiments consistently (Zhang et al., 2019; Yu et al., 2021).
-This challenge can be alleviated with better documentation, training, and platforms that simplify deployment for teams with less expertise (Bairyev, 2023).
-Third, many chaos engineering tools lack automation, making setup and execution time-consuming and prone to errors.
-Automating processes like fault injection and monitoring can reduce the burden on engineers, minimize mistakes, and improve scalability.
-Organizations can address this through commercial tools or internal solutions (Bairyev, 2023; Cotroneo et al., 2022; Shah, 2021).
-Finally, while Artificial Intelligence (AI) holds promise for further streamlining experiment design and execution, research in this area is still limited.
-Using AI could reduce errors, improve efficiency, and help organizations scale their resilience efforts more effectively (Zhang et al., 2023).
+Our review highlights several ongoing challenges that limit the adoption and scalability of chaos engineering.
+These include cultural resistance, skill gaps, and resource overhead.
+The following subsections outline three key themes based on current literature and practice.
+8.2.1. Organizational and Cultural Challenges
+The findings highlight that organizational resistance, risk aversion, and a blame culture can pose significant barriers to the adoption of chaos engineering (Krishnamurthy, 2021; Bocetta, 2019).
+Although chaos engineering promotes resilience, many teams perceive fault injection, especially in production, as disruptive and risky (Sbiai et al., 2023; Zhang et al., 2019; Dedousis et al., 2023).
+This perception discourages widespread adoption, as teams are often reluctant to participate in experiments that could disrupt production or damage service reliability (Krishnamurthy, 2021; Ji et al., 2023).
+Adopting chaos engineering requires a mindset shift: from protecting systems at all costs to embracing failure as a tool for improvement (Bocetta, 2019; Bremmers, 2021).
+However, this shift is rarely supported by leadership and is often undermined by siloed teams, poor communication, and inconsistent reliability practices (Katirtzis, 2022; Dua, 2024; Abdul, 2024).
+Without cultural alignment and collaboration, chaos engineering remains confined to isolated efforts with limited organizational impact (Sbiai et al., 2023; Tivey et al., 2019).
+While these challenges are frequently noted in industry reports, academic research has yet to explore them systematically.
+There is a need for empirical studies that examine how factors like culture, team structure, and leadership influence chaos engineering adoption.
+Research could also contribute by developing and validating readiness frameworks or maturity models to guide organizations in scaling these practices effectively.
+8.2.2. Skill and Expertise Gaps
+Chaos Engineering requires specialized skills that many teams currently lack, creating a major barrier to adoption (Torkura et al., 2020; Kostic, 2024).
+First, effective implementation depends on understanding distributed systems, runtime behavior, and formal models like metric temporal logic, areas in which most practitioners have limited training (Frank et al., 2023; Nikolaidis et al., 2021).
+Next, managing experiment configurations, such as complex YAML files with interdependent settings, demands precision and familiarity with system internals (Kassab et al., 2022).
+Furthermore, many chaos engineering practices rely on intuition rather than standardized procedures, leading to inconsistent outcomes and reduced reliability (Nikolaidis et al., 2021; Vinisky, 2024).
+Adding to this challenge, some experiments require interaction with low-level components, such as kernel modules or orchestration layers, which increases the technical entry barrier (Katirtzis, 2022; Mooney, 2023).
+Although these challenges are recognized in both academic and practitioner sources, there is limited empirical research on how teams acquire and apply these skills in practice.
+Future work should focus on identifying specific knowledge gaps, evaluating effective training strategies, and designing practical tools and reusable templates to support broader and more consistent adoption (Bairyev, 2023; Cotroneo et al., 2022).
+8.2.3. Resource and Operational Constraints
+Chaos engineering demands significant time, computing resources, expertise, and cost, which often limits routine adoption  (Vu et al., 2022; Poltronieri et al., 2022; Zhang et al., 2021a).
+Setting up environments, managing telemetry, and iterating experiments require sustained effort, especially for smaller or time-constrained teams (Gill, 2021; Zhang et al., 2021a).
+Deploying testbeds can take considerable time and need expert oversight (Kassab et al., 2022).
+These challenges scale with system size and are intensified by limited infrastructure, cloud expertise, and the financial overhead of tooling and operations (Katirtzis, 2022; Durai, 2022; Kostic, 2024).
+As systems become increasingly complex, chaos engineering becomes more challenging to scale due to service heterogeneity and coordination overhead.
+Security and data concerns also arise.
+For example, experiments may expose vulnerabilities and generate sensitive telemetry.
+Managing this data poses storage, privacy, and analysis challenges.
+These factors—including time, complexity, and cost—often clash with delivery timelines, leading to chaos engineering being deprioritized despite its long-term value (Krivas and Portela, 2020; Kostic, 2024; Bairyev, 2023).
+Although frequently reported in practice, there is little empirical research on how teams manage these constraints.
+Future work could investigate how organizations reduce the setup overhead of test environments or prioritize experiments to align with available resources.
+Such studies would help tailor chaos engineering to practical operational limits.
 8.3. Threats to Validity
-We considered several potential threats to the external, construct, internal, and conclusion validity (Wohlin et al., 2012) that may impact our study.
+We considered potential threats to the external, construct, internal, and conclusion validity (Wohlin et al., 2012) that may impact our study.
 8.3.1. Threats to External Validity
-External validity refers to the extent to which the study results are applicable in a broader context (Wohlin et al., 2012).
-Since our multivocal review incorporated studies from various online sources, our findings may have limited applicability across the wider landscape of practices and general disciplines in chaos engineering, thus posing a threat to external validity.
-To enhance the external validity of our findings, we conducted multiple feedback sessions throughout the analysis of the selected studies.
-The discussions of each session were analyzed, providing qualitative insights that allowed us to refine our research methods and the applicability of our results.
-Furthermore, we compiled an online appendix with all artifacts generated during our analysis, including a complete list of sources, codes, and groups (see Section  2), making them publicly accessible for further examination.
-Another potential threat is the risk of missing relevant literature (grey and scientific) due to alternative terminology for related concepts that may not match our search terms.
-We explicitly included relevant associated synonyms and terms in our search strings to address this threat.
-We also leveraged search engine features that inherently consider variations of terms in a search string.
-The items identified were rigorously assessed based on multiple quality dimensions (Wohlin, 2014).
-Finally, while there are numerous chaos engineering tools, we limited our analysis to ten due to practicality.
-To mitigate this limitation, we selected the ten most relevant tools currently used by practitioners.
-A detailed comparison of these chosen tools may provide additional insight, which we propose for future research.
+External validity refers to the extent to which our findings can be generalized beyond the specific context of our study (Wohlin et al., 2012).
+While this multivocal literature review (MLR) included both academic and grey sources, there is a possibility that some perspectives or practices in chaos engineering may not have been captured.
+A particular threat arises from the risk of excluding relevant literature that uses alternative terminology, such as resilience testing or fault injection, rather than chaos engineering.
+While we crafted inclusive search strings using multiple synonyms and related terms (see Table 3), some relevant works may still have been omitted.
+This limitation is common in MLRs and is explicitly acknowledged.
+Future studies may improve coverage by expanding search terms to encompass other reliability testing practices.
+Another limitation stems from the taxonomy itself, which has not yet undergone formal external validation.
+Although developed through thematic analysis and informed by structured internal review sessions, it remains untested in practical or industry settings.
+As such, its generalizability should be interpreted with caution until further validation studies are conducted.
+Finally, our tool analysis focused on ten widely adopted chaos engineering tools due to feasibility constraints.
+While these tools represent current practitioner usage, broader tool coverage, including lesser-known or homegrown solutions, could yield deeper insights.
+We consider this an avenue for future work.
 8.3.2. Threats to Construct and Internal Validity
-Construct validity pertains to the generalizability of the study’s constructs, while internal validity relates to the credibility of the methods used for data analysis, including possible biases (Wohlin et al., 2012).
-To reduce potential threats to construct and internal validity, we implemented multiple triangulation rounds, inter-rater reliability assessments, and quality control measures (see Section 2).
-These methods helped minimize potential biases, such as observer and interpretive biases, thereby enhancing the validity of our data analysis.
-Additional studies and codes were incorporated, although no further inter-rater evaluation was conducted.
-In particular, some risk of observer bias remains inherent in this methodology (Kumara et al., 2021).
+Construct validity concerns whether the study accurately captures the concepts it aims to investigate, while internal validity refers to the soundness of the methods used and the control over potential biases (Wohlin et al., 2012).
+To address these concerns, we conducted at least four structured feedback sessions during our analysis.
+Discussions from these sessions were qualitatively examined and used to fine-tune both our methodological choices and the relevance of our findings.
+We have also compiled an online appendix (see Section 3.6) that contains all key artifacts, including the full list of sources, extracted codes, thematic groupings, and taxonomy classifications, to ensure transparency and reproducibility.
+We also adopted triangulation by synthesizing both academic and grey literature sources, thereby reducing our reliance on a single perspective.
+Inter-rater reliability was formally assessed during the study selection phase, using Cohen’s Kappa, which indicated strong agreement for both academic and grey literature.
+This ensured consistency in applying inclusion and exclusion criteria.
+The first author conducted manual coding, which was independently reviewed by the second author.
+Disagreements were resolved collaboratively through discussion.
+This collaborative process was applied throughout all coding phases to ensure consistency and reduce interpretation bias.
+Despite these precautions, some risk of observer bias remains inherent in qualitative research.
+However, we believe that our triangulation of sources, structured collaboration, and transparent artifact release significantly mitigate this threat.
 8.3.3. Threats to Conclusion Validity
-Conclusion validity refers to the extent to which the conclusions of the study are credibly supported by the available data  (Wohlin et al., 2012).
-To address this threat, we used thematic coding and inter-rater reliability assessments to reduce observer and interpretive biases, aiming to analyze the data thoroughly.
-Furthermore, the study conclusions were independently derived by different authors and subsequently reviewed and validated against the primary studies selected during a feedback session.
-8. Related Surveys
-In this section, we consider key surveys on resilience, robustness, and fault injection techniques, exploring their relationship to the proactive focus of chaos engineering.
-For example, Natella et al.
-conducted an extensive survey on Software Fault Injection (SFI) techniques, categorizing methods by fault types, injection approaches, and application domains.
-These techniques are valuable for assessing software reliability in controlled settings.
-However, SFI techniques are generally applied in pre-production or isolated environments rather than in the live, iterative fault-testing scenarios central to chaos engineering (Natella et al., 2016).
-Another study, conducted by Colman-Meixner et al.
-(Colman-Meixner et al., 2016), explores resilience strategies across cloud infrastructure layers, such as physical servers, virtualization, and application components.
-Their work proposes a multi-layered framework with redundancy and failover mechanisms to ensure continuity during failures.
-While this approach aligns with the resilience goals of chaos engineering, the focus remains on structured recovery strategies within controlled environments rather than proactive fault injection in live settings.
-Laranjeiro et al.
-(Laranjeiro et al., 2021) reviewed robustness assessment techniques for systems such as operating systems and middleware and discussed methods such as fault injection and API testing to evaluate dependability under stress.
-These approaches aim to identify weaknesses but emphasize static testing, which does not include the proactive, live-system experiments that chaos engineering relies on.
-More recently, Rouholamini et al.
-(Rouholamini et al., 2020) presented a study on proactive self-healing techniques in cloud computing, categorizing strategies such as automated fault detection, anomaly recognition, and self-correction mechanisms.
-These techniques improve system resilience by addressing faults as they arise, but differ from the chaos engineering approach of using intentional fault injection to reveal latent vulnerabilities preemptively.
-In general, there is a significant knowledge gap in the area of proactive and iterative resilience testing directly within live production environments.
-Moreover, there are no comprehensive reviews of the literature on chaos engineering.
-Therefore, this review paper systematically analyzes, assesses, and summarizes the existing literature to clarify the unique contributions of chaos engineering to the domain of system resilience.
+Threats to conclusion validity concern the degree to which the study’s conclusions are reasonably based on the data analyzed (Wohlin et al., 2012).
+To mitigate this, we applied thematic coding using ATLAS.ti tool, and employed a structured analysis process to reduce observer and interpretation biases.
+Coding and classification were guided by a consensus-driven protocol between the first and second authors, ensuring consistency in the derivation of findings.
+Both authors independently formulated the conclusions presented in this study and later validated them through collaborative discussion.
+These conclusions were cross-checked against the original source material to ensure they remained grounded in the data.
+While our synthesis integrated both academic and grey literature, we acknowledge that our interpretations are limited to the scope of the selected articles.
+In particular, the absence of external expert validation for the taxonomy remains a limitation.
+Future research may benefit from follow-up evaluations, broader empirical testing, or practitioner feedback to confirm the applicability of our findings in real-world settings.
+We also acknowledge that our results are limited to the quality of the grey literature and peer-reviewed publications, which we mitigate by conducting a quality assessment analysis.
 9. Conclusion and Future Work
 As modern distributed systems become increasingly complex, chaos engineering offers a proactive approach to ensure stability by introducing controlled failures that reveal system weaknesses, prevent outages, and support high performance under stress.
-This study bridges the gap between research insight and industry practices through a multivocal literature review (MLR) of 88 sources from academic and grey literature.
+This study bridges the gap between research insight and industry practices through a multivocal literature review (MLR) of 96 sources from academic and grey literature.
 To promote clarity across academia and industry, we proposed a unified definition of chaos engineering, reinforcing its role as a vital complement to traditional testing.
-We identified the main functionalities of a chaos engineering platform and compiled a set of principles that can guide systematically planning and conducting chaos experiments.
-In addition, we presented the key components of a chaos engineering platform and identified 12 quality requirements to guide the selection and evaluation of the platform.
+We identified the main functionalities of a chaos engineering platform and compiled a set of activities that can guide systematic planning and conducting chaos experiments.
+In addition, we presented the key components of a chaos engineering platform and identified eleven quality requirements to guide the selection and evaluation of the platform.
 We also identified eight qualitative and four quantitative metrics essential for evaluating the impact of chaos experiments, ensuring that chaos engineering meaningfully enhances system robustness.
-Our review highlights four technical and three socio-technical challenges that chaos engineering seeks to resolve.
+Our review highlights three technical and four socio-technical challenges that chaos engineering seeks to resolve.
 To support organizations in structuring chaos engineering practices, we developed a taxonomy that organizes tools and techniques by environment, automation mode, automation strategy, and deployment stage while identifying the ten most commonly used tools within this framework.
@@
 These practices and potential pitfalls guide organizations in optimizing their chaos engineering initiatives to foster safer and more robust systems.
-Our findings also reveal that solution-oriented research in academia and practical experience in industry converge, underscoring chaos engineering’s essential role in strengthening modern distributed systems.
-Finally, we outlined several research and practical challenges that must be addressed to realize chaos engineering’s vision entirely.
+Our findings also reveal that solution-oriented research in academia and practical experience in industry converge, underscoring the essential role of chaos engineering in strengthening modern distributed systems.
+Finally, we outlined several research and practical challenges that must be addressed to realize the vision of chaos engineering fully.
 We believe that these identified open issues can guide future research in the domain of chaos engineering.
@@
 This exploration will complement the ten tools analyzed in this literature review, offering broader insights into the prevalence and practical application of chaos engineering in diverse projects.
-Additionally, we aim to apply chaos engineering to improve the robustness of AI-enabled systems by evaluating how such systems handle disruptions in accuracy, adaptability, and consistency, ensuring that they continue to deliver reliable performance even under challenging conditions.
+Additionally, we aim to apply chaos engineering to enhance the robustness of AI-enabled systems by evaluating how these systems respond to disruptions in accuracy, adaptability, and consistency, ensuring they continue to deliver reliable performance even under challenging conditions.